### v1

# Object-Centric Image to Video Generation with Language Guidance

Abstract
Accurate and flexible world models are crucial for autonomous systems to understand their environment and predict future events. Object-centric models, with structured latent spaces, have shown promise in modeling object dynamics and interactions, but often face challenges in scaling to complex datasets and incorporating external guidance, limiting their applicability in robotics. To address these limitations, we propose TextOCVP, an object-centric model for image-to-video generation guided by textual descriptions. TextOCVP parses an observed scene into object representations, called slots, and utilizes a text-conditioned transformer predictor to forecast future object states and video frames. Our approach jointly models object dynamics and interactions while incorporating textual guidance, thus leading to accurate and controllable predictions. Our method’s structured latent space offers enhanced control over the prediction process, outperforming several image-to-video generative baselines. Additionally, we demonstrate that structured object-centric representations provide superior controllability and interpretability, facilitating the modeling of object dynamics and enabling more precise and understandable predictions. Videos and code are available at https://play-slot.github.io/TextOCVP/.
Object-Centric Learning, Text-Conditioned Video Prediction, Object-Centric Video Prediction, Representation Learning, Transformers
Refer to captiona) Object-Centric DecompositionLearned ObjectsRefer to captionb) Text-Conditioned Image-to-Video Generation
‘Put the green block
in the cyan bowl.’
$\mathcal{C}$
Figure 1: Overview of TextOCVP. (a) Our model first parses a reference frame $\textbf{X}_{1}$ into its object components $\textbf{S}_{1}$. (b) Our TextOCVP predictor jointly models object dynamics and interactions while integrating textual guidance, generating future object states and video frames that align with the provided textual instructions.
1 Introduction
Understanding and reasoning about the real world is essential for enabling autonomous systems to better comprehend their surroundings, predict future events, and adapt their actions accordingly. Humans achieve these capabilities by perceiving the environment as a structured composition of individual objects that interact and evolve dynamically over time (Kahneman et al., 1992). Neural networks equipped with such compositional inductive biases have shown the ability to learn structured object-centric representations of the world, which enable desirable properties, such as out-of-distribution generalization (Dittadi et al., 2022), compositionality (Greff et al., 2020), or sample efficiency (Mosbach et al., 2024).
Recent advances in unsupervised object-centric representation learning have progressed from extracting object representations in synthetic images (Locatello et al., 2020; Lin et al., 2020b) to modeling objects in video (Kipf et al., 2022; Singh et al., 2022) and scaling to real-world scenes (Seitzer et al., 2023; Zadaianchuk et al., 2024). These developments have enabled object-level dynamics modeling for future prediction and planning. Notably, models like SlotFormer (Wu et al., 2023) or OCVP (Villar-Corrales et al., 2023) introduced object-centric world models that explicitly model spatio-temporal relationships between objects, shifting away from image-level approaches that ignore scene compositionality. Despite these advancements, current approaches struggle with complex object appearances and dynamics, and lack the ability to incorporate external guidance, thus limiting their effectiveness as world models in robotic applications.
To address these challenges, we propose TextOCVP, a novel object-centric model for image-to-video generation with language guidance, illustrated in Fig. 1. Given a reference image and text instruction, TextOCVP extracts object representations and predicts their evolution using a novel text-conditioned object-centric transformer. This module forecasts future object representations by modeling object dynamics and interactions, while also incorporating textual information through a text-to-slot attention mechanism. By jointly modeling spatio-temporal object relationships and integrating textual guidance, TextOCVP generates future object states and video frames that align with the provided instructions.
We evaluate our approach through extensive experiments focusing on image-to-video generation tasks with varying levels of complexity. Our results show that TextOCVP outperforms existing text-conditioned methods by effectively leveraging object-centric representations. This demonstrates the significant advantage of incorporating textual guidance into structured object-centric representations, particularly for scenes featuring multiple moving objects.
Through an in-depth model analysis, we demonstrate the importance of our proposed object-centric latent space, which allows TextOCVP to generate video sequences that closely align with language instructions. We further demonstrate how TextOCVP adapts its generated video sequences based on different captions by routing the text information to the relevant object representations.
In summary, our contributions are as follows:
  • We propose TextOCVP, a novel text-driven image-to-video generation model, featuring a text-conditioned object-centric predictor that integrates textual guidance via a text-to-slot attention mechanism.
  • Through extensive evaluation, we show that TextOCVP outperforms existing text-conditioned models by leveraging object-centric representations.
  • We demonstrate that TextOCVP is controllable, seamlessly adapting to diverse textual instructions.
2 Related Work
2.1 Object-Centric Learning
Representation learning, the ability to extract meaningful features from data, often improves model performance by enhancing its understanding of the input space (Bengio et al., 2013). Object-centric representation learning focuses on simultaneously learning representations that characterize individual objects within an input image or video.
Recently, object-centric models have progressed from learning object representations from synthetic images (Greff et al., 2019; Lin et al., 2020a; Villar-Corrales & Behnke, 2022; Locatello et al., 2020) to videos (Kipf et al., 2022; Singh et al., 2022; Elsayed et al., 2022), and real-world scenes (Seitzer et al., 2023; Aydemir et al., 2023; Zadaianchuk et al., 2024; Kakogeorgiou et al., 2024). The learned object representations benefit downstream tasks, such as reinforcement learning (Mosbach et al., 2024) or visual-question answering (Wu et al., 2023), among others.
2.2 Video Prediction and Generation
Future frame video prediction (VP) is the task of forecasting the upcoming $T$ video frames conditioned on the preceding $C$ seed frames (Oprea et al., 2020). When the number of seed frames is $C=1$, this task is often referred to as image-to-video generation. Several methods have been proposed to address this challenge, leveraging 2D convolutions (Gao et al., 2022; Chiu et al., 2020), 3D convolutions (Vondrick et al., 2016; Tulyakov et al., 2018), recurrent neural networks (RNNs) (Denton & Fergus, 2018; Villar-Corrales et al., 2022; Wang et al., 2022; Guen & Thome, 2020; Franceschi et al., 2020), Transformers (Rakhimov et al., 2021; Ye & Bilodeau, 2022, 2023), or diffusion models (Höppe et al., 2022; Ho et al., 2022).
2.2.1 Object-Centric Video Prediction
Object-centric VP presents a structured approach that explicitly models the dynamics and interactions of individual objects to forecast future video frames. These methods typically involve three main steps: decomposing seed frames into object representations, forecasting future object states using a dynamics model, and rendering video frames from the predicted object representations. Various approaches have been proposed for this task, using different architectural designs such as RNNs (Creswell et al., 2021; Zoran et al., 2021; Nguyen et al., 2024) or transformers (Wu et al., 2021, 2023; Villar-Corrales et al., 2023; Song et al., 2023; Daniel & Tamar, 2024). Despite promising results, these models are currently limited to simple deterministic datasets or rely on action-conditioning (Mosbach et al., 2024) or inferred latent vectors (Villar-Corrales & Behnke, 2025). In contrast, our proposed model forecasts future video frames conditioned on past object slots and textual descriptions.
2.2.2 Text-Conditioned Video Prediction
This category of VP models leverages text descriptions to provide appearance, motion and action cues that guide the generation of future frames. This task was first proposed by Hu et al. (2022), who utilized a VQ-VAE to encode the images into visual token representations, and modeled the image dynamics with an axial transformer that jointly processes such tokens along with text descriptions. TVP (Song et al., 2024) leverages RNNs to learn text representations, which condition a GAN-based framework in order to generate videos from a single frame. MMVG (Fu et al., 2023) combines a VQ-GAN (Esser et al., 2021) with a masked transformer predictor. More recently, several methods leverage diffusion models with text conditioning (Gu et al., 2024; Ni et al., 2023; Chen et al., 2023).
Concurrently with our work, Wang et al. (2024) combine object-centric learning with text-conditional image-to-video generation. The authors propose an autoregressive diffusion model conditioned on object slots and text descriptions and evaluate on simple datasets. In contrast, we explicitly model the object dynamics with an autoregressive transformer, and evaluate our approach on more complex robotics simulations.
3 Method
We propose TextOCVP – a novel object-centric model for text-conditioned image-to-video generation. Given an initial reference image $\textbf{X}_{1}$ and a text caption $\mathcal{C}$, TextOCVP generates the subsequent $T$ video frames $\hat{\mathbf{X}}_{2:{T+1}}$, which maintain a similar appearance and structural composition as the reference image, and follow the motion described in the text caption.
TextOCVP, which is illustrated in Fig. 2, implements an object-centric approach, in which the reference frame $\textbf{X}_{1}$ is first decomposed with a scene parsing module (Sec. 3.1) into a set of $N_{\mathbf{S}}$ $D$-dimensional object representations called slots $\textbf{S}_{1}\in{\mathbb{R}}^{N_{\mathbf{S}}\times D}$, where each slot represents a single object in the image. The object slots are fed to a text-conditioned transformer predictor (Sec. 3.2), which jointly models their spatio-temporal relations, and incorporates the textual information from the caption $\mathcal{C}$ as a guidance for predicting the future object slots $\hat{\mathbf{S}}_{2:T+1}$. Finally, the predicted object slots are decoded in order to render object representations and video frames (Sec. 3.3).
We propose two different TextOCVP variants, which differ in the underlying object-centric decomposition modules. Specifically, $\text{TextOCVP}_{\text{SAVi}}$ leverages SAVi (Kipf et al., 2022), whereas $\text{TextOCVP}_{\text{DINO}}$ extends the DINOSAUR (Seitzer et al., 2023) framework for recursive object-centric video decomposition and video rendering.
Figure 2: Overview of TextOCVP. TextOCVP parses the reference frame $\textbf{X}_{1}$ into object representations $\textbf{S}_{1}$. The text-conditioned object-centric predictor models object dynamics and interactions, incorporating information from the description $\mathbf{C}$ to predict future object states and frames.
3.1 Scene Parsing
The scene parsing module decomposes a video sequence $\mathbf{X}_{1:{\tau}}$ into a set of permutation-invariant object representations called slots $\mathbf{S}_{1:\tau}=(\textbf{S}_{1},...,\textbf{S}_{\tau})$, with $\textbf{S}_{t}=(\textbf{s}_{t}^{1},...,\textbf{s}_{t}^{N_{\mathbf{S}}})$, where each slot s represents a single object.
For learning the object representations, we leverage the object-centric video decomposition framework proposed by Kipf et al. (2022). At time step $t$, the corresponding input frame $\textbf{X}_{t}$ is encoded with a feature extractor module into a set of $D_{h}$-dimensional feature maps $\textbf{h}_{t}\in{\mathbb{R}}^{L\times D_{h}}$ representing $L$ spatial locations. The feature extractor is convolutional neural network in our $\text{TextOCVP}_{\text{SAVi}}$ variant and a DINO-pretrained vision transformer (Caron et al., 2021) in $\text{TextOCVP}_{\text{DINO}}$. These feature maps are processed with a Slot Attention (Locatello et al., 2020) corrector, which updates the previous slots $\textbf{S}_{t-1}$ based on visual features from the current frame following an iterative attention mechanism. Namely, Slot Attention performs cross-attention, with the attention weights normalized over the slot dimension, thus encouraging the slots to compete to represent parts of the input. It then updates the slots using a Gated Recurrent Unit (Cho et al., 2014) (GRU). Formally, Slot Attention updates the previous slots $\textbf{S}_{t-1}$ by:
$\textbf{A}=\operatorname*{softmax}_{N_{\mathbf{S}}}\left(\frac{q(\textbf{S}_{t-1})k(\textbf{h}_{t})^{T}}{\sqrt{D}}\right)\in{\mathbb{R}}^{N_{\mathbf{S}}\times L},$ (1)
$\textbf{S}_{t}=\text{GRU}(W_{t}v(\textbf{h}),\textbf{S}_{t-1})\;\text{with}\;W_{i,j}=\frac{\textbf{A}_{i,j}}{\sum_{l=1}^{L}\textbf{A}_{i,l}},$ (2)
where $k,q$ and $v$ are learned linear projections that map input features and slots into a common dimension. The final output of this module is the set of slots $\textbf{S}_{t}$, representing the objects of the input frame.
3.2 Text-Conditioned Object-Centric Predictor
Our proposed text-conditioned predictor module, depicted in Fig. 3, autoregressively forecasts future object states conditioned on the object slots from the reference frame $\textbf{S}_{1}$ and a textual description $\mathcal{C}$.
To condition the prediction process, the text description $\mathcal{C}$ is encoded into text token embeddings $\mathbf{C}$ employing an encoder-only transformer module. We experiment with different variants of this module, including a vanilla transformer encoder (Vaswani et al., 2017) or a pretrained T5 (Raffel et al., 2020) module.
At time step $t$, the predictor receives as input the corresponding text embeddings $\mathbf{C}$, as well as the previous object slots $\mathbf{S}_{1:t}$, which are initially mapped via an MLP to the predictor token dimensionality $D_{\text{Pred}}$. Additionally, these tokens are augmented with a temporal positional encoding, which applies the same sinusoidal positional embedding to all tokens from the same time-step, thus preserving the inherent permutation-equivariance of the objects.
Each layer of our predictor module mirrors the transformer decoder architecture (Vaswani et al., 2017). First, a self-attention layer enables every slot to attend to all other object representations in the sequence, modeling the spatio-temporal relations between objects. Subsequently, a text-to-slot cross-attention layer enhances the slot representations by incorporating important features from the text embeddings, such as motion or appearance information. Finally, an MLP is independently applied for each token. This process is repeated in every predictor layer, resulting in the predicted object slots of the subsequent time step $\hat{\textbf{S}}_{t+1}$. Furthermore, we apply a residual connection from $\textbf{S}_{t}$ to $\hat{\textbf{S}}_{t+1}$, which improves the temporal consistency of the predictions. This process is repeated autoregressively to obtain slot predictions for $T$ subsequent time steps.
3.3 Video Rendering
The video rendering module decodes the predicted slots $\hat{\textbf{S}}_{t}$ to render the corresponding video frame $\hat{\textbf{X}}_{t}$. We leverage two variants of the video rendering module, for our $\text{TextOCVP}_{\text{SAVi}}$ and $\text{TextOCVP}_{\text{DINO}}$ variants, respectively.
$\text{TextOCVP}_{\text{SAVi}}$ Decoder   This variant independently decodes each slot in $\hat{\textbf{S}}_{t}$ with a CNN-based Spatial Broadcast Decoder (Watters et al., 2019), rendering an object image $\textbf{o}_{t}^{n}$ and mask $\textbf{m}_{t}^{n}$ for each slot $\textbf{s}_{t}^{n}$. The object masks are normalized across the slot dimension, and the representations are combined via a weighted sum to render video frames:
$\hat{\textbf{X}}_{t}=\sum_{n=1}^{N_{\mathbf{S}}}\textbf{o}_{t}^{n}\cdot\tilde{\textbf{m}}_{t}^{n}\;\;\;\text{with}\;\;\;\tilde{\textbf{m}}_{t}^{n}=\operatorname*{softmax}_{N_{\mathbf{S}}}(\textbf{m}_{t}^{n}).$ (3)
Figure 3: Text-conditioned object-centric predictor.
$\text{TextOCVP}_{\text{DINO}}$ Decoder   This decoder variant decodes the object slots in two distinct stages. First, following DINOSAUR (Seitzer et al., 2023), an MLP-based Spatial Broadcast Decoder (Watters et al., 2019) is used to generate object features along with their corresponding masks. Similar to the $\text{TextOCVP}_{\text{SAVi}}$ decoder, the object masks are normalized and combined with the object features in order to reconstruct the encoded features $\hat{\textbf{h}}_{t}\in{\mathbb{R}}^{L\times D_{h}}$. In the second stage, the reconstructed features $\hat{\textbf{h}}_{t}$ are arranged into a grid format and processed with a CNN decoder to generate the corresponding video frame $\hat{\textbf{X}}_{t}$.
3.4 Training and Inference
Our proposed TextOCVP is trained in two different stages.
Object-Centric Learning   We first train the scene parsing and video rendering modules for decomposing video frames into object-centric representations by minimizing a reconstruction loss. In the $\text{TextOCVP}_{\text{SAVi}}$ variant, these modules are trained simply by reconstructing the input images, whereas in $\text{TextOCVP}_{\text{DINO}}$ they are trained by jointly minimizing an image and a feature reconstruction loss:
$\mathcal{L}_{\text{SAVi}}=\frac{1}{T}\sum_{t=1}^{T}||\hat{\textbf{X}}_{t}-\textbf{X}_{t}||,$ (4)
$\mathcal{L}_{\text{DINO}}=\frac{1}{T}\sum_{t=1}^{T}||\hat{\textbf{X}}_{t}-\textbf{X}_{t}||+||\hat{\textbf{h}}_{t}-\textbf{h}_{t}||.$ (5)
Predictor Training   Given the pretrained scene parsing and rendering modules, we train our TextOCVP predictor for text-conditioned video generation using a dataset containing paired videos and text descriptions. Namely, given the object representations from a reference frame $\textbf{S}_{1}$ and the textual description $\mathcal{C}$, our TextOCVP predictor forecasts the subsequent object slots $\hat{\textbf{S}}_{2}$, which are decoded into a predicted video frame $\hat{\textbf{X}}_{2}$. This process is repeated autoregressively, i.e. the predicted slots are appended to the input in the next time step, in order to generate the set of slots for the subsequent $T$ time steps. This autoregressive training, in contrast to teacher forcing, enforces our predictor to operate with imperfect inputs, leading to better modeling of long-term dynamics at inference time.
Our predictor is trained by minimizing the following combined loss:
$\mathcal{L}_{\text{TextOCVP}}=\frac{1}{T}\sum_{t=2}^{T+1}\lambda_{\text{Img}}\mathcal{L}_{\text{Img}}+\lambda_{\text{Slot}}\mathcal{L}_{\text{Slot}},$ (6)
with $\mathcal{L}_{\text{Img}}=||\hat{\textbf{X}}_{t}-\textbf{X}_{t}||_{2}^{2}\;\;\text{and}\;\;\mathcal{L}_{\text{Slot}}=||\hat{\textbf{S}}_{t}-\textbf{S}_{t}||_{2}^{2},$ (7)
where $\mathcal{L}_{\text{Img}}$ measures the future frame prediction error, and $\mathcal{L}_{\text{Slot}}$ enforces the alignment of the predicted object slots with the actual inferred object-centric representations.
Inference At inference time, TextOCVP receives as input a single reference frame and a language instruction. Our model parses the seed frame into object slots and autoregressively predicts future object states and video frames conditioned on the given textual description. By modifying the language instruction, TextOCVP can generate a new sequence continuation that performs the specified task while preserving a consistent scene composition.
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets
We evaluate the performance of TextOCVP for text-driven image-to-video generation on two different datasets, namely CATER and CLIPort. Further dataset details are provided in Appendix D.
CATER   CATER (Girdhar & Ramanan, 2020) is a dataset that consists of long video sequences, featuring two simultaneously moving 3D objects, with the motion described by a textual caption. We used the CATER-Hard variant introduced by Hu et al. (2022). This dataset contains 30,000 sequences featuring between three and eight objects, two of which follow a predefined action pattern. In our experiments, we resize the images to $64\times 64$.
CLIPort   CLIPort (Shridhar et al., 2022) is a robot manipulation dataset consisting of video-caption pairs. We employ 21,000 $336\times 336$ sequences of the Put-Block-In-Bowl task. Each sequence features a table with multiple bowls and blocks of different color, placed at random positions on the table. The corresponding caption describes the action of the robot arm picking a block of a specific color and putting it into a specific bowl.
4.1.2 Baselines
[…]
4.1.3 Implementation Details
All our models are implemented in PyTorch and trained on a single NVIDIA A6000 (48Gb) GPU. $\text{TextOCVP}_{\text{SAVi}}$ closely follows Kipf et al. (2022) for the design of the scene parsing and video rendering modules. $\text{TextOCVP}_{\text{DINO}}$ leverages DINOv2 (Oquab et al., 2023) as image encoder, uses a four-layer MLP-based Spatial-Broadcast Decoder (Watters et al., 2019), which is shared among all slots, to decode the object slots into object features and masks. Additionally, it uses a CNN decoder to map the reconstructed scene features back to images. On the CATER dataset, we use the $\text{TextOCVP}_{\text{SAVi}}$ variant with eight 128-dimensional object slots; whereas on CLIPort we employ the $\text{TextOCVP}_{\text{DINO}}$ variant with ten 128-dimensional slots. Our predictor module is an eight-layer transformer with 512-dimensional tokens, eight attention heads, and a hidden dimension of 1024. Further implementation details are provided in Appendix B.
4.2 Results
4.2.1 CATER Results
Table 1: Quantitative evaluation on CATER for prediction horizons of $T=9$ and $T=19$. TextOCVP outperforms all baselines. Best two results are highlighted in boldface and underlined, respectively.
CATER1→9 CATER1→19
[…]
OCVP 0.874 0.078 0.854 0.101
Non-OC TF 0.874 0.092 0.849 0.112
SEER 0.723 0.245 0.535 0.299
MAGE 0.877 0.108 0.871 0.111
TextOCVP 0.924 0.035 0.904 0.042
On the CATER dataset, we train the models to generate nine future frames given a single reference frame and a text caption. In Tab. 1 we report quantitative evaluations on CATER using the same setting as in training, i.e. predicting $T=9$ frames, as well as when predicting $T=19$ future frames. In both settings, TextOCVP outperforms all other models, demonstrating superior perceptual quality.
[…]
‘the medium green metal sphere is sliding
to (2, 1). the small brown metal cube
is picked up and placed to (-3, 1)’
Figure 4: Qualitative evaluation for text-driven image-to-video generation on CATER. Top row shows ground truth frames. TextOCVP generates sharp future frames, whereas MAGE blurs and misses objects.
Fig. 4 shows a qualitative comparison between TextOCVP and MAGE. Our proposed method generates a sequence that is closely aligned to the ground-truth, whereas MAGE predictions feature multiple errors and artifacts, including missing objects, blurry contours, and significant changes on object shapes. This highlights the effectiveness of object-centric representations to accurately represent and model object dynamics. Additional qualitative results are provided in Appendix E.
4.2.2 CLIPort Results
[…]
CLIPort1→9 CLIPort1→19
[…]
Non-OC TF 0.901 0.184 0.872 0.210
SEER 0.887 0.141 0.622 0.331
$\text{MAGE}_{\text{DINO}}$ 0.940 0.064 0.931 0.075
TextOCVP 0.950 0.062 0.931 0.078
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
‘put the gray block in the brown bowl
[…]
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
‘put the blue block in the gray bowl
[…]
Figure 5: Qualitative evaluation on CLIPort. Top rows depict the ground truth frames. TextOCVP successfully completes the pick-and-place task, whereas MAGE fails to predict robot motion (a) or the block disappears (b).
[…]
However, we identify that TextOCVP often generates artifacts in the background and lacks textured details in certain objects, thus degrading its quantitative performance despite generating accurate future frames. Additional qualitative results are provided in Appendix E.
4.3 Model Analysis
4.3.1 Ablation Studies
Table 3: Ablation studies. We ablate different architectural design choices in our proposed TextOCVP model.
(a) Effect of number of layers in predictor.
CATER1→9
[…]
2 31.83 0.912 0.041
4 32.58 0.921 0.036
8 32.82 0.924 0.035
(b) Evaluation of different text encoders on CATER and CLIPort.
CATER1→9 CLIPort1→9
[…]
Custom TF 31.62 0.909 0.045 - - -
Frozen T5 31.83 0.912 0.041 26.99 0.950 0.062
FT T5 - - - 26.85 0.947 0.065
(c) Effect of residual connection in predictor.
CLIPort1→9
[…]
No 26.55 0.946 0.066
Yes 26.99 0.950 0.062
(d) Effect of the number of slots $N_{\mathbf{S}}$.
CLIPort1→9
[…]
8 25.95 0.890 0.079
10 26.99 0.950 0.062
We perform several ablation studies to support and validate the architectural choices of our model components and their impact on TextOCVP’s image-to-video generation performance. The results are presented in Tab. 6(d).
Number of Layers (Tab. 6(a))   TextOCVP performs best with $N_{\text{Pred}}=8$ predictor layers. Scaling beyond eight layers was not explored, as the performance improvements when increasing from four to eight layers were marginal.
Residual Connection (Tab. 6(c))   Applying a residual connection to the predictor output, $\hat{\textbf{S}}_{t+1}:=\hat{\textbf{S}}_{t+1}+\textbf{S}_{t}$, enhances the model’s performance by improving the temporal consistency of the predicted slots.
Text Encoder (Tab. 6(b))   We evaluate the performance of different text encoders, including a vanilla transformer, a frozen T5 encoder, and a fine-tuned (FT) T5. A frozen T5 module led to the best TextOCVP performance on both CATER and CLIPort datasets. We argue that given the relatively small vocabulary size on CLIPort, fine-tuning T5 did not prove to be beneficial.
Number of Slots (Tab. 6(d))   Using $N_{\mathbf{S}}=10$ slots resulted in better performance for TextOCVP on CLIPort, even though each scene can be described with eight slots (six objects, robot arm and background). This observation suggests that additional slots can be used as registers in order to help with internal computations (Darcet et al., 2024).
Table 4: Quantitative evaluation on a CLIPort test set with unseen colors. We report the video generation performance and its drop relative to the known-color dataset.
CLIPort1→9 CLIPort1→19
[…]
$\text{MAGE}_{\text{DINO}}$ 0.935 -0.5% 0.076 -19% 0.924 -0.7% 0.087 -16%
TextOCVP 0.946-0.4% 0.066-6.4% 0.927-0.4% 0.083-6.4%
4.3.2 Model Robustness
We evaluate the performance of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set featuring color variations in the text instructions that were not encountered during training. This evaluation measures how well the models generalize to unseen scene features, highlighting their ability to handle novel configurations.
Tab. 4 presents the results on this evaluation set. We report both the video generation performance as well as the performance drop relative to the evaluation set with known colors. Our model outperforms $\text{MAGE}_{\text{DINO}}$ for both prediction horizons. Most notably, TextOCVP demonstrates significantly higher robustness to objects with previously unseen colors compared to the baseline, showing a much smaller drop in performance for the perceptual LPIPS metric. This highlights the effectiveness of leveraging object-centric representations for image-to-video generation and planning, in contrast to relying on holistic scene representations that struggle with novel scene compositions.
4.3.3 Controllability
‘the large purple rubber cone is picked up
and placed to (2, 3). the small gold metal
snitch is picked up and placed to (-1, 1).’
Original
Caption
a)
‘the medium cyan rubber sphere is picked up
and placed to (-2, 2). the medium purple
metal cone is sliding to (2, 3).’
Changed
Objects
& Actions
b)
‘the large purple rubber cone is sliding to
(-1, 1). the small gold metal snitch is
rotating. the medium purple metal cone
is picked up and placed to (-1, -3).’
More
Moving
Objects
c)
[…]
Figure 7: Qualitative evaluation of TextOCVP controllability on CATER. Top row shows the ground truth sequence. We underline the changed actions with respect to the original caption. We demonstrate TextOCVP’s controllability by generating multiple sequences from the same reference frame, each conditioned on a different text instruction.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the cyan block in the brown bowl
Original
Caption
Refer to captionRefer to captionRefer to caption‘put the blue block in the brown bowl
Changed
Block
[…]
Figure 8: Qualitative evaluation of TextOCVP controllability on CLIPort. Top row depicts ground-truth frames. TextOCVP correctly generates a sequence where the robot picks up the correct block and places it into the bowl as specified in the textual instruction.
A key objective of text-driven image-to-video generation is to provide control over the model generations. This is achieved through language instructions that describe the objects in the scene and their expected motion. We qualitatively assess the controllability of our model on both the CATER and CLIPort datasets.
CATER   In Fig. 7, we showcase the control that TextOCVP provides over its generations. We qualitatively evaluate this ability by generating multiple videos conditioned on the same reference frame while varying the language instruction. We experiment with altering both the moving and target objects and their actions in the language instruction, as well as providing instructions that include a number of tasks beyond what the model was trained on.
As shown in Fig. 7, TextOCVP successfully identifies the objects described in the text and executes the instructions accordingly. Notably, our model can distinguish between two nearly identical purple cones in the scene, despite their identical shapes and color, and correctly generates a sequence with the specified motion.
This demonstrates that key object attributes, such as size or color, are effectively captured through the text-to-slot attention mechanism. Leveraging this information, the text-conditioned predictor accurately predicts the future motion of each individual object as specified in the instruction. These results highlight the benefits of combining object-centric representations with text-based conditioning.
CLIPort   We perform a similar experiment on the CLIPort dataset, as shown in Fig. 8. A single frame containing multiple colored blocks and bowls is provided to TextOCVP. By modifying the textual instruction, we can control which block the robot arm picks up and the bowl where it is placed. In both examples, TextOCVP successfully selects the correct block and places it in the specified bowl. Furthermore, the movement of the robot arm adapts to the motion described in the instruction, demonstrating accurate and responsive control.
5 Conclusion
We presented TextOCVP, a novel object-centric model for text-driven image-to-video generation. Given a single input image and a text description, TextOCVP generates a sequence that matches the description by parsing the environment into object representations and modeling their dynamics conditioned on the textual instruction. This is achieved with a novel text-conditioned object-centric transformer, which predicts future object states by modeling the spatio-temporal relationships between objects while incorporating the guidance from text. Through extensive evaluations, we demonstrated that TextOCVP outperforms existing text-conditioned models for the image-to-video generation, highlighting our model’s ability to generate consistent long sequences and adapt its predictions based on language input. Moreover, we validated our architectural choices through ablation studies, highlighting the importance of combining textual and object-centric information. With its structured latent space and superior controllability, TextOCVP offers potential for robotics applications, where it can serve as a controllable world model, enabling efficient planning, reasoning, and decision-making.
Acknowledgment
This work was funded by grant BE 2556/16-2 (Research Unit FOR 2535 Anticipating Human Behavior) of the German Research Foundation (DFG)
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Limitations and Future Work
While TextOCVP demonstrates promising results for text-guided object-centric video prediction, it presents some limitations that we plan to address in future work:
Prediction Artifacts   TextOCVP occasionally generates artifacts in the predicted frames, such as blurriness, inconsistent object appearances, lack of textured details in the objects, or visual artifacts in the background. We believe that these limitations stem from the video rendering module, which might lack the representational power to reconstruct precise image details from the latent space representation.
Poor Temporal Consistency   We observe that TextOCVP’s predictions often lack temporal consistency when forecasting for long prediction horizons ($T>30$). We attribute this limitation to the fact that TextOCVP is trained by optimizing reconstruction losses, which do not penalize temporal inconsistencies.
Future Work   To address these limitations, we plan to extend our TextOCVP framework with more powerful decoder modules, such as autoregressive transformers (Singh et al., 2022) or diffusion models (Jiang et al., 2023), as well as scale our predictor module. Furthermore, we plan to incorporate temporal discriminators (Clark et al., 2020) to improve the temporal consistency of the predicted video frames. We believe that exploring these architectural modifications will enable us to utilize TextOCVP as a world model in complex real-world robotic environments.
Appendix B Implementation Details
We employ $\text{TextOCVP}_{\text{SAVi}}$ for the experiments on CATER and $\text{TextOCVP}_{\text{DINO}}$ for experiments on CLIPort. Below we discuss the implementation details for each of these variants.
B.1 $\text{TextOCVP}_{\text{DINO}}$
$\text{TextOCVP}_{\text{DINO}}$ variant consists of our proposed text-conditioned predictor module and an object-centric decomposition module that extends the DINOSAUR (Seitzer et al., 2023) framework for recursive object-centric video decomposition and video rendering.
Text-Conditioned Predictor   The predictor is composed of $N_{\text{Pred}}=8$ identical layers, each containing 8-head attention mechanisms and an MLP with a single hidden layer of dimension 1024 and a ReLU activation function. Furthermore, the predictor uses an embedding dimensionality of 512, context window size of ten frames, and applies a residual connection from the predictor input to its output.
Text Encoder   $\text{TextOCVP}_{\text{DINO}}$ leverages a pretrained and frozen small version of T5 encoder (Raffel et al., 2020), which consists of 6 T5 blocks. This text encoder uses a vocabulary with size 32,128.
Scene Parsing   The scene parsing module generates $N_{\mathbf{S}}=10$ slots of dimension 128. As feature extractor, we use DINOv2 ViT-Base (Oquab et al., 2023), featuring 12 layers, using a patch size of 14, and producing patch features with dimension $D_{h}=768$. The Slot Attention corrector module processes the first video frame with three iterations in order to obtain a good initial object-centric decomposition, and a single iteration for subsequent frames, which suffices to recursively update the slot representation. The initial object slots $\textbf{S}_{0}$ are randomly sampled from a Gaussian distribution with learned mean and covariance. We use a single Transformer encoder block as the transition function, which consists of a four attention heads and an MLP with hidden dimension 512.
Video Rendering   The video rendering module consists of two distinct decoders. First, a four-layer MLP-based Spatial Broadcast Decoder (Watters et al., 2019) with hidden dimension 1024 reconstructs the patch features from the slots. Then, to reconstruct the images from the features, we implement a CNN-based decoder with four convolutional layers, where each layer uses $3\times 3$ kernels. Every layer is followed by bilinear upsampling and a ReLU activation function. A final convolutional layer is applied to map to the RGB channels of the image.
Training   We train our model for object-centric decomposition using videos sequences of length five frames for 1000 epochs. We use batch size of 16, the Adam optimizer (Kingma & Ba, 2015), and a base learning rate of $4\times 10^{-4}$, which is linearly warmed-up for the first 10000 steps, followed by cosine annealing for the remaining of the training process. Moreover, we clip the gradients to a maximum norm of 0.05. The predictor module is trained given a frozen and pretrained object-centric decomposition model for 700 epochs. The predictor is trained using the same hyper-parameters as for object-centric decomposition. In the predictor loss function $\mathcal{L}_{\text{TextOCVP}}$, we set $\lambda_{\text{Img}}=1$ and $\lambda_{\text{Slot}}=1$.
B.2 $\text{TextOCVP}_{\text{SAVi}}$
$\text{TextOCVP}_{\text{SAVi}}$ uses the same text-conditioned predictor and text-encoder architectures as $\text{TextOCVP}_{\text{DINO}}$, but employs SAVi (Kipf et al., 2022) as the object-centric decomposition module.
Scene Parsing   The scene parsing module generates $N_{\mathbf{S}}=8$ slots of dimension 128. Following Kipf et al. (2022), we use as feature extractor a four-layer CNN with ReLU activation function, where each convolutional layer features 32 $5\times 5$ kernels, $\text{stride}=1$, and $\text{padding}=2$. The Slot Attention corrector follows the same structure as in $\text{TextOCVP}_{\text{DINO}}$.
Video Rendering   Following Kipf et al. (2022), we utilize a CNN-based Spatial Broadcast Decoder (Watters et al., 2019) with four convolutional layers with 32 kernels of size $5\times 5$ , $\text{stride}=1$, and $\text{padding}=2$, and a final convolutional layer which maps to four channels (RGB + alpha mask).
Training   We train our model for object-centric decomposition using video sequences of length ten frames for 1000 epochs, using batch size of 64, and an initial learning rate of $10^{-4}$, which is warmed up for 2500 steps, followed by cosine annealing for the remaining of the training process. Moreover, we clip the gradients to a maximum norm of 0.05. The predictor module is trained given a frozen and pretrained object-centric decomposition model for 1400 epochs. The predictor is trained using the same hyper-parameters as for object-centric decomposition. In the predictor loss function $\mathcal{L}_{\text{TextOCVP}}$, we set $\lambda_{\text{Img}}=1$ and $\lambda_{\text{Slot}}=1$.
Appendix C Baselines
We employ four different baselines to compare againts our TextOCVP model on the image-to-video generation task on CATER and CLIPort datasets. To emphasize the importance of incorporating textual information, we include a comparison with OCVP-Seq (Villar-Corrales et al., 2023), a recent object-centric video prediction model that does not utilize text conditioning. Additionally, we evaluate a non-object-centric TextOCVP variant (Non-OC TF) that processes the input image into a single high-dimensional slot representation, instead of multiple object-centric slots, thus allowing us to evaluate the effect of object-centric representations. Moreover, we compare TextOCVP with two popular text-conditioned image-to-video generation baselines that do not incorporate object-centricity: MAGE (Hu et al., 2022) and SEER (Gu et al., 2024). We train both models on CATER and CLIPort closely following the original implementation details [^1] https://github.com/Youncy-Hu/MAGE [^2] https://github.com/seervideodiffusion/SeerVideoLDM/tree/main.
C.1 MAGE
MAGE is an autoregressive text-guided image-to-video generation framework that utilizes a VQ-VAE (Russakovsky et al., 2015) encoder-decoder architecture to learn efficient visual token representations. A cross-attention module aligns textual and visual embeddings to produce a spatially-aligned motion representation termed Motion Anchor (MA), which is fused with visual tokens via an axial transformer for video generation. For experiments on CATER, we use a codebook size of $512\times 256$ with a downsampling ratio of four, whereas on CLIPort we use a codebook size of $512\times 1024$. Moreover, on CLIPort we replace MAGE’s standard CNN encoder and decoder with the DINOv2 ViT encoder and CNN decoder used in our $\text{TextOCVP}_{\text{DINO}}$ approach. This adjustment ensures a fair comparison and significantly enhances MAGE’s performance on CLIPort. We refer to this modified version as $\text{MAGE}_{\text{DINO}}$. Additionally, to align with our experimental setup, we omit the speed parameter.
C.2 SEER
SEER is a diffusion-based model for language-guided video prediction. It employs an Inflated 3D U-Net derived from a pretrained text-to-image 2D latent diffusion model (Rombach et al., 2022), extending it along the temporal axis and integrating temporal attention layers to simultaneously model spatial and temporal dynamics. For the language conditioning module, SEER introduces a novel Frame Sequential Text (FSText) Decomposer, which decomposes global instructions generated by the CLIP text encoder (Radford et al., 2021) into frame-specific sub-instructions. These are aligned with frames using a transformer-based temporal network and injected into the diffusion process via cross-attention layers. We initialize SEER from a checkpoint pretrained on the Something-Something V2 dataset (Goyal et al., 2017), and further fine-tune it for a few epochs. We observed that incorporating a text loss enhanced SEER’s performance, while other hyper parameters were kept consistent with its original implementation.
Appendix D Datasets
CATER   CATER (Girdhar & Ramanan, 2020) is a dataset that consists of long video sequences, each described by a textual caption. The video scenes consist of multiple 3D geometric objects in a 2D table plane, which is split into a $6\times 6$ grid with fixed axis, allowing the exact description of object’s positions using coordinates. The text instruction describes the movement of specific objects through four atomic actions: ‘rotate’, ‘pick-place’, ‘slide’, and ‘contain’. The caption follows a template consisting of the subject, action, and an optional object or end-point coordinate, depending on the action. The movement of the objects starts at the same time step. Furthermore, the initial positions are randomly selected from the plane grid, and the camera position is fixed for every sequence.
In our work, we employ CATER-hard, which is a complete version of the CATER dataset, containing 30000 video-caption pairs. It includes 5 possible objects: cone, cube, sphere, cylinder, or snitch, which is a special small object in metallic gold color, shaped like three intertwined toruses. Furthermore, every object is described by its size (small, medium, or large), material (metal or rubber), and color (red, blue, green, yellow, gray, brown, purple, cyan, or gold if the object is the snitch), and this description is included in the textual caption. Every atomic action is available. The ‘rotate’ action is afforded by cubes, cylinders and the snitch, the ‘contain’ action is only afforded by the cones, while the other two actions are afforded by every object. Every video has between 3 and 8 objects, and two actions happen to different objects at the same time. The vocabulary size is 50.
CLIPort   CLIPort (Shridhar et al., 2022) is a robot manipulation dataset, consisting of video-caption pairs, i.e. long videos whose motion is described by a textual video caption. There are many variants of the CLIPort dataset, but we focus on the Put-Block-In-Bowl variant. We generate 21000 video-caption pairs with resolution $336\times 336$. Every video contains 6 objects on a 2D table plane, and a robot arm. Objects can be either a block or a bowl, and there is at least one of them in every sequence. The starting position of each object is random, with the only constrain to be placed on the table. Each video describes the action of the robot arm picking a block, and putting it in a specific bowl. The video caption follows the template ‘put the [color] block in the [color] bowl’. Each individual object in the scene has a different color. In the train and validation set, the block and the bowl that are part of the caption can have one of the following colors: blue, green, red, brown, cyan, gray, or yellow, while in the test set they can have blue, green, red, pink, purple, white, or orange color. The other 4 objects, called distractors, can have any color. During a video sequence, it can be possible that the robot arm goes out of frame, and comes back in later frames, thus requiring the model to leverage long range dependencies. The vocabulary size is 15.
Appendix E Additional Results
[…]
Ext. DINO
Masks
Figure 9: Comparison between SAVi and our Extended DINOSAUR (Ext. DINO) decomposition modules in reconstructing an CLIPort video sequence.
E.1 SAVi vs. DINO
Current object-centric approaches for video generation are limited to relatively simple synthetic datasets. We attribute this limitation primarily to the object-centric modules used for learning object representations. Motivated by this observation, we extend DINOSAUR (Seitzer et al., 2023) to handle video data and reconstruct images effectively.
To demonstrate the significance of the object-centric module in scaling to more complex datasets, we input the same video sequence to both SAVi and our extended DINOSAUR trained on CLIPort. As illustrated in Fig. 9, SAVi struggles to accurately model objects on the 2D plane, whereas our proposed object-centric model successfully reconstructs the scene, closely resembling the input. The visual features extracted by the DINOv2 (Oquab et al., 2023) encoder contain high-level semantic information, and during training, the slots are specifically optimized to efficiently encode this information. This design enables the model to scale and handle more complex object-centric video data effectively.
E.2 Robustness to Number of Objects
We evaluate the performance of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set consisting of scenes with a larger number of objects than encountered during training, i.e. the 2D table plane contains 8 instead of 6 objects. This evaluation demonstrates the robustness of models when generalizing to scenes with more objects. The quantitative results are presented in Tab. 5. We report both the video generation performance and the performance drop relative to the original evaluation set.
We observe that our model outperforms $\text{MAGE}_{\text{DINO}}$ on every metric, for both prediction horizons. Notably, TextOCVP demonstrates significantly higher robustness to scenes with more objects, as reflected by the smaller decline in performance for the perceptual LPIPS metric. This result is further illustrated in Fig. 10, where we show a qualitative comparison of video generations for scenes with 8 objects. As observed, our model correctly illustrates the motion described in the text instructions, while $\text{MAGE}_{\text{DINO}}$ fails to generate accurate sequences, missing the target block, a behaviour observed in various examples.
These results highlight once again the effectiveness of object-centric representations in image-to-video generation, as TextOCVP is able to generalize to scenes with more objects by simply increasing the number of slots.
Table 5: Quantitative evaluation on a CLIPort test set with more objects. We report the video generation performance and its drop relative to the original dataset.
CLIPort1→9 CLIPort1→19
[…]
$\text{MAGE}_{\text{DINO}}$ 0.929 (-1.2%) 0.088 (-37.5%) 0.920 (-1.2%) 0.094 (-25.3%)
TextOCVP 0.936 (-1.5%) 0.076 (-22.6%) 0.921 (-1.1%) 0.090 (-15.4%)
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the gray block in the brown bowl Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
$\text{MAGE}_{\text{DINO}}$
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
TextOCVP
[…]
Figure 10: Qualitative evaluation of $\text{MAGE}_{\text{DINO}}$ and TextOCVP on a CLIPort sequence with more objects than those seen during training. TextOCVP correctly generates a sequence where the robot picks up the gray block and places it in the brown bowl, while $\text{MAGE}_{\text{DINO}}$ fails by missing the target bowl.
E.3 Text-to-Slot Attention Visualizations
An additional advantage of using object-centric representations is the improved interpretability. This can be shown in the visualizations of text-to slot attention weights, which help us understand how the textual information influences and guides the model predictions. First, in Fig. 11, we visualize the text-to-slot attention weights for different cross-attention heads for a single slot that represents the rotating red cube. We observe that the slot attends to relevant text tokens from the input, such as the object shape, size and the action taking place.
‘the small red rubber cube is rotating.
the small gray cylinder is rotating.’
Refer to captionSceneRefer to captionAttended ObjectRefer to captionRefer to caption1.00.50.0
Head 1
the
small
red
rubber
cu
be
is
rotating
the
small
gray
metal
cylinder
is
rotating
Refer to captionRefer to caption1.00.50.0
Head 2
the
small
red
rubber
cu
be
is
rotating
the
small
gray
metal
cylinder
is
rotating
Refer to captionRefer to caption1.00.50.0
Head 3
the
small
red
rubber
cu
be
is
rotating
the
small
gray
metal
cylinder
is
rotating
Refer to captionRefer to caption
1.00.50.0
Head 4
the
small
red
rubber
cu
be
is
rotating
the
small
gray
metal
cylinder
is
rotating
Figure 11: Visualization of the text-to-slot attention between an object and the textual caption. We visualize the text-to-slot attention weights for different heads. Our predictor module focuses on relevant text tokens to predict the scene dynamics, including the object size, shape or the action.
In Fig. 12, we additionally visualize the text-to-slot attention weights, averaged across attention heads, for different slots in a CATER sequence. We observe that slots that represent objects in the textual description attend to relevant text tokens, such as their target coordinate locations. These results demonstrate that the text-to-slot attention mechanism effectively aligns the textual information with the object-centric representations, enabling the model to generate accurate video sequences based on the given text instructions, while providing superior interpretability.
[…]
‘the cone is sliding to (-3, -3).
the snitch is picked up and placed to (-1, 2 ).’
Refer to captionRefer to caption1.00.50.0Refer to caption
Slot 3
the
cone
is
sliding
to
-3
-3
the
snitch
is
picked
up
and
placed
to
-1
2
Refer to captionRefer to caption1.00.50.0Refer to caption
Slot 4
the
cone
is
sliding
to
-3
-3
the
snitch
is
picked
up
and
placed
to
-1
2
Refer to captionRefer to caption1.00.50.0Refer to caption
Slot 6
the
cone
is
sliding
to
-3
-3
the
snitch
is
picked
up
and
placed
to
-1
2
Figure 12: Text-to-slot attention weights, averaged across attention heads, for different objects in the sequence. Slots that represent objects in the textual description attend to relevant text tokens, such as their target locations.
E.4 Qualitative Evaluations
Figs. 13 and 14  show qualitative evaluations on CATER in which both MAGE and TextOCVP successfully generate a sequence following the instructions from the textual description. Fig. 15 illustrates an example where MAGE fails to generate a correct sequence, while TextOCVP successfully completes the task described by the text.
Figs. 16 and 17  show examples of TextOCVP’s control over the predictions. In both sequences, TextOCVP generates a correct sequence given the text instructions, and seamlessly adapts its generations to a modified version of the textual instructions.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
‘the large brown metal cone is picked up and containing the
medium red rubber cone. the small gold metal snitch is rotating.’
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
MAGE
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
TextOCVP
[…]
Figure 13: Qualitative evaluation on CATER. Both MAGE and TextOCVP successfully generate a sequence following the instructions from the textual description.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
‘the medium green rubber cone is picked up and
containing the small gold metal snitch.
the large purple rubber cone is picked up and placed to (-1, 3).’
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
MAGE
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
TextOCVP
[…]
Figure 14: Qualitative evaluation on CATER. Both MAGE and TextOCVP successfully generate a sequence that illustrates the motion described in the text, but MAGE’s predictions are of a lower resolution.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
‘the large yellow rubber cone is sliding to (2, 3).
the small gold metal snitch is picked up and placed to (-3, 1).’
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
MAGE
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
TextOCVP
[…]
Figure 15: Qualitative evaluation on CATER. MAGE fails to generate a sequence that accurately follows the motion described in the text. Specifically, the yellow cone does not slide as expected, and artifacts such as the merging of two small objects are introduced. On the other hand, the sequence generated by TextOCVP is closely aligned with the ground truth, accurately capturing the motion of the objects.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the blue block in the brown bowl
Original
Caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the blue block in the red bowl
Changed
Bowl
[…]
Figure 16: Qualitative evaluation of TextOCVP controllability on CLIPort. TextOCVP correctly generates a sequence where the robot picks up and places the block specified in the textual instruction.
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the green block in the cyan bowl
Original
Caption
Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the green block in the red bowl
Changed
Bowl
[…]
Figure 17: Qualitative evaluation of TextOCVP controllability on CLIPort. TextOCVP correctly generates a sequence where the robot picks up and places the block specified in the textual instruction.

### v2

@@
-Object-Centric Image to Video Generation with Language Guidance
+TextOCVP: Object-Centric Video Prediction with Language Guidance
 Abstract
-Accurate and flexible world models are crucial for autonomous systems to understand their environment and predict future events.
-Object-centric models, with structured latent spaces, have shown promise in modeling object dynamics and interactions, but often face challenges in scaling to complex datasets and incorporating external guidance, limiting their applicability in robotics.
-To address these limitations, we propose TextOCVP, an object-centric model for image-to-video generation guided by textual descriptions.
+Understanding and forecasting future scene states is critical for autonomous agents to plan and act effectively in complex environments.
+Object-centric models, with structured latent spaces, have shown promise in modeling object dynamics and predicting future scene states, but often struggle to scale beyond simple synthetic datasets and to integrate external guidance, limiting their applicability in robotics.
+To address these limitations, we propose TextOCVP, an object-centric model for video prediction guided by textual descriptions.
 TextOCVP parses an observed scene into object representations, called slots, and utilizes a text-conditioned transformer predictor to forecast future object states and video frames.
-Our approach jointly models object dynamics and interactions while incorporating textual guidance, thus leading to accurate and controllable predictions.
-Our method’s structured latent space offers enhanced control over the prediction process, outperforming several image-to-video generative baselines.
-Additionally, we demonstrate that structured object-centric representations provide superior controllability and interpretability, facilitating the modeling of object dynamics and enabling more precise and understandable predictions.
-Videos and code are available at https://play-slot.github.io/TextOCVP/.
-Object-Centric Learning, Text-Conditioned Video Prediction, Object-Centric Video Prediction, Representation Learning, Transformers
-Refer to captiona) Object-Centric DecompositionLearned ObjectsRefer to captionb) Text-Conditioned Image-to-Video Generation
+Our approach jointly models object dynamics and interactions while incorporating textual guidance, enabling accurate and controllable predictions.
+TextOCVP’s structured latent space offers a more precise control of the forecasting process, outperforming several video prediction baselines on two datasets.
+Additionally, we show that structured object-centric representations provide superior robustness to novel scene configurations, as well as improved controllability and interpretability, enabling more precise and understandable predictions.
+Videos and code are available at https://play-slot.github.io/TextOCVP.
+1 Introduction
+Understanding and reasoning about the environment is essential for enabling autonomous systems to better comprehend their surroundings, predict future events, and adapt their actions accordingly.
+Humans achieve these capabilities by perceiving the environment as a structured composition of individual objects that interact and evolve dynamically over time (Kahneman, Treisman, and Gibbs 1992).
+Neural networks equipped with such compositional inductive biases have shown the ability to learn structured object-centric representations of the world, which enable desirable properties, such as out-of-distribution generalization (Dittadi et al.
+2022), compositionality (Greff, Van Steenkiste, and Schmidhuber 2020), or sample efficiency (Mosbach et al.
+2025).
+Recent advances in unsupervised object-centric representation learning have progressed from extracting object representations in synthetic images (Locatello et al.
+2020; Lin et al.
+2020) to modeling objects in video (Kipf et al.
+2022; Singh, Wu, and Ahn 2022) and scaling to real-world scenes (Seitzer et al.
+2023; Zadaianchuk, Seitzer, and Martius 2024).
+These developments have enabled object-level dynamics modeling for future prediction and planning.
+Notably, approaches like SlotFormer (Wu et al.
+2023) or OCVP (Villar-Corrales, Wahdan, and Behnke 2023) introduced object-centric prediction models that explicitly model spatio-temporal relationships between objects, shifting away from image-level approaches that ignore scene compositionality.
+Despite these advancements, current object-centric methods struggle with complex object appearances and dynamics, and lack mechanisms to incorporate external guidance, thus limiting their scalability and broader applicability.
+Refer to captiona) Object-Centric ParsingLearned ObjectsRefer to captionb) Text-Conditioned Video Prediction
 ‘Put the green block
 in the cyan bowl.’
-$\mathcal{C}$
+‘Put the green block
+in the red bowl.’
 Figure 1: Overview of TextOCVP.
-(a) Our model first parses a reference frame $\textbf{X}_{1}$ into its object components $\textbf{S}_{1}$.
-(b) Our TextOCVP predictor jointly models object dynamics and interactions while integrating textual guidance, generating future object states and video frames that align with the provided textual instructions.
-1 Introduction
-Understanding and reasoning about the real world is essential for enabling autonomous systems to better comprehend their surroundings, predict future events, and adapt their actions accordingly.
-Humans achieve these capabilities by perceiving the environment as a structured composition of individual objects that interact and evolve dynamically over time (Kahneman et al., 1992).
-Neural networks equipped with such compositional inductive biases have shown the ability to learn structured object-centric representations of the world, which enable desirable properties, such as out-of-distribution generalization (Dittadi et al., 2022), compositionality (Greff et al., 2020), or sample efficiency (Mosbach et al., 2024).
-Recent advances in unsupervised object-centric representation learning have progressed from extracting object representations in synthetic images (Locatello et al., 2020; Lin et al., 2020b) to modeling objects in video (Kipf et al., 2022; Singh et al., 2022) and scaling to real-world scenes (Seitzer et al., 2023; Zadaianchuk et al., 2024).
-These developments have enabled object-level dynamics modeling for future prediction and planning.
-Notably, models like SlotFormer (Wu et al., 2023) or OCVP (Villar-Corrales et al., 2023) introduced object-centric world models that explicitly model spatio-temporal relationships between objects, shifting away from image-level approaches that ignore scene compositionality.
-Despite these advancements, current approaches struggle with complex object appearances and dynamics, and lack the ability to incorporate external guidance, thus limiting their effectiveness as world models in robotic applications.
-To address these challenges, we propose TextOCVP, a novel object-centric model for image-to-video generation with language guidance, illustrated in Fig.
+(a) Our model parses a reference frame $\textbf{X}_{1}$ into its object components $\textbf{S}_{1}$.
+(b) Our TextOCVP predictor jointly models object dynamics and interactions guided by text, generating future object states and frames that align with the provided textual instructions.
+To address these challenges, we propose TextOCVP, a novel object-centric model for video prediction guided by textual instructions, illustrated in Fig.
 1.
-Given a reference image and text instruction, TextOCVP extracts object representations and predicts their evolution using a novel text-conditioned object-centric transformer.
-This module forecasts future object representations by modeling object dynamics and interactions, while also incorporating textual information through a text-to-slot attention mechanism.
-By jointly modeling spatio-temporal object relationships and integrating textual guidance, TextOCVP generates future object states and video frames that align with the provided instructions.
-We evaluate our approach through extensive experiments focusing on image-to-video generation tasks with varying levels of complexity.
-Our results show that TextOCVP outperforms existing text-conditioned methods by effectively leveraging object-centric representations.
-This demonstrates the significant advantage of incorporating textual guidance into structured object-centric representations, particularly for scenes featuring multiple moving objects.
-Through an in-depth model analysis, we demonstrate the importance of our proposed object-centric latent space, which allows TextOCVP to generate video sequences that closely align with language instructions.
-We further demonstrate how TextOCVP adapts its generated video sequences based on different captions by routing the text information to the relevant object representations.
+Given a reference image and text instruction, TextOCVP extracts object representations and predicts their evolution using a text-conditioned object-centric transformer.
+This predictor forecasts future object states by explicitly modeling their dynamics and interactions over time, while integrating textual information via a text-to-slot attention mechanism.
+By jointly modeling spatio-temporal object relationships and incorporating textual guidance, TextOCVP predicts future object states and frames aligned with the input instruction.
+We evaluate our approach for text-conditioned video prediction through extensive experiments on two distinct datasets.
+Our results show that TextOCVP outperforms existing text-conditioned video prediction methods by effectively leveraging structured object-centric representations, particularly on scenes featuring multiple moving objects.
+To verify the effectiveness of object-centric representations for text-guided video prediction, we conduct an in-depth analysis of TextOCVP.
+Specifically, we demonstrate that its structured latent space enables accurate and controllable video prediction by aligning language instructions with the corresponding objects, outperforming baselines that rely on holistic scene representations.
+Beyond generation quality, TextOCVP offers improved interpretability and exhibits strong robustness to novel scene configurations, including varying numbers of objects or previously unseen colors.
 In summary, our contributions are as follows:
-  • We propose TextOCVP, a novel text-driven image-to-video generation model, featuring a text-conditioned object-centric predictor that integrates textual guidance via a text-to-slot attention mechanism.
-  • Through extensive evaluation, we show that TextOCVP outperforms existing text-conditioned models by leveraging object-centric representations.
-  • We demonstrate that TextOCVP is controllable, seamlessly adapting to diverse textual instructions.
+  • We propose TextOCVP, a text-guided video prediction model, featuring a text-conditioned object-centric predictor that integrates textual guidance into the prediction process via a text-to-slot attention mechanism.
+  • Through extensive evaluations, we show that TextOCVP outperforms existing text-conditioned video prediction models by leveraging object-centric representations.
+  • We demonstrate that TextOCVP is controllable, seamlessly adapting to diverse textual instructions, while exhibiting interpretability and robustness to novel setups.
 2 Related Work
 2.1 Object-Centric Learning
-Representation learning, the ability to extract meaningful features from data, often improves model performance by enhancing its understanding of the input space (Bengio et al., 2013).
-Object-centric representation learning focuses on simultaneously learning representations that characterize individual objects within an input image or video.
-Recently, object-centric models have progressed from learning object representations from synthetic images (Greff et al., 2019; Lin et al., 2020a; Villar-Corrales & Behnke, 2022; Locatello et al., 2020) to videos (Kipf et al., 2022; Singh et al., 2022; Elsayed et al., 2022), and real-world scenes (Seitzer et al., 2023; Aydemir et al., 2023; Zadaianchuk et al., 2024; Kakogeorgiou et al., 2024).
-The learned object representations benefit downstream tasks, such as reinforcement learning (Mosbach et al., 2024) or visual-question answering (Wu et al., 2023), among others.
-2.2 Video Prediction and Generation
-Future frame video prediction (VP) is the task of forecasting the upcoming $T$ video frames conditioned on the preceding $C$ seed frames (Oprea et al., 2020).
-When the number of seed frames is $C=1$, this task is often referred to as image-to-video generation.
-Several methods have been proposed to address this challenge, leveraging 2D convolutions (Gao et al., 2022; Chiu et al., 2020), 3D convolutions (Vondrick et al., 2016; Tulyakov et al., 2018), recurrent neural networks (RNNs) (Denton & Fergus, 2018; Villar-Corrales et al., 2022; Wang et al., 2022; Guen & Thome, 2020; Franceschi et al., 2020), Transformers (Rakhimov et al., 2021; Ye & Bilodeau, 2022, 2023), or diffusion models (Höppe et al., 2022; Ho et al., 2022).
-2.2.1 Object-Centric Video Prediction
+Representation learning, the ability to extract meaningful features from data, often improves model performance by enhancing its understanding of the input space (Bengio, Courville, and Vincent 2013).
+Object-centric representation methods aim to parse an image or video into a set of object components in an unsupervised manner.
+These objects are typically represented as unconstrained embeddings (called slots) (Locatello et al.
+2020; Kipf et al.
+2022; Singh, Wu, and Ahn 2022), patch-based representations (Lin et al.
+2020), factored latent vectors (Greff et al.
+2019), or explicit object prototypes (Villar-Corrales and Behnke 2022; Monnier, Groueix, and Aubry 2020).
+These methods have demonstrated promise in learning object representations across diverse domains, ranging from synthetic images (Locatello et al.
+2020; Lin et al.
+2020) to videos (Kipf et al.
+2022; Singh, Wu, and Ahn 2022; Elsayed et al.
+2022), and real-world scenes (Seitzer et al.
+2023; Zadaianchuk, Seitzer, and Martius 2024).
+The learned object representations benefit downstream tasks, such as reinforcement learning (Mosbach et al.
+2025) or visual-question answering (Wu et al.
+2023).
+2.2 Video Prediction
+Video prediction (VP) is the task of forecasting the upcoming $T$ video frames conditioned on the preceding $C$ seed frames (Oprea et al.
+2020).
+Several methods have been proposed to address this challenge, leveraging 2D convolutions (Gao et al.
+2022; Chiu, Adeli, and Niebles 2020), 3D convolutions (Tulyakov et al.
+2018), recurrent neural networks (RNNs) (Denton and Fergus 2018; Villar-Corrales et al.
+2022; Wang et al.
+2022), Transformers (Rakhimov et al.
+2021; Ye and Bilodeau 2022), or diffusion models (Höppe et al.
+2022; Ho et al.
+2022).
+Object-Centric Video Prediction
 Object-centric VP presents a structured approach that explicitly models the dynamics and interactions of individual objects to forecast future video frames.
-These methods typically involve three main steps: decomposing seed frames into object representations, forecasting future object states using a dynamics model, and rendering video frames from the predicted object representations.
-Various approaches have been proposed for this task, using different architectural designs such as RNNs (Creswell et al., 2021; Zoran et al., 2021; Nguyen et al., 2024) or transformers (Wu et al., 2021, 2023; Villar-Corrales et al., 2023; Song et al., 2023; Daniel & Tamar, 2024).
-Despite promising results, these models are currently limited to simple deterministic datasets or rely on action-conditioning (Mosbach et al., 2024) or inferred latent vectors (Villar-Corrales & Behnke, 2025).
-In contrast, our proposed model forecasts future video frames conditioned on past object slots and textual descriptions.
-2.2.2 Text-Conditioned Video Prediction
-This category of VP models leverages text descriptions to provide appearance, motion and action cues that guide the generation of future frames.
-This task was first proposed by Hu et al.
-(2022), who utilized a VQ-VAE to encode the images into visual token representations, and modeled the image dynamics with an axial transformer that jointly processes such tokens along with text descriptions.
-TVP (Song et al., 2024) leverages RNNs to learn text representations, which condition a GAN-based framework in order to generate videos from a single frame.
-MMVG (Fu et al., 2023) combines a VQ-GAN (Esser et al., 2021) with a masked transformer predictor.
-More recently, several methods leverage diffusion models with text conditioning (Gu et al., 2024; Ni et al., 2023; Chen et al., 2023).
+These methods typically involve three main steps: parsing seed frames into object representations, forecasting future object states using a dynamics model, and rendering video frames from the predicted object representations.
+Various approaches have addressed this task using different architectural priors, such as RNNs (Creswell et al.
+2021; Nguyen et al.
+2024) or transformers (Wu, Yoon, and Ahn 2021; Wu et al.
+2023; Villar-Corrales, Wahdan, and Behnke 2023).
+Despite promising results, these models are limited to simple deterministic datasets or rely on action-conditioning (Mosbach et al.
+2025; Villar-Corrales and Behnke 2025).
+In contrast, our model forecasts future frames conditioned on past object slots and text descriptions.
+Text-Conditioned Video Prediction
+Text-conditioned VP models leverage text descriptions to provide appearance, motion and action cues that guide the generation of future frames.
+This task was first proposed by Hu, Luo, and Chen (2022), who utilized a VQ-VAE to encode images into visual token representations, and modeled the scene dynamics with an axial transformer to jointly process visual tokens with text descriptions.
+Similarly, approaches like TVP (Song et al.
+2024) and MMVG (Fu et al.
+2023) address this task using RNNs or masked transformers, respectively.
+More recently, several methods leverage diffusion models for text-guided VP (Blattmann et al.
+2023; Ni et al.
+2023; Chen et al.
+2024; Xing et al.
+2024; Chen et al.
+2023; Gu et al.
+2024).
+These approaches encode video frames into discrete token sequences via pretrained quantized autoencoders, and leverage pretrained language models to guide the diffusion-based generation with text features.
+To further improve temporal coherence and semantic alignment, several diffusion-based methods leverage specialized conditioning strategies (Xing et al.
+2024; Chen et al.
+2023) or attention mechanisms (Gu et al.
+2024), among others.
+While effective, these models operate on holistic or spatial representations, and require large-scale compute and data.
+In contrast, TextOCVP adopts a more structured and efficient approach, explicitly modeling object dynamics using slot representations, where each object in the scene is represented by a distinct embedding.
 Concurrently with our work, Wang et al.
-(2024) combine object-centric learning with text-conditional image-to-video generation.
-The authors propose an autoregressive diffusion model conditioned on object slots and text descriptions and evaluate on simple datasets.
-In contrast, we explicitly model the object dynamics with an autoregressive transformer, and evaluate our approach on more complex robotics simulations.
+(2025) and Jeong et al.
+(2025) combine object-centric learning with autoregressive diffusion models and transformers for text-guided video prediction on simple synthetic datasets.
+In contrast, we model object dynamics with an autoregressive transformer, evaluate on more complex robotics simulations, and perform an in-depth analysis of the properties of object-centric representations for text-conditioned video prediction.
 3 Method
-We propose TextOCVP – a novel object-centric model for text-conditioned image-to-video generation.
-Given an initial reference image $\textbf{X}_{1}$ and a text caption $\mathcal{C}$, TextOCVP generates the subsequent $T$ video frames $\hat{\mathbf{X}}_{2:{T+1}}$, which maintain a similar appearance and structural composition as the reference image, and follow the motion described in the text caption.
+We propose TextOCVP, a novel object-centric model for text-conditioned video prediction.
+Given an initial reference image $\textbf{X}_{1}$ and a text caption $\mathcal{C}$, TextOCVP generates the $T$ subsequent video frames $\hat{\mathbf{X}}_{2:{T+1}}$, which maintain a similar appearance and structural composition as the reference image, and follow the motion described in the text caption.
 TextOCVP, which is illustrated in Fig.
@@
 The object slots are fed to a text-conditioned transformer predictor (Sec.
-3.2), which jointly models their spatio-temporal relations, and incorporates the textual information from the caption $\mathcal{C}$ as a guidance for predicting the future object slots $\hat{\mathbf{S}}_{2:T+1}$.
-Finally, the predicted object slots are decoded in order to render object representations and video frames (Sec.
+3.2), which jointly models their spatio-temporal relations, and incorporates the textual information from the caption $\mathcal{C}$ as guidance for predicting the future object slots $\hat{\mathbf{S}}_{2:T+1}$.
+Finally, the predicted slots are decoded to render future video frames (Sec.
 3.3).
 We propose two different TextOCVP variants, which differ in the underlying object-centric decomposition modules.
-Specifically, $\text{TextOCVP}_{\text{SAVi}}$ leverages SAVi (Kipf et al., 2022), whereas $\text{TextOCVP}_{\text{DINO}}$ extends the DINOSAUR (Seitzer et al., 2023) framework for recursive object-centric video decomposition and video rendering.
+Specifically, $\text{TextOCVP}_{\text{SAVi}}$ leverages SAVi (Kipf et al.
+2022), whereas $\text{TextOCVP}_{\text{DINO}}$ extends the recent DINOSAUR (Seitzer et al.
+2023) framework for recursive object-centric video decomposition and video rendering.
+3.1 Scene Parsing
+The scene parsing module decomposes a video sequence $\mathbf{X}_{1:{\tau}}$ into a set of permutation-invariant object representations called slots $\mathbf{S}_{1:\tau}=(\textbf{S}_{1},...,\textbf{S}_{\tau})$, with $\textbf{S}_{t}=(\textbf{s}_{t}^{1},...,\textbf{s}_{t}^{N_{\mathbf{S}}})$, where each slot $\textbf{s}\in{\mathbb{R}}^{D}$ represents a single object.
+For scene parsing, we adopt the recursive object-centric video decomposition framework from Kipf et al.
+(2022).
+At time step $t$, the corresponding input frame $\textbf{X}_{t}$ is encoded with a feature extractor module into a set of $D_{h}$-dimensional feature maps $\textbf{h}_{t}\in{\mathbb{R}}^{L\times D_{h}}$ representing $L$ spatial locations.
+The feature extractor is a convolutional neural network in our $\text{TextOCVP}_{\text{SAVi}}$ variant and a DINO-pretrained vision transformer (Caron et al.
+2021) in $\text{TextOCVP}_{\text{DINO}}$.
+These feature maps are processed with Slot Attention (Locatello et al.
+2020), which updates the previous slots $\textbf{S}_{t-1}$ based on visual features from the current frame following an iterative attention mechanism.
+Namely, Slot Attention performs cross-attention, with the attention weights normalized over the slot dimension, thus encouraging the slots to compete to represent parts of the input.
+It then updates the slots using a Gated Recurrent Unit (Cho et al.
+2014) (GRU).
+Formally, Slot Attention updates the previous slots $\textbf{S}_{t-1}$ by:
+$\textbf{A}=\operatorname*{softmax}_{N_{\mathbf{S}}}\left(\frac{q(\textbf{S}_{t-1})k(\textbf{h}_{t})^{T}}{\sqrt{D}}\right)\in{\mathbb{R}}^{N_{\mathbf{S}}\times L},$ (1)
+$\textbf{S}_{t}=\text{GRU}(W_{t}v(\textbf{h}_{t}),\textbf{S}_{t-1})\;\text{with}\;W_{i,j}=\frac{\textbf{A}_{i,j}}{\sum_{l=1}^{L}\textbf{A}_{i,l}},$ (2)
+where $k,q$ and $v$ are learned linear projections that map input features and slots into a common dimension.
+The output of this module is a set of slots $\textbf{S}_{t}$ that represents the objects of the input frame.
 Figure 2: Overview of TextOCVP.
 TextOCVP parses the reference frame $\textbf{X}_{1}$ into object representations $\textbf{S}_{1}$.
-The text-conditioned object-centric predictor models object dynamics and interactions, incorporating information from the description $\mathbf{C}$ to predict future object states and frames.
-3.1 Scene Parsing
-The scene parsing module decomposes a video sequence $\mathbf{X}_{1:{\tau}}$ into a set of permutation-invariant object representations called slots $\mathbf{S}_{1:\tau}=(\textbf{S}_{1},...,\textbf{S}_{\tau})$, with $\textbf{S}_{t}=(\textbf{s}_{t}^{1},...,\textbf{s}_{t}^{N_{\mathbf{S}}})$, where each slot s represents a single object.
-For learning the object representations, we leverage the object-centric video decomposition framework proposed by Kipf et al.
-(2022).
-At time step $t$, the corresponding input frame $\textbf{X}_{t}$ is encoded with a feature extractor module into a set of $D_{h}$-dimensional feature maps $\textbf{h}_{t}\in{\mathbb{R}}^{L\times D_{h}}$ representing $L$ spatial locations.
-The feature extractor is convolutional neural network in our $\text{TextOCVP}_{\text{SAVi}}$ variant and a DINO-pretrained vision transformer (Caron et al., 2021) in $\text{TextOCVP}_{\text{DINO}}$.
-These feature maps are processed with a Slot Attention (Locatello et al., 2020) corrector, which updates the previous slots $\textbf{S}_{t-1}$ based on visual features from the current frame following an iterative attention mechanism.
-Namely, Slot Attention performs cross-attention, with the attention weights normalized over the slot dimension, thus encouraging the slots to compete to represent parts of the input.
-It then updates the slots using a Gated Recurrent Unit (Cho et al., 2014) (GRU).
-Formally, Slot Attention updates the previous slots $\textbf{S}_{t-1}$ by:
-$\textbf{A}=\operatorname*{softmax}_{N_{\mathbf{S}}}\left(\frac{q(\textbf{S}_{t-1})k(\textbf{h}_{t})^{T}}{\sqrt{D}}\right)\in{\mathbb{R}}^{N_{\mathbf{S}}\times L},$ (1)
-$\textbf{S}_{t}=\text{GRU}(W_{t}v(\textbf{h}),\textbf{S}_{t-1})\;\text{with}\;W_{i,j}=\frac{\textbf{A}_{i,j}}{\sum_{l=1}^{L}\textbf{A}_{i,l}},$ (2)
-where $k,q$ and $v$ are learned linear projections that map input features and slots into a common dimension.
-The final output of this module is the set of slots $\textbf{S}_{t}$, representing the objects of the input frame.
+The text-conditioned object-centric predictor models object dynamics and interactions, incorporating information from the description $\mathbf{C}$ to predict future object states $\hat{\mathbf{S}}_{2:T+1}$, which can be used to render future video frames $\hat{\mathbf{X}}_{2:{T+1}}$.
 3.2 Text-Conditioned Object-Centric Predictor
@@
 To condition the prediction process, the text description $\mathcal{C}$ is encoded into text token embeddings $\mathbf{C}$ employing an encoder-only transformer module.
-We experiment with different variants of this module, including a vanilla transformer encoder (Vaswani et al., 2017) or a pretrained T5 (Raffel et al., 2020) module.
-At time step $t$, the predictor receives as input the corresponding text embeddings $\mathbf{C}$, as well as the previous object slots $\mathbf{S}_{1:t}$, which are initially mapped via an MLP to the predictor token dimensionality $D_{\text{Pred}}$.
-Additionally, these tokens are augmented with a temporal positional encoding, which applies the same sinusoidal positional embedding to all tokens from the same time-step, thus preserving the inherent permutation-equivariance of the objects.
-Each layer of our predictor module mirrors the transformer decoder architecture (Vaswani et al., 2017).
+We experiment with different variants of this module, including a vanilla transformer encoder (Vaswani et al.
+2017) or a pretrained T5 (Raffel et al.
+2020) module.
+At time step $t$, the predictor receives as input the corresponding text embeddings $\mathbf{C}$, as well as the previous object slots $\mathbf{S}_{1:t}$, which are initially mapped via an MLP to the predictor token dimensionality.
+Additionally, these tokens are augmented with a temporal positional encoding, which applies the same sinusoidal positional embedding to all tokens from the same time step, thus preserving the inherent permutation-equivariance of the objects.
+Each layer of our predictor module mirrors the transformer decoder architecture (Vaswani et al.
+2017).
 First, a self-attention layer enables every slot to attend to all other object representations in the sequence, modeling the spatio-temporal relations between objects.
-Subsequently, a text-to-slot cross-attention layer enhances the slot representations by incorporating important features from the text embeddings, such as motion or appearance information.
-Finally, an MLP is independently applied for each token.
+Subsequently, a text-to-slot cross-attention layer enhances the slot representations by incorporating relevant features from the text embeddings, guiding the prediction process to align with the motion and dynamics described in the textual caption.
+Finally, an MLP is applied for each token.
 This process is repeated in every predictor layer, resulting in the predicted object slots of the subsequent time step $\hat{\textbf{S}}_{t+1}$.
-Furthermore, we apply a residual connection from $\textbf{S}_{t}$ to $\hat{\textbf{S}}_{t+1}$, which improves the temporal consistency of the predictions.
+Furthermore, we apply a residual connection from $\textbf{S}_{t}$ to $\hat{\textbf{S}}_{t+1}$, which improves the prediction temporal consistency.
 This process is repeated autoregressively to obtain slot predictions for $T$ subsequent time steps.
@@
 We leverage two variants of the video rendering module, for our $\text{TextOCVP}_{\text{SAVi}}$ and $\text{TextOCVP}_{\text{DINO}}$ variants, respectively.
-$\text{TextOCVP}_{\text{SAVi}}$ Decoder   This variant independently decodes each slot in $\hat{\textbf{S}}_{t}$ with a CNN-based Spatial Broadcast Decoder (Watters et al., 2019), rendering an object image $\textbf{o}_{t}^{n}$ and mask $\textbf{m}_{t}^{n}$ for each slot $\textbf{s}_{t}^{n}$.
+$\text{TextOCVP}_{\text{SAVi}}$ Decoder
+This variant independently decodes each slot in $\hat{\textbf{S}}_{t}$ with a CNN-based Spatial Broadcast Decoder (Watters et al.
+2019), rendering an object image $\textbf{o}_{t}^{n}$ and mask $\textbf{m}_{t}^{n}$ for each slot $\textbf{s}_{t}^{n}$.
 The object masks are normalized across the slot dimension, and the representations are combined via a weighted sum to render video frames:
 $\hat{\textbf{X}}_{t}=\sum_{n=1}^{N_{\mathbf{S}}}\textbf{o}_{t}^{n}\cdot\tilde{\textbf{m}}_{t}^{n}\;\;\;\text{with}\;\;\;\tilde{\textbf{m}}_{t}^{n}=\operatorname*{softmax}_{N_{\mathbf{S}}}(\textbf{m}_{t}^{n}).$ (3)
-Figure 3: Text-conditioned object-centric predictor.
-$\text{TextOCVP}_{\text{DINO}}$ Decoder   This decoder variant decodes the object slots in two distinct stages.
-First, following DINOSAUR (Seitzer et al., 2023), an MLP-based Spatial Broadcast Decoder (Watters et al., 2019) is used to generate object features along with their corresponding masks.
+Figure 3: Overview of our text-conditioned object-centric video predictor.
+Our predictor employs a transformer-based architecture to model the spatio-temporal dynamics of object slots via self-attention, while incorporating textual guidance through a text-to-slot cross-attention mechanism.
+$\text{TextOCVP}_{\text{DINO}}$ Decoder
+This decoder variant decodes the object slots in two distinct stages.
+First, following DINOSAUR (Seitzer et al.
+2023), an MLP-based Spatial Broadcast Decoder (Watters et al.
+2019) is used to generate object features along with their corresponding masks.
 Similar to the $\text{TextOCVP}_{\text{SAVi}}$ decoder, the object masks are normalized and combined with the object features in order to reconstruct the encoded features $\hat{\textbf{h}}_{t}\in{\mathbb{R}}^{L\times D_{h}}$.
@@
 Our proposed TextOCVP is trained in two different stages.
-Object-Centric Learning   We first train the scene parsing and video rendering modules for decomposing video frames into object-centric representations by minimizing a reconstruction loss.
-In the $\text{TextOCVP}_{\text{SAVi}}$ variant, these modules are trained simply by reconstructing the input images, whereas in $\text{TextOCVP}_{\text{DINO}}$ they are trained by jointly minimizing an image and a feature reconstruction loss:
-$\mathcal{L}_{\text{SAVi}}=\frac{1}{T}\sum_{t=1}^{T}||\hat{\textbf{X}}_{t}-\textbf{X}_{t}||,$ (4)
-$\mathcal{L}_{\text{DINO}}=\frac{1}{T}\sum_{t=1}^{T}||\hat{\textbf{X}}_{t}-\textbf{X}_{t}||+||\hat{\textbf{h}}_{t}-\textbf{h}_{t}||.$ (5)
-Predictor Training   Given the pretrained scene parsing and rendering modules, we train our TextOCVP predictor for text-conditioned video generation using a dataset containing paired videos and text descriptions.
-Namely, given the object representations from a reference frame $\textbf{S}_{1}$ and the textual description $\mathcal{C}$, our TextOCVP predictor forecasts the subsequent object slots $\hat{\textbf{S}}_{2}$, which are decoded into a predicted video frame $\hat{\textbf{X}}_{2}$.
+Object-Centric Learning
+We first train the scene parsing and video rendering modules for parsing video frames into object-centric representations by minimizing a reconstruction loss.
+In the $\text{TextOCVP}_{\text{SAVi}}$ variant ($\mathcal{L}_{\text{SAVi}}$), these modules are trained by reconstructing the input images, whereas in $\text{TextOCVP}_{\text{DINO}}$ ($\mathcal{L}_{\text{DINO}}$) they are trained by jointly minimizing an image and a feature reconstruction loss:
+[…]
+[…]
+Predictor Training
+Given the pretrained scene parsing and rendering modules, we train our TextOCVP predictor for text-conditioned video prediction using a dataset containing paired videos and text descriptions.
+Namely, given the object representations from a reference frame $\textbf{S}_{1}$ and the textual embeddings $\mathbf{C}$, our TextOCVP predictor forecasts the subsequent object slots $\hat{\textbf{S}}_{2}$, which are decoded into a predicted video frame $\hat{\textbf{X}}_{2}$.
 This process is repeated autoregressively, i.e.
@@
 This autoregressive training, in contrast to teacher forcing, enforces our predictor to operate with imperfect inputs, leading to better modeling of long-term dynamics at inference time.
-Our predictor is trained by minimizing the following combined loss:
-$\mathcal{L}_{\text{TextOCVP}}=\frac{1}{T}\sum_{t=2}^{T+1}\lambda_{\text{Img}}\mathcal{L}_{\text{Img}}+\lambda_{\text{Slot}}\mathcal{L}_{\text{Slot}},$ (6)
-with $\mathcal{L}_{\text{Img}}=||\hat{\textbf{X}}_{t}-\textbf{X}_{t}||_{2}^{2}\;\;\text{and}\;\;\mathcal{L}_{\text{Slot}}=||\hat{\textbf{S}}_{t}-\textbf{S}_{t}||_{2}^{2},$ (7)
+Our predictor is trained by minimizing the following loss:
+$\mathcal{L}_{\text{TextOCVP}}=\frac{1}{T}\sum_{t=2}^{T+1}\left(\lambda_{\text{Img}}\mathcal{L}_{\text{Img}}+\lambda_{\text{Slot}}\mathcal{L}_{\text{Slot}}\right),$ (6)
+[…]
 where $\mathcal{L}_{\text{Img}}$ measures the future frame prediction error, and $\mathcal{L}_{\text{Slot}}$ enforces the alignment of the predicted object slots with the actual inferred object-centric representations.
-Inference At inference time, TextOCVP receives as input a single reference frame and a language instruction.
+Inference
+At inference time, TextOCVP receives as input a single reference frame and a language instruction.
 Our model parses the seed frame into object slots and autoregressively predicts future object states and video frames conditioned on the given textual description.
@@
 4.1 Experimental Setup
-4.1.1 Datasets
-We evaluate the performance of TextOCVP for text-driven image-to-video generation on two different datasets, namely CATER and CLIPort. Further dataset details are provided in Appendix D.
-CATER   CATER (Girdhar & Ramanan, 2020) is a dataset that consists of long video sequences, featuring two simultaneously moving 3D objects, with the motion described by a textual caption.
-We used the CATER-Hard variant introduced by Hu et al.
-(2022).
-This dataset contains 30,000 sequences featuring between three and eight objects, two of which follow a predefined action pattern.
-In our experiments, we resize the images to $64\times 64$.
-CLIPort   CLIPort (Shridhar et al., 2022) is a robot manipulation dataset consisting of video-caption pairs.
-We employ 21,000 $336\times 336$ sequences of the Put-Block-In-Bowl task.
-Each sequence features a table with multiple bowls and blocks of different color, placed at random positions on the table.
-The corresponding caption describes the action of the robot arm picking a block of a specific color and putting it into a specific bowl.
-4.1.2 Baselines
+Datasets
+We evaluate TextOCVP for text-conditioned video prediction on the CATER and CLIPort datasets:
+CATER (Girdhar and Ramanan 2020) is a dataset of long video sequences showing 3D objects in motion, each accompanied by a descriptive caption.
+We used the CATER-hard variant (Hu, Luo, and Chen 2022), containing 30,000 sequences with $64\times 64$ frames featuring between two and eight objects, two of which follow the scripted motion.
+CLIPort (Shridhar, Manuelli, and Fox 2022) is a robot manipulation dataset featuring video-caption pairs.
+We employ 21,000 $336\times 336$ sequences of the Put-Block-In-Bowl task, where the caption describes placing a block into a bowl.
+Baselines
 […]
-4.1.3 Implementation Details
-All our models are implemented in PyTorch and trained on a single NVIDIA A6000 (48Gb) GPU.
-$\text{TextOCVP}_{\text{SAVi}}$ closely follows Kipf et al.
-(2022) for the design of the scene parsing and video rendering modules.
-$\text{TextOCVP}_{\text{DINO}}$ leverages DINOv2 (Oquab et al., 2023) as image encoder, uses a four-layer MLP-based Spatial-Broadcast Decoder (Watters et al., 2019), which is shared among all slots, to decode the object slots into object features and masks.
-Additionally, it uses a CNN decoder to map the reconstructed scene features back to images.
-On the CATER dataset, we use the $\text{TextOCVP}_{\text{SAVi}}$ variant with eight 128-dimensional object slots; whereas on CLIPort we employ the $\text{TextOCVP}_{\text{DINO}}$ variant with ten 128-dimensional slots.
-Our predictor module is an eight-layer transformer with 512-dimensional tokens, eight attention heads, and a hidden dimension of 1024.
-Further implementation details are provided in Appendix B.
-4.2 Results
-4.2.1 CATER Results
-Table 1: Quantitative evaluation on CATER for prediction horizons of $T=9$ and $T=19$. TextOCVP outperforms all baselines. Best two results are highlighted in boldface and underlined, respectively.
-CATER1→9 CATER1→19
-[…]
-OCVP 0.874 0.078 0.854 0.101
-Non-OC TF 0.874 0.092 0.849 0.112
-SEER 0.723 0.245 0.535 0.299
-MAGE 0.877 0.108 0.871 0.111
-TextOCVP 0.924 0.035 0.904 0.042
-On the CATER dataset, we train the models to generate nine future frames given a single reference frame and a text caption.
-In Tab.
-1 we report quantitative evaluations on CATER using the same setting as in training, i.e.
-predicting $T=9$ frames, as well as when predicting $T=19$ future frames.
-In both settings, TextOCVP outperforms all other models, demonstrating superior perceptual quality.
-[…]
+Refer to captiont=1t=122101020203030MAGE TextOCVP
 ‘the medium green metal sphere is sliding
@@
 is picked up and placed to (-3, 1)’
-Figure 4: Qualitative evaluation for text-driven image-to-video generation on CATER. Top row shows ground truth frames. TextOCVP generates sharp future frames, whereas MAGE blurs and misses objects.
+Figure 4: Text-guided video prediction result on CATER. Top row shows ground truth. TextOCVP predicts sharp and accurate frames, whereas MAGE blurs and misses objects.
+Implementation Details
+All models are implemented in PyTorch and trained on a single NVIDIA A6000 (48Gb) GPU.
+$\text{TextOCVP}_{\text{SAVi}}$ closely follows Kipf et al.
+(2022) for scene parsing and video rendering.
+$\text{TextOCVP}_{\text{DINO}}$ uses DINOv2 (Oquab et al.
+2024) as image encoder, a four-layer MLP Spatial-Broadcast Decoder (Watters et al.
+2019) to decode the slots into object features and masks, and a CNN decoder to map the reconstructed scene features back to images.
+On CATER, we use the $\text{TextOCVP}_{\text{SAVi}}$ variant with eight 128-dimensional object slots, whereas on CLIPort we employ the $\text{TextOCVP}_{\text{DINO}}$ variant with ten 128-dimensional slots.
+Our predictor module is an eight-layer transformer with 512-dimensional tokens, eight attention heads, and a hidden dimension of 1024.
+Further experiment details are provided in Appendix B-E.
+4.2 Results
+CATER Results
+On CATER, we train the models to predict nine future frames given a single reference frame and a text caption.
+In Table 1(a) we report quantitative evaluations on CATER using the same setting as in training, i.e.
+predicting $T=9$ frames, as well as when predicting $T=19$ future frames.
+In both settings, TextOCVP outperforms all other models, demonstrating superior perceptual quality.
 Fig.
-4 shows a qualitative comparison between TextOCVP and MAGE.
+4 qualitatively compares TextOCVP and MAGE.
 Our proposed method generates a sequence that is closely aligned to the ground-truth, whereas MAGE predictions feature multiple errors and artifacts, including missing objects, blurry contours, and significant changes on object shapes.
 This highlights the effectiveness of object-centric representations to accurately represent and model object dynamics.
-Additional qualitative results are provided in Appendix E.
-4.2.2 CLIPort Results
 […]
+(a) CATER Evaluation
+CATER1→9 CATER1→19
+[…]
+OCVP 0.874 0.078 4.16 0.854 0.101 8.08
+Non-OC 0.874 0.092 3.04 0.849 0.112 8.62
+SEER 0.723 0.245 11.23 0.535 0.299 17.29
+MAGE 0.877 0.108 3.46 0.871 0.111 5.88
+TextOCVP 0.922 0.036 2.16 0.902 0.044 5.09
+(b) CLIPort Evaluation
 CLIPort1→9 CLIPort1→19
 […]
-Non-OC TF 0.901 0.184 0.872 0.210
-SEER 0.887 0.141 0.622 0.331
-$\text{MAGE}_{\text{DINO}}$ 0.940 0.064 0.931 0.075
-TextOCVP 0.950 0.062 0.931 0.078
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-‘put the gray block in the brown bowl
-[…]
-[…]
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-‘put the blue block in the gray bowl
-[…]
-[…]
-Figure 5: Qualitative evaluation on CLIPort.
-Top rows depict the ground truth frames.
-TextOCVP successfully completes the pick-and-place task, whereas MAGE fails to predict robot motion (a) or the block disappears (b).
-[…]
-[…]
-However, we identify that TextOCVP often generates artifacts in the background and lacks textured details in certain objects, thus degrading its quantitative performance despite generating accurate future frames.
-Additional qualitative results are provided in Appendix E.
-4.3 Model Analysis
-4.3.1 Ablation Studies
+Non-OC 0.901 0.184 8.13 0.872 0.210 13.23
+SEER 0.887 0.141 6.80 0.622 0.331 8.29
+$\text{MAGE}_{\text{DINO}}$ 0.940 0.064 2.11 0.931 0.075 2.59
+TextOCVP 0.950 0.062 1.36 0.931 0.078 2.23
 Table 2: Ablation studies. We ablate different architectural design choices in our proposed TextOCVP model.
@@
 […]
-2 31.83 0.912 0.041
-4 32.58 0.921 0.036
-8 32.82 0.924 0.035
+2 0.908 0.045 2.26
+4 0.911 0.043 2.20
+8 0.922 0.036 2.16
 (b) Evaluation of different text encoders on CATER and CLIPort.
-CATER1→9 CLIPort1→9
+CATER1→19 CLIPort1→19
 […]
-Custom TF 31.62 0.909 0.045 - - -
-Frozen T5 31.83 0.912 0.041 26.99 0.950 0.062
-FT T5 - - - 26.85 0.947 0.065
+Custom TF 0.903 0.045 5.39 - - -
+Frozen T5 0.902 0.044 5.09 0.931 0.078 2.23
+FT T5 0.901 0.043 5.14 0.928 0.082 2.67
 (c) Effect of residual connection in predictor.
@@
 […]
-No 26.55 0.946 0.066
-Yes 26.99 0.950 0.062
+No 0.946 0.066 1.58
+Yes 0.950 0.062 1.36
 (d) Effect of the number of slots $N_{\mathbf{S}}$.
@@
 […]
-8 25.95 0.890 0.079
-10 26.99 0.950 0.062
-We perform several ablation studies to support and validate the architectural choices of our model components and their impact on TextOCVP’s image-to-video generation performance.
-The results are presented in Tab.
-6(d).
-Number of Layers (Tab.
-6(a))   TextOCVP performs best with $N_{\text{Pred}}=8$ predictor layers.
-Scaling beyond eight layers was not explored, as the performance improvements when increasing from four to eight layers were marginal.
-Residual Connection (Tab.
-6(c))   Applying a residual connection to the predictor output, $\hat{\textbf{S}}_{t+1}:=\hat{\textbf{S}}_{t+1}+\textbf{S}_{t}$, enhances the model’s performance by improving the temporal consistency of the predicted slots.
-Text Encoder (Tab.
-6(b))   We evaluate the performance of different text encoders, including a vanilla transformer, a frozen T5 encoder, and a fine-tuned (FT) T5.
-A frozen T5 module led to the best TextOCVP performance on both CATER and CLIPort datasets.
-We argue that given the relatively small vocabulary size on CLIPort, fine-tuning T5 did not prove to be beneficial.
-Number of Slots (Tab.
-6(d))   Using $N_{\mathbf{S}}=10$ slots resulted in better performance for TextOCVP on CLIPort, even though each scene can be described with eight slots (six objects, robot arm and background).
-This observation suggests that additional slots can be used as registers in order to help with internal computations (Darcet et al., 2024).
-Table 4: Quantitative evaluation on a CLIPort test set with unseen colors. We report the video generation performance and its drop relative to the known-color dataset.
+8 0.946 0.079 5.62
+10 0.950 0.062 1.36
+CLIPort Results
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+‘put the blue block in the gray bowl
+[…]
+TextOCVPt=1t=12020303040405050Missing
+[…]
+[…]
+[…]
+Nevertheless, we identify that TextOCVP often generates visual artifacts and lacks fine-grained textures, which can affect its quantitative scores despite accurate predictions.
+Further figures and results are provided in Appendix F.
+4.3 Model Analysis
+Ablation Studies
+We perform several ablation studies to support and validate the architectural choices of our model components and their impact on TextOCVP’s video prediction performance.
+The results are presented in Table 2(d).
+Number of Layers (Table 2(a))   TextOCVP performs best with $N_{\text{P}}=8$ predictor layers, outperforming shallower variants.
+We did not explore scaling beyond eight layers due to the significant increase in the number of parameters.
+Residual Connection (Table 2(c))   Incorporating a residual connection in the predictor, $\hat{\textbf{S}}_{t+1}:=\hat{\textbf{S}}_{t+1}+\textbf{S}_{t}$, enhances the model’s performance by improving the temporal consistency of the predicted slots.
+Text Encoder (Table 2(b))   We evaluate the performance of different text encoders, including a vanilla transformer, a frozen T5 encoder, and a fine-tuned (FT) T5.
+A frozen T5 module led to the best TextOCVP performance on the CLIPort dataset.
+Although all encoders led to similar performance on the CATER dataset, we opted for the frozen T5 to avoid unnecessarily increasing the model’s parameter count.
+Number of Slots (Table 2(d))   Using $N_{\mathbf{S}}=10$ slots resulted in better performance for TextOCVP on CLIPort, even though each scene can be described with eight slots (six objects, robot arm and background).
+This observation suggests that additional slots can be used as registers in order to help with internal computations (Darcet et al.
+2024).
+Table 3: Quantitative evaluation on a CLIPort test set with unseen colors. We report the video prediction performance and its drop relative to the known-color dataset.
 CLIPort1→9 CLIPort1→19
@@
 TextOCVP 0.946-0.4% 0.066-6.4% 0.927-0.4% 0.083-6.4%
-4.3.2 Model Robustness
-We evaluate the performance of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set featuring color variations in the text instructions that were not encountered during training.
-This evaluation measures how well the models generalize to unseen scene features, highlighting their ability to handle novel configurations.
-Tab.
-4 presents the results on this evaluation set.
-We report both the video generation performance as well as the performance drop relative to the evaluation set with known colors.
-Our model outperforms $\text{MAGE}_{\text{DINO}}$ for both prediction horizons.
-Most notably, TextOCVP demonstrates significantly higher robustness to objects with previously unseen colors compared to the baseline, showing a much smaller drop in performance for the perceptual LPIPS metric.
-This highlights the effectiveness of leveraging object-centric representations for image-to-video generation and planning, in contrast to relying on holistic scene representations that struggle with novel scene compositions.
-4.3.3 Controllability
+Model Robustness
 ‘the large purple rubber cone is picked up
@@
 c)
-[…]
-Figure 7: Qualitative evaluation of TextOCVP controllability on CATER.
+t=1t=12215153030
+Figure 6: Evaluation of TextOCVP’s controllability on CATER.
 Top row shows the ground truth sequence.
-We underline the changed actions with respect to the original caption.
-We demonstrate TextOCVP’s controllability by generating multiple sequences from the same reference frame, each conditioned on a different text instruction.
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the cyan block in the brown bowl
-Original
-Caption
-Refer to captionRefer to captionRefer to caption‘put the blue block in the brown bowl
-Changed
-Block
-[…]
-Figure 8: Qualitative evaluation of TextOCVP controllability on CLIPort.
-Top row depicts ground-truth frames.
-TextOCVP correctly generates a sequence where the robot picks up the correct block and places it into the bowl as specified in the textual instruction.
-A key objective of text-driven image-to-video generation is to provide control over the model generations.
-This is achieved through language instructions that describe the objects in the scene and their expected motion.
-We qualitatively assess the controllability of our model on both the CATER and CLIPort datasets.
-CATER   In Fig.
-7, we showcase the control that TextOCVP provides over its generations.
-We qualitatively evaluate this ability by generating multiple videos conditioned on the same reference frame while varying the language instruction.
-We experiment with altering both the moving and target objects and their actions in the language instruction, as well as providing instructions that include a number of tasks beyond what the model was trained on.
+TextOCVP enables predicting different sequence continuations conditioned on distinct text instructions.
+We evaluate the generalization ability of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set featuring color variations in the text instructions that were not encountered during training.
+This setup assesses the models’ ability to generalize to novel scene configurations and adapt to unseen text-visual combinations.
+Table 3 presents the results on this evaluation set with unseen colors.
+We report both the video prediction performance as well as the performance drop relative to the evaluation set with known colors.
+Our model consistently outperforms $\text{MAGE}_{\text{DINO}}$ for both prediction horizons.
+Most notably, TextOCVP demonstrates significantly higher robustness to objects with previously unseen colors compared to the baseline, showing a much smaller drop in performance for the perceptual LPIPS metric.
+These results highlight the advantages of structured object-centric representations, which provide greater generalization and robustness in video prediction—particularly in contrast to holistic scene representations that struggle with novel scene compositions.
+Controllability
+A key objective of text-driven video prediction is to provide fine-grained control over the prediction process, which is enabled through language instructions describing the objects in the scene and their expected motion.
+In Fig.
+6, we qualitatively showcase the level of control that TextOCVP provides over its generations on the CATER dataset.
+We qualitatively evaluate this ability by generating multiple video continuations from the same reference frame while varying the language instruction.
+These variations include changes to both the target objects and their actions, as well as providing instructions that include a number of actions beyond what the model was trained on.
 As shown in Fig.
-7, TextOCVP successfully identifies the objects described in the text and executes the instructions accordingly.
-Notably, our model can distinguish between two nearly identical purple cones in the scene, despite their identical shapes and color, and correctly generates a sequence with the specified motion.
-This demonstrates that key object attributes, such as size or color, are effectively captured through the text-to-slot attention mechanism.
-Leveraging this information, the text-conditioned predictor accurately predicts the future motion of each individual object as specified in the instruction.
-These results highlight the benefits of combining object-centric representations with text-based conditioning.
-CLIPort   We perform a similar experiment on the CLIPort dataset, as shown in Fig.
-8.
-A single frame containing multiple colored blocks and bowls is provided to TextOCVP.
-By modifying the textual instruction, we can control which block the robot arm picks up and the bowl where it is placed.
-In both examples, TextOCVP successfully selects the correct block and places it in the specified bowl.
-Furthermore, the movement of the robot arm adapts to the motion described in the instruction, demonstrating accurate and responsive control.
+6, TextOCVP successfully identifies the relevant objects and executes the actions described in the text instructions.
+Notably, our model distinguishes between two nearly identical purple cones, despite the identical shapes and color, and correctly generates a sequence with the specified motion.
+These results show that key object attributes, such as size or color, are effectively captured through the text-to-slot attention mechanism, enabling the predictor to accurately forecast the motion of each individual object.
+This highlights the benefits of combining object-centric representations with text guidance for fine-grained, controllable video prediction.
+Interpretability
+Refer to captiont=1t=1Refer to caption55Refer to caption1010Refer to caption2020Refer to caption3030
+‘the cone is sliding to (-3, -3).
+the snitch is picked up and placed to (-1, 2 ).’
+Refer to captionRefer to caption1.00.50.0Refer to caption
+Slot 3
+the
+cone
+is
+sliding
+to
+-3
+-3
+the
+snitch
+is
+picked
+up
+and
+placed
+to
+-1
+2
+Refer to captionRefer to caption1.00.50.0Refer to caption
+Slot 4
+the
+cone
+is
+sliding
+to
+-3
+-3
+the
+snitch
+is
+picked
+up
+and
+placed
+to
+-1
+2
+Refer to captionRefer to caption1.00.50.0Refer to caption
+Slot 6
+the
+cone
+is
+sliding
+to
+-3
+-3
+the
+snitch
+is
+picked
+up
+and
+placed
+to
+-1
+2
+Figure 7: Text-to-slot attention weights for different objects in a video. Slots representing objects described in the caption focus on relevant text tokens, such as target locations.
+In Fig.
+7, we visualize the text-to-slot attention weights—averaged across attention heads—for different slots in a CATER sequence, revealing how textual information guides model predictions.
+We observe that slots corresponding to objects described in the textual instruction focus on relevant text tokens, such as target coordinate locations.
+These results demonstrate that the text-to-slot attention mechanism effectively aligns the textual information with the object-centric representations, enabling the model to generate accurate video sequences based on the given text instructions, while providing superior interpretability.
 5 Conclusion
-We presented TextOCVP, a novel object-centric model for text-driven image-to-video generation.
-Given a single input image and a text description, TextOCVP generates a sequence that matches the description by parsing the environment into object representations and modeling their dynamics conditioned on the textual instruction.
-This is achieved with a novel text-conditioned object-centric transformer, which predicts future object states by modeling the spatio-temporal relationships between objects while incorporating the guidance from text.
-Through extensive evaluations, we demonstrated that TextOCVP outperforms existing text-conditioned models for the image-to-video generation, highlighting our model’s ability to generate consistent long sequences and adapt its predictions based on language input.
-Moreover, we validated our architectural choices through ablation studies, highlighting the importance of combining textual and object-centric information.
-With its structured latent space and superior controllability, TextOCVP offers potential for robotics applications, where it can serve as a controllable world model, enabling efficient planning, reasoning, and decision-making.
+We presented TextOCVP, a novel object-centric model for text-conditioned video prediction.
+Given a single input image and a text description, TextOCVP generates frames by parsing the scene into object representations and modeling their dynamics conditioned on the textual instruction.
+This is accomplished through a text-conditioned object-centric transformer, which predicts future object states by modeling the spatio-temporal relationships between objects while incorporating textual guidance.
+Through extensive evaluations, we demonstrated that TextOCVP outperforms existing models for text-driven video prediction from a single frame, highlighting our model’s ability to predict over long prediction horizons and adapt its predictions based on the provided textual description.
+Moreover, we validated our architectural choices through ablation studies, highlighting the importance of combining textual and object-centric information, as well as demonstrating TextOCVP’s robustness and interpretability.
+With its structured latent space and superior controllability, TextOCVP offers potential for applications in robotics, where it can serve as a controllable world model, enabling efficient planning, reasoning, and decision-making.
 Acknowledgment
 This work was funded by grant BE 2556/16-2 (Research Unit FOR 2535 Anticipating Human Behavior) of the German Research Foundation (DFG)
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
 Appendix A Limitations and Future Work
+A.1 Limitations
 While TextOCVP demonstrates promising results for text-guided object-centric video prediction, it presents some limitations that we plan to address in future work:
-Prediction Artifacts   TextOCVP occasionally generates artifacts in the predicted frames, such as blurriness, inconsistent object appearances, lack of textured details in the objects, or visual artifacts in the background.
-We believe that these limitations stem from the video rendering module, which might lack the representational power to reconstruct precise image details from the latent space representation.
-Poor Temporal Consistency   We observe that TextOCVP’s predictions often lack temporal consistency when forecasting for long prediction horizons ($T>30$).
-We attribute this limitation to the fact that TextOCVP is trained by optimizing reconstruction losses, which do not penalize temporal inconsistencies.
-Future Work   To address these limitations, we plan to extend our TextOCVP framework with more powerful decoder modules, such as autoregressive transformers (Singh et al., 2022) or diffusion models (Jiang et al., 2023), as well as scale our predictor module.
-Furthermore, we plan to incorporate temporal discriminators (Clark et al., 2020) to improve the temporal consistency of the predicted video frames.
-We believe that exploring these architectural modifications will enable us to utilize TextOCVP as a world model in complex real-world robotic environments.
+Prediction Artifacts
+TextOCVP occasionally generates artifacts in the predicted frames, such as blurriness, inconsistent object appearances, lack of textured details, or visual artifacts in the background.
+We believe that these limitations stem from the video rendering module, which might lack the representational power to reconstruct precise image details from the object-centric latent space representation.
+Limited Temporal Consistency
+While TextOCVP produces plausible predictions, we observe that its temporal consistency can degrade when forecasting for long prediction horizons ($T>30$), occasionally resulting in object jittering or instability.
+We attribute this limitation to the fact that TextOCVP is trained to predict only up to nine future frames ($T=9$) and only optimizing reconstruction losses, which do not penalize such temporal inconsistencies.
+A.2 Future Work
+In future work we aim to address these limitations.
+Our TextOCVP model is designed with a modular architecture, enabling for both improvements and flexible swapping of components—such as the parsing, predictor, or video rendering modules—to seamlessly improve the entire pipeline.
+To address the prediction artifacts, we plan to extend our TextOCVP framework with more powerful decoder modules, such as autoregressive transformers (Singh, Wu, and Ahn 2022) or diffusion models (Jiang et al.
+2023), as well as scale our predictor module.
+Furthermore, we plan to incorporate temporal discriminators (Clark, Donahue, and Simonyan 2020) to improve the temporal consistency of the predicted video frames.
+We believe that exploring these architectural modifications can overcome the aforementioned limitations and will enable the scaling of TextOCVP to real-world robotic environments.
 Appendix B Implementation Details
@@
 B.1 $\text{TextOCVP}_{\text{DINO}}$
-$\text{TextOCVP}_{\text{DINO}}$ variant consists of our proposed text-conditioned predictor module and an object-centric decomposition module that extends the DINOSAUR (Seitzer et al., 2023) framework for recursive object-centric video decomposition and video rendering.
-Text-Conditioned Predictor   The predictor is composed of $N_{\text{Pred}}=8$ identical layers, each containing 8-head attention mechanisms and an MLP with a single hidden layer of dimension 1024 and a ReLU activation function.
+The $\text{TextOCVP}_{\text{DINO}}$ variant consists of our proposed text-conditioned predictor module and an object-centric decomposition module that extends the DINOSAUR (Seitzer et al.
+2023) framework for recursive object-centric video decomposition and video rendering.
+Text-Conditioned Predictor
+The predictor is composed of $N_{\text{P}}=8$ identical layers, each containing 8-head attention mechanisms and an MLP with a single hidden layer of dimension 1024 and a ReLU activation function.
 Furthermore, the predictor uses an embedding dimensionality of 512, context window size of ten frames, and applies a residual connection from the predictor input to its output.
-Text Encoder   $\text{TextOCVP}_{\text{DINO}}$ leverages a pretrained and frozen small version of T5 encoder (Raffel et al., 2020), which consists of 6 T5 blocks.
+Text Encoder
+$\text{TextOCVP}_{\text{DINO}}$ leverages a pretrained and frozen small version of T5 encoder (Raffel et al.
+2020), which consists of six T5 blocks.
 This text encoder uses a vocabulary with size 32,128.
-Scene Parsing   The scene parsing module generates $N_{\mathbf{S}}=10$ slots of dimension 128.
-As feature extractor, we use DINOv2 ViT-Base (Oquab et al., 2023), featuring 12 layers, using a patch size of 14, and producing patch features with dimension $D_{h}=768$.
+Scene Parsing
+The scene parsing module generates $N_{\mathbf{S}}=10$ slots of dimension 128.
+As feature extractor, we use DINOv2 ViT-Base (Oquab et al.
+2024), featuring 12 layers, using a patch size of 14, and producing patch features with dimension $D_{h}=768$.
 The Slot Attention corrector module processes the first video frame with three iterations in order to obtain a good initial object-centric decomposition, and a single iteration for subsequent frames, which suffices to recursively update the slot representation.
 The initial object slots $\textbf{S}_{0}$ are randomly sampled from a Gaussian distribution with learned mean and covariance.
-We use a single Transformer encoder block as the transition function, which consists of a four attention heads and an MLP with hidden dimension 512.
-Video Rendering   The video rendering module consists of two distinct decoders.
-First, a four-layer MLP-based Spatial Broadcast Decoder (Watters et al., 2019) with hidden dimension 1024 reconstructs the patch features from the slots.
-Then, to reconstruct the images from the features, we implement a CNN-based decoder with four convolutional layers, where each layer uses $3\times 3$ kernels.
-Every layer is followed by bilinear upsampling and a ReLU activation function.
-A final convolutional layer is applied to map to the RGB channels of the image.
-Training   We train our model for object-centric decomposition using videos sequences of length five frames for 1000 epochs.
-We use batch size of 16, the Adam optimizer (Kingma & Ba, 2015), and a base learning rate of $4\times 10^{-4}$, which is linearly warmed-up for the first 10000 steps, followed by cosine annealing for the remaining of the training process.
+We use a single Transformer encoder block as the transition function, which consists of four attention heads and an MLP with a hidden dimension of 512.
+Video Rendering
+The video rendering module consists of two distinct decoders.
+First, a four-layer MLP-based Spatial Broadcast Decoder (Watters et al.
+2019) with hidden dimension 1024 reconstructs the patch features from the slots.
+Then, a CNN-based decoder reconstructs full-resolution images from these features.
+It consists of four convolutional layers, each using $3\times 3$ kernels, a ReLU activation function and bilinear upsampling.
+A final convolutional layer and bilinear interpolation are applied to map the outputs to the RGB channels and spatial dimensions of the image.
+Training
+We train our model for object-centric decomposition using video sequences of length five frames for 1000 epochs.
+We use batch size of 16, the Adam optimizer (Kingma and Ba 2015), and a base learning rate of $4\times 10^{-4}$, which is linearly warmed-up for the first 10000 steps, followed by cosine annealing for the remaining of the training process.
 Moreover, we clip the gradients to a maximum norm of 0.05.
-The predictor module is trained given a frozen and pretrained object-centric decomposition model for 700 epochs.
+The predictor module is trained given the frozen and pretrained object-centric decomposition model for 700 epochs to predict the subsequent nine frames using a single seed frame.
 The predictor is trained using the same hyper-parameters as for object-centric decomposition.
@@
 B.2 $\text{TextOCVP}_{\text{SAVi}}$
-$\text{TextOCVP}_{\text{SAVi}}$ uses the same text-conditioned predictor and text-encoder architectures as $\text{TextOCVP}_{\text{DINO}}$, but employs SAVi (Kipf et al., 2022) as the object-centric decomposition module.
-Scene Parsing   The scene parsing module generates $N_{\mathbf{S}}=8$ slots of dimension 128.
+$\text{TextOCVP}_{\text{SAVi}}$ uses the same text-conditioned predictor and text-encoder architectures as $\text{TextOCVP}_{\text{DINO}}$, but employs SAVi (Kipf et al.
+2022) as the object-centric decomposition module.
+Scene Parsing
+The scene parsing module generates $N_{\mathbf{S}}=8$ slots of dimension 128.
 Following Kipf et al.
@@
 The Slot Attention corrector follows the same structure as in $\text{TextOCVP}_{\text{DINO}}$.
-Video Rendering   Following Kipf et al.
-(2022), we utilize a CNN-based Spatial Broadcast Decoder (Watters et al., 2019) with four convolutional layers with 32 kernels of size $5\times 5$ , $\text{stride}=1$, and $\text{padding}=2$, and a final convolutional layer which maps to four channels (RGB + alpha mask).
-Training   We train our model for object-centric decomposition using video sequences of length ten frames for 1000 epochs, using batch size of 64, and an initial learning rate of $10^{-4}$, which is warmed up for 2500 steps, followed by cosine annealing for the remaining of the training process.
+Video Rendering
+Following Kipf et al.
+(2022), we utilize a CNN-based Spatial Broadcast Decoder (Watters et al.
+2019) with four convolutional layers with 32 kernels of size $5\times 5$ , $\text{stride}=1$, and $\text{padding}=2$.
+A final convolutional layer maps from the hidden 32-channel representation to four output channels (RGB + alpha mask).
+Training
+We train our model for object-centric decomposition using video sequences of length ten frames for 1000 epochs, using batch size of 64, and an initial learning rate of $10^{-4}$, which is warmed up for 2500 steps, followed by cosine annealing for the remaining of the training process.
 Moreover, we clip the gradients to a maximum norm of 0.05.
-The predictor module is trained given a frozen and pretrained object-centric decomposition model for 1400 epochs.
+The predictor module is trained given the frozen and pretrained object-centric decomposition model for 1400 epochs to predict the subsequent nine frames using a single seed frame.
 The predictor is trained using the same hyper-parameters as for object-centric decomposition.
 In the predictor loss function $\mathcal{L}_{\text{TextOCVP}}$, we set $\lambda_{\text{Img}}=1$ and $\lambda_{\text{Slot}}=1$.
+Table 4: Number of learnable parameters in TextOCVP and baselines for experiments on the CLIPort dataset.
+Model # Parameters
+TextOCVP 33.76M
+Non-OC 34.16M
+$\text{MAGE}_{\text{DINO}}$ 32.11M
+SEER 405.89M
 Appendix C Baselines
-We employ four different baselines to compare againts our TextOCVP model on the image-to-video generation task on CATER and CLIPort datasets.
-To emphasize the importance of incorporating textual information, we include a comparison with OCVP-Seq (Villar-Corrales et al., 2023), a recent object-centric video prediction model that does not utilize text conditioning.
-Additionally, we evaluate a non-object-centric TextOCVP variant (Non-OC TF) that processes the input image into a single high-dimensional slot representation, instead of multiple object-centric slots, thus allowing us to evaluate the effect of object-centric representations.
-Moreover, we compare TextOCVP with two popular text-conditioned image-to-video generation baselines that do not incorporate object-centricity: MAGE (Hu et al., 2022) and SEER (Gu et al., 2024).
-We train both models on CATER and CLIPort closely following the original implementation details [^1] https://github.com/Youncy-Hu/MAGE [^2] https://github.com/seervideodiffusion/SeerVideoLDM/tree/main.
-C.1 MAGE
-MAGE is an autoregressive text-guided image-to-video generation framework that utilizes a VQ-VAE (Russakovsky et al., 2015) encoder-decoder architecture to learn efficient visual token representations.
+[…]
+[…]
+Table 5: Comparison of MAGE variants on CLIPort. $\text{MAGE}_{\text{DINO}}$ clearly outperforms the original MAGE variant.
+Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+MAGE 7.116 0.453 0.713
+$\text{MAGE}_{\text{DINO}}$ 23.723 0.940 0.064
+[…]
+MAGE is an autoregressive text-guided video prediction framework that utilizes a VQ-VAE (Van Den Oord, Vinyals et al.
+2017) encoder-decoder architecture to learn efficient visual token representations.
 A cross-attention module aligns textual and visual embeddings to produce a spatially-aligned motion representation termed Motion Anchor (MA), which is fused with visual tokens via an axial transformer for video generation.
 For experiments on CATER, we use a codebook size of $512\times 256$ with a downsampling ratio of four, whereas on CLIPort we use a codebook size of $512\times 1024$.
-Moreover, on CLIPort we replace MAGE’s standard CNN encoder and decoder with the DINOv2 ViT encoder and CNN decoder used in our $\text{TextOCVP}_{\text{DINO}}$ approach.
-This adjustment ensures a fair comparison and significantly enhances MAGE’s performance on CLIPort.
+To ensure a fair comparison with TextOCVP on CLIPort, we replace MAGE’s standard CNN encoder and decoder with the DINOv2 ViT encoder and CNN decoder used in our TextOCVP model.
 We refer to this modified version as $\text{MAGE}_{\text{DINO}}$.
-Additionally, to align with our experimental setup, we omit the speed parameter.
+Table 5 presents a comparison between the original MAGE model and $\text{MAGE}_{\text{DINO}}$ on CLIPort.
+The results demonstrate that $\text{MAGE}_{\text{DINO}}$ significantly outperforms the original variant, enabling a fair comparison with TextOCVP and other baselines on this benchmark.
 C.2 SEER
 SEER is a diffusion-based model for language-guided video prediction.
-It employs an Inflated 3D U-Net derived from a pretrained text-to-image 2D latent diffusion model (Rombach et al., 2022), extending it along the temporal axis and integrating temporal attention layers to simultaneously model spatial and temporal dynamics.
-For the language conditioning module, SEER introduces a novel Frame Sequential Text (FSText) Decomposer, which decomposes global instructions generated by the CLIP text encoder (Radford et al., 2021) into frame-specific sub-instructions.
+It employs an Inflated 3D U-Net derived from a pretrained text-to-image 2D latent diffusion model (Rombach et al.
+2022), extending it along the temporal axis and integrating temporal attention layers to simultaneously model spatial and temporal dynamics.
+For the language conditioning module, SEER introduces a novel Frame Sequential Text (FSText) Decomposer, which decomposes global instructions generated by the CLIP text encoder (Radford et al.
+2021) into frame-specific sub-instructions.
 These are aligned with frames using a transformer-based temporal network and injected into the diffusion process via cross-attention layers.
-We initialize SEER from a checkpoint pretrained on the Something-Something V2 dataset (Goyal et al., 2017), and further fine-tune it for a few epochs.
-We observed that incorporating a text loss enhanced SEER’s performance, while other hyper parameters were kept consistent with its original implementation.
+We initialize SEER from a checkpoint pretrained on the Something-Something V2 dataset (Goyal et al.
+2017), and further fine-tune it for a few epochs.
+We observed that incorporating a text loss enhanced SEER’s performance, while other hyper-parameters were kept consistent with its original implementation.
 Appendix D Datasets
-CATER   CATER (Girdhar & Ramanan, 2020) is a dataset that consists of long video sequences, each described by a textual caption.
+D.1 CATER
+CATER (Girdhar and Ramanan 2020) is a dataset that consists of long video sequences, each described by a textual caption.
 The video scenes consist of multiple 3D geometric objects in a 2D table plane, which is split into a $6\times 6$ grid with fixed axis, allowing the exact description of object’s positions using coordinates.
@@
 Furthermore, the initial positions are randomly selected from the plane grid, and the camera position is fixed for every sequence.
-In our work, we employ CATER-hard, which is a complete version of the CATER dataset, containing 30000 video-caption pairs.
+In our work, we employ CATER-hard, which is a complete version of the CATER dataset, containing 30000 video-caption pairs, with video frames resized to $64\times 64$.
 It includes 5 possible objects: cone, cube, sphere, cylinder, or snitch, which is a special small object in metallic gold color, shaped like three intertwined toruses.
@@
 The vocabulary size is 50.
-CLIPort   CLIPort (Shridhar et al., 2022) is a robot manipulation dataset, consisting of video-caption pairs, i.e.
+D.2 CLIPort
+CLIPort (Shridhar, Manuelli, and Fox 2022) is a robot manipulation dataset, consisting of video-caption pairs, i.e.
 long videos whose motion is described by a textual video caption.
@@
 Objects can be either a block or a bowl, and there is at least one of them in every sequence.
-The starting position of each object is random, with the only constrain to be placed on the table.
+The starting position of each object is random, with the only constraint being that it must be placed on the table.
 Each video describes the action of the robot arm picking a block, and putting it in a specific bowl.
@@
 Each individual object in the scene has a different color.
-In the train and validation set, the block and the bowl that are part of the caption can have one of the following colors: blue, green, red, brown, cyan, gray, or yellow, while in the test set they can have blue, green, red, pink, purple, white, or orange color.
+In the train and validation set, the block and the bowl that are part of the caption can have one of the following colors: blue, green, red, brown, cyan, gray, or yellow, while in the evaluation set with unseen colors they can have blue, green, red, pink, purple, white, or orange color.
 The other 4 objects, called distractors, can have any color.
@@
 The vocabulary size is 15.
-Appendix E Additional Results
 […]
+CATER1→9 CATER1→19 CLIPort1→9 CLIPort1→19
+[…]
+OCVP 29.08 0.874 0.078 4.16 28.11 0.854 0.101 8.08
+Non-OC 29.68 0.874 0.092 3.04 28.39 0.849 0.112 8.62 23.44 0.901 0.184 8.13 20.14 0.872 0.210 13.23
+SEER 22.05 0.723 0.245 11.23 16.05 0.535 0.299 17.29 21.01 0.887 0.141 6.80 11.30 0.622 0.331 8.29
+MAGE 34.91 0.877 0.108 3.46 34.76 0.871 0.111 5.88 23.72 0.940 0.064 2.11 22.27 0.931 0.075 2.59
+TextOCVP 32.98 0.922 0.036 2.16 31.29 0.902 0.044 5.09 26.99 0.950 0.062 1.36 23.88 0.931 0.078 2.23
+Appendix E Evaluation Metrics
+To measure TextOCVP’s video prediction performance and compare it with existing approaches, we evaluate the visual quality of predicted video frames using popular image- and video-based metrics.
+PSNR and SSIM (Wang et al. 2004) measure pixel-wise and statistical differences between the predicted and ground-truth video frames, respectively.
+LPIPS (Zhang et al.
+2018) is a perceptual metric that measures the visual similarity between two images based on deep features from pretrained neural networks, usually VGG (Simonyan and Zisserman 2015).
+Unlike pixel-wise metrics, LPIPS compares activations from multiple layers and captures differences in texture, structure, and semantics, thus aligning closely with human perception.
+While these metrics focus on frame-level visual quality, video prediction evaluation requires quantifying temporal consistency and motion realism.
+JEDi (Luo et al.
+2025) measures the quality and realism of generated videos by comparing feature distributions of generated and real videos, capturing both visual fidelity and motion dynamics.
+We prefer JEDi over FVD (Unterthiner et al.
+2019), as JEDi is less sensitive to small evaluation datasets, and more stable in low-motion synthetic scenarios.
+In our evaluations and comparisons with baselines, we favor the LPIPS and JEDi scores, which correlate well with human perception, while reporting other metrics for completeness.
+Appendix F Additional Results
+F.1 SAVi vs. DINOSAUR
+Current object-centric approaches for video prediction are limited to relatively simple synthetic datasets, and struggle to scale beyond scenes featuring simple 3D shapes with simple deterministic motion.
+We attribute this limitation primarily to the object-centric modules used for learning object representations.
+Motivated by this observation, we extend the recent DINOSAUR (Seitzer et al.
+2023) framework to recursively process video data and reconstruct video frames from their corresponding object-centric representations.
+To demonstrate the significance of the object-centric module in scaling to more complex datasets, we compare both SAVi and our Extended DINOSAUR trained on CLIPort.
+As illustrated in Fig.
+8, SAVi struggles to accurately represent the objects on the table, missing multiple objects and changing their shape and color.
+In contrast, the Extended DINOSAUR model successfully reconstructs the scene, closely resembling the input, while accurately representing each object.
+The visual features extracted by the DINOv2 (Oquab et al.
+2024) encoder contain high-level semantic information, and during training, the slots are specifically optimized to efficiently encode this information.
+This design enables the model to scale and handle more complex object-centric video data effectively.
+$t=1$
+22
+33
+44
+55
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+GT
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+SAVi
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Ext. DINO
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
 Ext. DINO
 Masks
-Figure 9: Comparison between SAVi and our Extended DINOSAUR (Ext. DINO) decomposition modules in reconstructing an CLIPort video sequence.
-E.1 SAVi vs. DINO
-Current object-centric approaches for video generation are limited to relatively simple synthetic datasets.
-We attribute this limitation primarily to the object-centric modules used for learning object representations.
-Motivated by this observation, we extend DINOSAUR (Seitzer et al., 2023) to handle video data and reconstruct images effectively.
-To demonstrate the significance of the object-centric module in scaling to more complex datasets, we input the same video sequence to both SAVi and our extended DINOSAUR trained on CLIPort.
-As illustrated in Fig.
-9, SAVi struggles to accurately model objects on the 2D plane, whereas our proposed object-centric model successfully reconstructs the scene, closely resembling the input.
-The visual features extracted by the DINOv2 (Oquab et al., 2023) encoder contain high-level semantic information, and during training, the slots are specifically optimized to efficiently encode this information.
-This design enables the model to scale and handle more complex object-centric video data effectively.
-E.2 Robustness to Number of Objects
-We evaluate the performance of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set consisting of scenes with a larger number of objects than encountered during training, i.e.
-the 2D table plane contains 8 instead of 6 objects.
-This evaluation demonstrates the robustness of models when generalizing to scenes with more objects.
-The quantitative results are presented in Tab.
-5.
-We report both the video generation performance and the performance drop relative to the original evaluation set.
-We observe that our model outperforms $\text{MAGE}_{\text{DINO}}$ on every metric, for both prediction horizons.
-Notably, TextOCVP demonstrates significantly higher robustness to scenes with more objects, as reflected by the smaller decline in performance for the perceptual LPIPS metric.
-This result is further illustrated in Fig.
-10, where we show a qualitative comparison of video generations for scenes with 8 objects.
-As observed, our model correctly illustrates the motion described in the text instructions, while $\text{MAGE}_{\text{DINO}}$ fails to generate accurate sequences, missing the target block, a behaviour observed in various examples.
-These results highlight once again the effectiveness of object-centric representations in image-to-video generation, as TextOCVP is able to generalize to scenes with more objects by simply increasing the number of slots.
+Figure 8: Comparison between SAVi and our Extended DINOSAUR (Ext.
+DINO) decomposition modules for reconstructing a CLIPort sequence.
+We visualize the reconstructed frames, as well as the slot masks obtained by Extended DINOSAUR.
+SAVi fails to reconstruct most objects, whereas Extended DINOSAUR accurately reconstructs the scene, while representing each object.
+F.2 Quantitative Comparison
+In the main paper, we present quantitative evaluations on the CATER and CLIPort datasets using three evaluation metrics, shown in two separate tables.
+For completeness, we report in Table 6 a text-guided video prediction evaluation of TextOCVP and multiple baseline models on CATER and CLIPort using four distinct evaluation metrics for different prediction horizons.
+[…]
+F.3 Computational Efficiency
+In Table 7 we report the average inference time for $T=1$ and $T=9$ frame predictions on CATER and CLIPort using a NVIDIA-A6000-48Gb GPU.
+[…]
+Table 7: Average inference (Inf.) time for $T=1$ and $T=9$ frame predictions on CATER and CLIPort.
+CATER Inf. [ms] CLIPort Inf. [ms]
+Model $T=1$ $T=9$ $T=1$ $T=9$
+MAGE 15.4 $\pm$ 1 111.1 $\pm$ 11 34.4 $\pm$ 3 198.8 $\pm$ 2
+[…]
 Table 8: Quantitative evaluation on a CLIPort test set with more objects. We report the video generation performance and its drop relative to the original dataset.
@@
 […]
-$\text{MAGE}_{\text{DINO}}$ 0.929 (-1.2%) 0.088 (-37.5%) 0.920 (-1.2%) 0.094 (-25.3%)
-TextOCVP 0.936 (-1.5%) 0.076 (-22.6%) 0.921 (-1.1%) 0.090 (-15.4%)
+$\text{MAGE}_{\text{DINO}}$ 0.929 -1.2% 0.088 -37% 0.920 -1.2% 0.094 -25%
+TextOCVP 0.936 -1.5% 0.076 -23% 0.921 -1.1% 0.090 -15%
+F.4 Robustness to Number of Objects
+We evaluate the performance of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set consisting of scenes with a larger number of objects than encountered during training, i.e.
+the 2D table plane contains 8 objects instead of the default 6 objects.
+This evaluation benchmarks the robustness of models when generalizing to scenes with a previously unseen number of objects.
+The quantitative results are presented in Table 8.
+We report both the video prediction performance, as well as the performance drop relative to the original evaluation set.
+We observe that our model outperforms $\text{MAGE}_{\text{DINO}}$ on every metric for both prediction horizons.
+Notably, TextOCVP demonstrates significantly higher robustness to scenes with more objects, as reflected by the smaller decline in performance for the perceptual LPIPS metric.
+This result is further illustrated in Fig.
+9, where we show qualitative comparisons of video generations for scenes with 8 objects.
+As observed, our model correctly predicts sequences following the motion described in the text instructions, whereas $\text{MAGE}_{\text{DINO}}$ fails to generate accurate sequences according to the descriptions.
+These results further demonstrate the effectiveness of object-centric representations for video prediction, as TextOCVP is able to generalize to scenes with more objects by simply increasing the number of slots.
+This flexibility is enabled by initializing object slots via sampling from a learned Gaussian, allowing the use of a variable number of slots at test time while breaking symmetry and preserving permutation invariance.
+Although the object slots do not specialize during training, they reliably bind to meaningful entities through iterative attention during inference.
+This results in robust and scalable scene decomposition, enabling accurate modeling of complex scenes with varying object counts.
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the gray block in the brown bowl Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
 TextOCVP
+t=1t=1551010202030304040a)
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the yellow block in the gray bowl Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+$\text{MAGE}_{\text{DINO}}$
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+TextOCVP
+t=1t=110102020303040405050b)
+Figure 9: Qualitative evaluation of $\text{MAGE}_{\text{DINO}}$ and TextOCVP on CLIPort sequences with more objects than those seen during training.
+TextOCVP correctly generates sequences according to the text instructions, whereas $\text{MAGE}_{\text{DINO}}$ misses the target bowl.
+Table 9: Impact of visual artifacts on TextOCVP’s performance on CLIPort. Cropping the artifact-prone bottom part of the predicted frames leads to a notable improvement in TextOCVP’s performance.
+CLIPort1→9 CLIPort1→19
 […]
-Figure 10: Qualitative evaluation of $\text{MAGE}_{\text{DINO}}$ and TextOCVP on a CLIPort sequence with more objects than those seen during training.
-TextOCVP correctly generates a sequence where the robot picks up the gray block and places it in the brown bowl, while $\text{MAGE}_{\text{DINO}}$ fails by missing the target bowl.
-E.3 Text-to-Slot Attention Visualizations
-An additional advantage of using object-centric representations is the improved interpretability.
-This can be shown in the visualizations of text-to slot attention weights, which help us understand how the textual information influences and guides the model predictions.
-First, in Fig.
-11, we visualize the text-to-slot attention weights for different cross-attention heads for a single slot that represents the rotating red cube.
-We observe that the slot attends to relevant text tokens from the input, such as the object shape, size and the action taking place.
+Full Image 0.950 0.062 0.931 0.078
+Excluded Bottom 0.953 0.050 0.932 0.069
+F.5 Impact of Visual Artifacts
+As already discussed in the main paper, TextOCVP occasionally generates visual artifacts on the CLIPort dataset, most noticeably in the bottom-center region of the frame, where blurry patches often appear.
+To assess the impact of these visual artifacts on TextOCVP’s quantitative performance, we evaluate our model after removing the last bottom rows from the predicted frames—an area that contains only background pixels and the artifacts.
+As shown in Table 9, removing these rows leads to improved results, particularly for the perceptual LPIPS metric.
+These findings verify that TextOCVP generates future frame predictions that closely follow the text description, and that its overall performance is underestimated due to localized visual artifacts.
+Mitigating such artifacts through improved architectures or training strategies presents an interesting direction for future work.
+F.6 Text-To-Slot Attention
+An additional advantage of using object-centric representations is improved interpretability.
+This is illustrated by visualizing the text-to-slot attention weights, which reveal how textual information guides the model’s predictions.
+In Fig.
+11, we visualize the text-to-slot attention weights from different cross-attention heads for a single slot that represents a rotating red cube.
+We observe that the slot attends to distinct relevant text tokens from the language description on each attention head, including the object shape (head 2), size (head 1) and the action taking place (head 4).
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+‘put the gray block in the brown bowl
+[…]
+TextOCVPt=1t=12020303040405050
+Figure 10: Qualitative result on CLIPort. Top row shows ground truth frames. TextOCVP completes the pick-and-place task, whereas MAGE fails to predict the robot motion.
+F.7 Additional Qualitative Evaluations
+Fig.
+10 shows an example where we evaluate TextOCVP and $\text{MAGE}_{\text{DINO}}$ for text-guided video prediction over a long prediction horizon of 50 frames.
+$\text{MAGE}_{\text{DINO}}$ fails to complete the task outlined in the textual description, as it stops generating consistent robot motion after 30 frames.
+In contrast, TextOCVP successfully predicts future frames where the robot completes the pick-and-place task.
+Figs. 12 and 13  show qualitative evaluations on CATER in which both MAGE and TextOCVP successfully predict sequence continuations following the instructions from the textual description.
+Fig. 14 illustrates an example where MAGE fails to generate a correct sequence, while TextOCVP successfully completes the task described by the text.
+Figs.
+15 and 16  show examples of TextOCVP’s control over the predictions.
+In both sequences, TextOCVP generates a correct sequence given the text instructions, and seamlessly adapts its generations to a modified version of the textual instructions.
 ‘the small red rubber cube is rotating.
@@
 rotating
-Figure 11: Visualization of the text-to-slot attention between an object and the textual caption.
-We visualize the text-to-slot attention weights for different heads.
-Our predictor module focuses on relevant text tokens to predict the scene dynamics, including the object size, shape or the action.
-In Fig.
-12, we additionally visualize the text-to-slot attention weights, averaged across attention heads, for different slots in a CATER sequence.
-We observe that slots that represent objects in the textual description attend to relevant text tokens, such as their target coordinate locations.
-These results demonstrate that the text-to-slot attention mechanism effectively aligns the textual information with the object-centric representations, enabling the model to generate accurate video sequences based on the given text instructions, while providing superior interpretability.
-[…]
-‘the cone is sliding to (-3, -3).
-the snitch is picked up and placed to (-1, 2 ).’
-Refer to captionRefer to caption1.00.50.0Refer to caption
-Slot 3
-the
-cone
-is
-sliding
-to
--3
--3
-the
-snitch
-is
-picked
-up
-and
-placed
-to
--1
-2
-Refer to captionRefer to caption1.00.50.0Refer to caption
-Slot 4
-the
-cone
-is
-sliding
-to
--3
--3
-the
-snitch
-is
-picked
-up
-and
-placed
-to
--1
-2
-Refer to captionRefer to caption1.00.50.0Refer to caption
-Slot 6
-the
-cone
-is
-sliding
-to
--3
--3
-the
-snitch
-is
-picked
-up
-and
-placed
-to
--1
-2
-Figure 12: Text-to-slot attention weights, averaged across attention heads, for different objects in the sequence.
-Slots that represent objects in the textual description attend to relevant text tokens, such as their target locations.
-E.4 Qualitative Evaluations
-Figs.
-13 and 14  show qualitative evaluations on CATER in which both MAGE and TextOCVP successfully generate a sequence following the instructions from the textual description.
-Fig.
-15 illustrates an example where MAGE fails to generate a correct sequence, while TextOCVP successfully completes the task described by the text.
-Figs.
-16 and 17  show examples of TextOCVP’s control over the predictions.
-In both sequences, TextOCVP generates a correct sequence given the text instructions, and seamlessly adapts its generations to a modified version of the textual instructions.
+Figure 11: Visualization of the text-to-slot attention weights between an object slot and the textual tokens.
+We visualize the text-to-slot attention weights for four different heads.
+Our predictor module focuses on distinct and relevant text tokens on each head, including the object size, shape or the action.
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
@@
 TextOCVP
-[…]
+t=1t=1551010151520203030
 Figure 12: Qualitative evaluation on CATER. Both MAGE and TextOCVP successfully generate a sequence following the instructions from the textual description.
@@
 TextOCVP
-[…]
+t=1t=1551010151520203030
 Figure 13: Qualitative evaluation on CATER. Both MAGE and TextOCVP successfully generate a sequence that illustrates the motion described in the text, but MAGE’s predictions are of a lower resolution.
@@
 TextOCVP
-[…]
+t=1t=1551010151520203030
 Figure 14: Qualitative evaluation on CATER.
@@
 Bowl
-[…]
+t=1t=110102020303040405050
 Figure 15: Qualitative evaluation of TextOCVP controllability on CLIPort. TextOCVP correctly generates a sequence where the robot picks up and places the block specified in the textual instruction.
@@
 Bowl
-[…]
+t=1t=110102020303040405050
 Figure 16: Qualitative evaluation of TextOCVP controllability on CLIPort. TextOCVP correctly generates a sequence where the robot picks up and places the block specified in the textual instruction.

### v3

@@
 Understanding and forecasting future scene states is critical for autonomous agents to plan and act effectively in complex environments.
-Object-centric models, with structured latent spaces, have shown promise in modeling object dynamics and predicting future scene states, but often struggle to scale beyond simple synthetic datasets and to integrate external guidance, limiting their applicability in robotics.
+Object-centric models, with structured latent spaces, have shown promise in modeling object dynamics and interactions in order to predict future scene states, but often struggle to scale beyond simple synthetic datasets and to integrate external guidance, limiting their applicability in robotic environments.
 To address these limitations, we propose TextOCVP, an object-centric model for video prediction guided by textual descriptions.
@@
 Additionally, we show that structured object-centric representations provide superior robustness to novel scene configurations, as well as improved controllability and interpretability, enabling more precise and understandable predictions.
-Videos and code are available at https://play-slot.github.io/TextOCVP.
+Videos and code are available on our project website.
+footnotetext: $\dagger$ denotes equal contribution.
 1 Introduction
 Understanding and reasoning about the environment is essential for enabling autonomous systems to better comprehend their surroundings, predict future events, and adapt their actions accordingly.
-Humans achieve these capabilities by perceiving the environment as a structured composition of individual objects that interact and evolve dynamically over time (Kahneman, Treisman, and Gibbs 1992).
-Neural networks equipped with such compositional inductive biases have shown the ability to learn structured object-centric representations of the world, which enable desirable properties, such as out-of-distribution generalization (Dittadi et al.
-2022), compositionality (Greff, Van Steenkiste, and Schmidhuber 2020), or sample efficiency (Mosbach et al.
-2025).
-Recent advances in unsupervised object-centric representation learning have progressed from extracting object representations in synthetic images (Locatello et al.
-2020; Lin et al.
-2020) to modeling objects in video (Kipf et al.
-2022; Singh, Wu, and Ahn 2022) and scaling to real-world scenes (Seitzer et al.
-2023; Zadaianchuk, Seitzer, and Martius 2024).
+Humans achieve these capabilities by perceiving the environment as a structured composition of individual objects that interact and evolve dynamically over time (Kahneman et al., 1992).
+Neural networks equipped with such compositional inductive biases have shown the ability to learn structured object-centric representations of the world, which enable desirable properties, such as out-of-distribution generalization (Dittadi et al., 2022), compositionality (Greff et al., 2020), or sample efficiency (Mosbach et al., 2025).
+Recent advances in unsupervised object-centric representation learning have progressed from extracting object representations in synthetic images (Locatello et al., 2020; Lin et al., 2020) to modeling objects in video (Kipf et al., 2022; Singh et al., 2022) and scaling to real-world scenes (Seitzer et al., 2023).
 These developments have enabled object-level dynamics modeling for future prediction and planning.
-Notably, approaches like SlotFormer (Wu et al.
-2023) or OCVP (Villar-Corrales, Wahdan, and Behnke 2023) introduced object-centric prediction models that explicitly model spatio-temporal relationships between objects, shifting away from image-level approaches that ignore scene compositionality.
+Notably, approaches like SlotFormer (Wu et al., 2023) or OCVP (Villar-Corrales et al., 2023) introduced object-centric prediction models that explicitly model spatio-temporal relationships between objects, shifting away from image-level approaches that ignore scene compositionality.
 Despite these advancements, current object-centric methods struggle with complex object appearances and dynamics, and lack mechanisms to incorporate external guidance, thus limiting their scalability and broader applicability.
-Refer to captiona) Object-Centric ParsingLearned ObjectsRefer to captionb) Text-Conditioned Video Prediction
-‘Put the green block
-in the cyan bowl.’
-‘Put the green block
-in the red bowl.’
+[…]
 Figure 1: Overview of TextOCVP.
@@
 We evaluate our approach for text-conditioned video prediction through extensive experiments on two distinct datasets.
-Our results show that TextOCVP outperforms existing text-conditioned video prediction methods by effectively leveraging structured object-centric representations, particularly on scenes featuring multiple moving objects.
-To verify the effectiveness of object-centric representations for text-guided video prediction, we conduct an in-depth analysis of TextOCVP.
+Our results show that TextOCVP outperforms several other text-conditioned video prediction methods by effectively leveraging structured object-centric representations, particularly on scenes featuring multiple moving objects.
+Furthermore, we conduct an in-depth analysis of TextOCVPto verify the effectiveness of object-centric representations for text-guided video prediction.
 Specifically, we demonstrate that its structured latent space enables accurate and controllable video prediction by aligning language instructions with the corresponding objects, outperforming baselines that rely on holistic scene representations.
@@
   • We propose TextOCVP, a text-guided video prediction model, featuring a text-conditioned object-centric predictor that integrates textual guidance into the prediction process via a text-to-slot attention mechanism.
-  • Through extensive evaluations, we show that TextOCVP outperforms existing text-conditioned video prediction models by leveraging object-centric representations.
+  • Through extensive evaluations, we show that TextOCVP outperforms other existing text-conditioned video prediction models by leveraging object-centric representations.
   • We demonstrate that TextOCVP is controllable, seamlessly adapting to diverse textual instructions, while exhibiting interpretability and robustness to novel setups.
@@
 2.1 Object-Centric Learning
-Representation learning, the ability to extract meaningful features from data, often improves model performance by enhancing its understanding of the input space (Bengio, Courville, and Vincent 2013).
+Representation learning—the ability to extract meaningful features from data—often improves model performance by enhancing its understanding of the input space (Bengio et al., 2013).
 Object-centric representation methods aim to parse an image or video into a set of object components in an unsupervised manner.
-These objects are typically represented as unconstrained embeddings (called slots) (Locatello et al.
-2020; Kipf et al.
-2022; Singh, Wu, and Ahn 2022), patch-based representations (Lin et al.
-2020), factored latent vectors (Greff et al.
-2019), or explicit object prototypes (Villar-Corrales and Behnke 2022; Monnier, Groueix, and Aubry 2020).
-These methods have demonstrated promise in learning object representations across diverse domains, ranging from synthetic images (Locatello et al.
-2020; Lin et al.
-2020) to videos (Kipf et al.
-2022; Singh, Wu, and Ahn 2022; Elsayed et al.
-2022), and real-world scenes (Seitzer et al.
-2023; Zadaianchuk, Seitzer, and Martius 2024).
-The learned object representations benefit downstream tasks, such as reinforcement learning (Mosbach et al.
-2025) or visual-question answering (Wu et al.
-2023).
+These objects are typically represented as unconstrained embeddings (called slots) (Locatello et al., 2020; Kipf et al., 2022; Singh et al., 2022), patch-based representations (Lin et al., 2020), factored latent vectors (Greff et al., 2019), or object prototypes (Villar-Corrales and Behnke, 2022; Monnier et al., 2020).
+These methods have demonstrated promise in learning object representations across diverse domains, ranging from synthetic images (Locatello et al., 2020; Lin et al., 2020) to videos (Kipf et al., 2022; Singh et al., 2022), and real-world scenes (Seitzer et al., 2023; Zadaianchuk et al., 2024).
+The learned object representations benefit downstream tasks, such as reinforcement learning (Mosbach et al., 2025) or visual-question answering (Wu et al., 2023).
 2.2 Video Prediction
-Video prediction (VP) is the task of forecasting the upcoming $T$ video frames conditioned on the preceding $C$ seed frames (Oprea et al.
-2020).
-Several methods have been proposed to address this challenge, leveraging 2D convolutions (Gao et al.
-2022; Chiu, Adeli, and Niebles 2020), 3D convolutions (Tulyakov et al.
-2018), recurrent neural networks (RNNs) (Denton and Fergus 2018; Villar-Corrales et al.
-2022; Wang et al.
-2022), Transformers (Rakhimov et al.
-2021; Ye and Bilodeau 2022), or diffusion models (Höppe et al.
-2022; Ho et al.
-2022).
-Object-Centric Video Prediction
+Video prediction (VP) is the task of forecasting the upcoming $T$ video frames conditioned on the preceding $C$ seed frames (Oprea et al., 2020).
+Several methods have been proposed to address this task, using 2D convolutions (Gao et al., 2022; Chiu et al., 2020), 3D convolutions (Tulyakov et al., 2018), recurrent neural networks (RNNs) (Denton and Fergus, 2018; Villar-Corrales et al., 2022; Wang et al., 2022), transformers (Rakhimov et al., 2021; Ye and Bilodeau, 2022), or diffusion models (Höppe et al., 2022; Ho et al., 2022).
+2.2.1 Object-Centric Video Prediction
 Object-centric VP presents a structured approach that explicitly models the dynamics and interactions of individual objects to forecast future video frames.
-These methods typically involve three main steps: parsing seed frames into object representations, forecasting future object states using a dynamics model, and rendering video frames from the predicted object representations.
-Various approaches have addressed this task using different architectural priors, such as RNNs (Creswell et al.
-2021; Nguyen et al.
-2024) or transformers (Wu, Yoon, and Ahn 2021; Wu et al.
-2023; Villar-Corrales, Wahdan, and Behnke 2023).
-Despite promising results, these models are limited to simple deterministic datasets or rely on action-conditioning (Mosbach et al.
-2025; Villar-Corrales and Behnke 2025).
+These methods typically involve three main steps: parsing seed frames into object representations, predicting future object states using a dynamics model, and rendering video frames from the predicted object representations.
+Various approaches have addressed this task using different architectural priors, such as RNNs (Creswell et al., 2021; Nguyen et al., 2024) or transformers (Wu et al., 2021; 2023; Villar-Corrales et al., 2023; Daniel and Tamar, 2024).
+Despite promising results, these models are limited to simple deterministic datasets or rely on action-conditioning (Mosbach et al., 2025; Villar-Corrales and Behnke, 2025).
 In contrast, our model forecasts future frames conditioned on past object slots and text descriptions.
-Text-Conditioned Video Prediction
+2.2.2 Text-Conditioned Video Prediction
 Text-conditioned VP models leverage text descriptions to provide appearance, motion and action cues that guide the generation of future frames.
-This task was first proposed by Hu, Luo, and Chen (2022), who utilized a VQ-VAE to encode images into visual token representations, and modeled the scene dynamics with an axial transformer to jointly process visual tokens with text descriptions.
-Similarly, approaches like TVP (Song et al.
-2024) and MMVG (Fu et al.
-2023) address this task using RNNs or masked transformers, respectively.
-More recently, several methods leverage diffusion models for text-guided VP (Blattmann et al.
-2023; Ni et al.
-2023; Chen et al.
-2024; Xing et al.
-2024; Chen et al.
-2023; Gu et al.
-2024).
+This task was first proposed by Hu et al.
+(2022), who utilized a VQ-VAE to encode images into visual token representations, and modeled the scene dynamics with an axial transformer to jointly process visual tokens with text descriptions.
+Similarly, approaches like TVP (Song et al., 2024) and MMVG (Fu et al., 2023) address this task using RNNs or masked transformers, respectively.
+More recently, several methods leverage diffusion models for text-guided VP (Blattmann et al., 2023; Ni et al., 2023; Chen et al., 2024; Xing et al., 2024; Chen et al., 2023; Gu et al., 2024).
 These approaches encode video frames into discrete token sequences via pretrained quantized autoencoders, and leverage pretrained language models to guide the diffusion-based generation with text features.
-To further improve temporal coherence and semantic alignment, several diffusion-based methods leverage specialized conditioning strategies (Xing et al.
-2024; Chen et al.
-2023) or attention mechanisms (Gu et al.
-2024), among others.
+To further improve temporal coherence and semantic alignment, several diffusion-based methods leverage specialized conditioning strategies (Xing et al., 2024; Chen et al., 2023) or attention mechanisms (Gu et al., 2024), among others.
 While effective, these models operate on holistic or spatial representations, and require large-scale compute and data.
@@
 (2025) and Jeong et al.
-(2025) combine object-centric learning with autoregressive diffusion models and transformers for text-guided video prediction on simple synthetic datasets.
-In contrast, we model object dynamics with an autoregressive transformer, evaluate on more complex robotics simulations, and perform an in-depth analysis of the properties of object-centric representations for text-conditioned video prediction.
+(2025) also combine object-centric learning with autoregressive diffusion models and transformers for text-guided video prediction on simple synthetic datasets.
+In contrast, we evaluate on more complex robotic simulations, and perform an in-depth analysis of the properties of object-centric representations for text-conditioned video prediction.
 3 Method
@@
 We propose two different TextOCVP variants, which differ in the underlying object-centric decomposition modules.
-Specifically, $\text{TextOCVP}_{\text{SAVi}}$ leverages SAVi (Kipf et al.
-2022), whereas $\text{TextOCVP}_{\text{DINO}}$ extends the recent DINOSAUR (Seitzer et al.
-2023) framework for recursive object-centric video decomposition and video rendering.
+Specifically, $\text{TextOCVP}_{\text{SAVi}}$ leverages SAVi (Kipf et al., 2022), whereas $\text{TextOCVP}_{\text{DINO}}$ extends DINOSAUR (Seitzer et al., 2023) for recursive object-centric video decomposition and video rendering.
+[…]
+Figure 2: Overview of TextOCVP.
+Our model parses the reference frame $\textbf{X}_{1}$ into object representations $\textbf{S}_{1}$.
+The text-conditioned object-centric predictor models object dynamics, incorporating information from the description $\mathbf{C}$ to predict future object states $\hat{\mathbf{S}}_{2:T+1}$, which can be decoded into video frames $\hat{\mathbf{X}}_{2:{T+1}}$.
 3.1 Scene Parsing
-The scene parsing module decomposes a video sequence $\mathbf{X}_{1:{\tau}}$ into a set of permutation-invariant object representations called slots $\mathbf{S}_{1:\tau}=(\textbf{S}_{1},...,\textbf{S}_{\tau})$, with $\textbf{S}_{t}=(\textbf{s}_{t}^{1},...,\textbf{s}_{t}^{N_{\mathbf{S}}})$, where each slot $\textbf{s}\in{\mathbb{R}}^{D}$ represents a single object.
+The scene parsing module decomposes a video sequence $\mathbf{X}_{1:{\tau}}$ into a set of permutation-invariant object representations called slots $\mathbf{S}_{1:\tau}=(\textbf{S}_{1},\ldots,\textbf{S}_{\tau})$, with $\textbf{S}_{t}=(\textbf{s}_{t}^{1},\ldots,\textbf{s}_{t}^{N_{\mathbf{S}}})$, where each slot $\textbf{s}\in{\mathbb{R}}^{D}$ represents a single object.
 For scene parsing, we adopt the recursive object-centric video decomposition framework from Kipf et al.
@@
 At time step $t$, the corresponding input frame $\textbf{X}_{t}$ is encoded with a feature extractor module into a set of $D_{h}$-dimensional feature maps $\textbf{h}_{t}\in{\mathbb{R}}^{L\times D_{h}}$ representing $L$ spatial locations.
-The feature extractor is a convolutional neural network in our $\text{TextOCVP}_{\text{SAVi}}$ variant and a DINO-pretrained vision transformer (Caron et al.
-2021) in $\text{TextOCVP}_{\text{DINO}}$.
-These feature maps are processed with Slot Attention (Locatello et al.
-2020), which updates the previous slots $\textbf{S}_{t-1}$ based on visual features from the current frame following an iterative attention mechanism.
+The feature extractor is a convolutional neural network in our $\text{TextOCVP}_{\text{SAVi}}$ variant and a DINO-pretrained vision transformer (Caron et al., 2021) in $\text{TextOCVP}_{\text{DINO}}$.
+These feature maps are processed with Slot Attention (Locatello et al., 2020), which updates the previous slots $\textbf{S}_{t-1}$ based on visual features from the current frame following an iterative attention mechanism.
 Namely, Slot Attention performs cross-attention, with the attention weights normalized over the slot dimension, thus encouraging the slots to compete to represent parts of the input.
-It then updates the slots using a Gated Recurrent Unit (Cho et al.
-2014) (GRU).
+It then updates the slots using a Gated Recurrent Unit (Cho et al., 2014) (GRU).
 Formally, Slot Attention updates the previous slots $\textbf{S}_{t-1}$ by:
@@
 The output of this module is a set of slots $\textbf{S}_{t}$ that represents the objects of the input frame.
-Figure 2: Overview of TextOCVP.
-TextOCVP parses the reference frame $\textbf{X}_{1}$ into object representations $\textbf{S}_{1}$.
-The text-conditioned object-centric predictor models object dynamics and interactions, incorporating information from the description $\mathbf{C}$ to predict future object states $\hat{\mathbf{S}}_{2:T+1}$, which can be used to render future video frames $\hat{\mathbf{X}}_{2:{T+1}}$.
 3.2 Text-Conditioned Object-Centric Predictor
-Our proposed text-conditioned predictor module, depicted in Fig.
-3, autoregressively forecasts future object states conditioned on the object slots from the reference frame $\textbf{S}_{1}$ and a textual description $\mathcal{C}$.
-To condition the prediction process, the text description $\mathcal{C}$ is encoded into text token embeddings $\mathbf{C}$ employing an encoder-only transformer module.
-We experiment with different variants of this module, including a vanilla transformer encoder (Vaswani et al.
-2017) or a pretrained T5 (Raffel et al.
-2020) module.
+Our proposed text-conditioned predictor module, illustrated in Fig.
+3, autoregressively forecasts future object states conditioned on the object slots from the reference frame $\textbf{S}_{1}$ and a text description $\mathcal{C}$.
+This design enables TextOCVP to predict temporally consistent object dynamics that are grounded not only in visual context, but also in high-level semantic intent specified via natural language.
+To condition the prediction process, the text description $\mathcal{C}$ is encoded into text token embeddings $\mathbf{C}$ using an encoder-only transformer.
+These tokens provide a global semantic prior guiding predictions at every time step.
+We experiment with different encoder variants, including vanilla transformer encoder (Vaswani et al., 2017) and pretrained T5 (Raffel et al., 2020).
 At time step $t$, the predictor receives as input the corresponding text embeddings $\mathbf{C}$, as well as the previous object slots $\mathbf{S}_{1:t}$, which are initially mapped via an MLP to the predictor token dimensionality.
 Additionally, these tokens are augmented with a temporal positional encoding, which applies the same sinusoidal positional embedding to all tokens from the same time step, thus preserving the inherent permutation-equivariance of the objects.
-Each layer of our predictor module mirrors the transformer decoder architecture (Vaswani et al.
-2017).
+[…]
+Figure 3: Overview of TextOCVP’s text-conditioned object-centric video predictor transformer.
+Each layer of our predictor module mirrors the transformer decoder architecture (Vaswani et al., 2017).
 First, a self-attention layer enables every slot to attend to all other object representations in the sequence, modeling the spatio-temporal relations between objects.
@@
 The video rendering module decodes the predicted slots $\hat{\textbf{S}}_{t}$ to render the corresponding video frame $\hat{\textbf{X}}_{t}$.
-We leverage two variants of the video rendering module, for our $\text{TextOCVP}_{\text{SAVi}}$ and $\text{TextOCVP}_{\text{DINO}}$ variants, respectively.
+We leverage two variants of the video rendering module, for our $\text{TextOCVP}_{\text{SAVi}}$ and $\text{TextOCVP}_{\text{DINO}}$ variants.
 $\text{TextOCVP}_{\text{SAVi}}$ Decoder
-This variant independently decodes each slot in $\hat{\textbf{S}}_{t}$ with a CNN-based Spatial Broadcast Decoder (Watters et al.
-2019), rendering an object image $\textbf{o}_{t}^{n}$ and mask $\textbf{m}_{t}^{n}$ for each slot $\textbf{s}_{t}^{n}$.
+This variant independently decodes each slot in $\hat{\textbf{S}}_{t}$ with a CNN-based Spatial Broadcast Decoder (Watters et al., 2019), rendering an object image $\textbf{o}_{t}^{n}$ and mask $\textbf{m}_{t}^{n}$ for each slot $\textbf{s}_{t}^{n}$.
 The object masks are normalized across the slot dimension, and the representations are combined via a weighted sum to render video frames:
 $\hat{\textbf{X}}_{t}=\sum_{n=1}^{N_{\mathbf{S}}}\textbf{o}_{t}^{n}\cdot\tilde{\textbf{m}}_{t}^{n}\;\;\;\text{with}\;\;\;\tilde{\textbf{m}}_{t}^{n}=\operatorname*{softmax}_{N_{\mathbf{S}}}(\textbf{m}_{t}^{n}).$ (3)
-Figure 3: Overview of our text-conditioned object-centric video predictor.
-Our predictor employs a transformer-based architecture to model the spatio-temporal dynamics of object slots via self-attention, while incorporating textual guidance through a text-to-slot cross-attention mechanism.
 $\text{TextOCVP}_{\text{DINO}}$ Decoder
 This decoder variant decodes the object slots in two distinct stages.
-First, following DINOSAUR (Seitzer et al.
-2023), an MLP-based Spatial Broadcast Decoder (Watters et al.
-2019) is used to generate object features along with their corresponding masks.
+First, following DINOSAUR (Seitzer et al., 2023), an MLP-based Spatial Broadcast Decoder (Watters et al., 2019) is used to generate object features along with their corresponding masks.
 Similar to the $\text{TextOCVP}_{\text{SAVi}}$ decoder, the object masks are normalized and combined with the object features in order to reconstruct the encoded features $\hat{\textbf{h}}_{t}\in{\mathbb{R}}^{L\times D_{h}}$.
@@
 […]
-[…]
 Predictor Training
 Given the pretrained scene parsing and rendering modules, we train our TextOCVP predictor for text-conditioned video prediction using a dataset containing paired videos and text descriptions.
-Namely, given the object representations from a reference frame $\textbf{S}_{1}$ and the textual embeddings $\mathbf{C}$, our TextOCVP predictor forecasts the subsequent object slots $\hat{\textbf{S}}_{2}$, which are decoded into a predicted video frame $\hat{\textbf{X}}_{2}$.
+Namely, given the object representations from a reference frame $\textbf{S}_{1}$ and the textual embeddings $\mathbf{C}$, the predictor forecasts subsequent object slots $\hat{\textbf{S}}_{2}$, which are decoded into a predicted video frame $\hat{\textbf{X}}_{2}$.
 This process is repeated autoregressively, i.e.
@@
 This autoregressive training, in contrast to teacher forcing, enforces our predictor to operate with imperfect inputs, leading to better modeling of long-term dynamics at inference time.
-Our predictor is trained by minimizing the following loss:
+TextOCVP predictor is trained by minimizing the following loss:
 $\mathcal{L}_{\text{TextOCVP}}=\frac{1}{T}\sum_{t=2}^{T+1}\left(\lambda_{\text{Img}}\mathcal{L}_{\text{Img}}+\lambda_{\text{Slot}}\mathcal{L}_{\text{Slot}}\right),$ (5)
@@
 4.1 Experimental Setup
-Datasets
+4.1.1 Datasets
 We evaluate TextOCVP for text-conditioned video prediction on the CATER and CLIPort datasets:
-CATER (Girdhar and Ramanan 2020) is a dataset of long video sequences showing 3D objects in motion, each accompanied by a descriptive caption.
-We used the CATER-hard variant (Hu, Luo, and Chen 2022), containing 30,000 sequences with $64\times 64$ frames featuring between two and eight objects, two of which follow the scripted motion.
-CLIPort (Shridhar, Manuelli, and Fox 2022) is a robot manipulation dataset featuring video-caption pairs.
-We employ 21,000 $336\times 336$ sequences of the Put-Block-In-Bowl task, where the caption describes placing a block into a bowl.
-Baselines
+CATER (Girdhar and Ramanan, 2020) is a dataset of long video sequences showing 3D objects in motion, each paired with a descriptive caption.
+We used the CATER-hard variant (Hu et al., 2022), containing 30,000 sequences with $64\times 64$ frames featuring two to eight objects, two of which follow the scripted motion.
+CLIPort (Shridhar et al., 2022) is a robot manipulation dataset featuring video-caption pairs.
+We employ 21,000 $336\times 336$ sequences of the Put-Block-In-Bowl task, where each scene contains six objects, either a block or a bowl, on a 2D table plane, and a robot arm.
+The caption describes placing a block into a bowl.
+4.1.2 Baselines
 […]
-Refer to captiont=1t=122101020203030MAGE TextOCVP
-‘the medium green metal sphere is sliding
-to (2, 1). the small brown metal cube
-is picked up and placed to (-3, 1)’
-Figure 4: Text-guided video prediction result on CATER. Top row shows ground truth. TextOCVP predicts sharp and accurate frames, whereas MAGE blurs and misses objects.
-Implementation Details
+4.1.3 Implementation Details
 All models are implemented in PyTorch and trained on a single NVIDIA A6000 (48Gb) GPU.
@@
 (2022) for scene parsing and video rendering.
-$\text{TextOCVP}_{\text{DINO}}$ uses DINOv2 (Oquab et al.
-2024) as image encoder, a four-layer MLP Spatial-Broadcast Decoder (Watters et al.
-2019) to decode the slots into object features and masks, and a CNN decoder to map the reconstructed scene features back to images.
+$\text{TextOCVP}_{\text{DINO}}$ uses DINOv2 (Oquab et al., 2024) as image encoder, a four-layer MLP Spatial-Broadcast Decoder (Watters et al., 2019) to decode slots into object features and masks, and a CNN decoder to map the reconstructed scene features back to images.
 On CATER, we use the $\text{TextOCVP}_{\text{SAVi}}$ variant with eight 128-dimensional object slots, whereas on CLIPort we employ the $\text{TextOCVP}_{\text{DINO}}$ variant with ten 128-dimensional slots.
 Our predictor module is an eight-layer transformer with 512-dimensional tokens, eight attention heads, and a hidden dimension of 1024.
-Further experiment details are provided in Appendix B-E.
+Further experimental details—including datasets, baselines, evaluation metrics, and implementation details—are provided in Appendices B-E.
 4.2 Results
-CATER Results
+4.2.1 CATER Results
+Table 1: Quantitative evaluation on CATER and CLIPort datasets for prediction horizons of $T=9$ and $T=19$.
+TextOCVP outperforms the baselines.
+Best two results are shown in bold and underlined, respectively.
+(a) Quantitative evaluation on CATER.
+CATER1→9 CATER1→19
+[…]
+OCVP (Villar-Corrales et al., 2023) 29.08 0.874 0.078 4.16 28.11 0.854 0.101 8.08
+Non-OC 29.68 0.874 0.092 3.04 28.39 0.849 0.112 8.62
+SEER (Gu et al., 2024) 22.05 0.723 0.245 11.23 16.05 0.535 0.299 17.29
+MAGE (Hu et al., 2022) 34.91 0.877 0.108 3.46 34.76 0.871 0.111 5.88
+TextOCVP (Ours) 32.98 0.922 0.036 2.16 31.29 0.902 0.044 5.09
+(b) Quantitative evaluation on CLIPort.
+CLIPort1→9 CLIPort1→19
+[…]
+Non-OC 23.44 0.901 0.184 8.13 20.14 0.872 0.210 13.23
+SEER (Gu et al., 2024) 21.01 0.887 0.141 6.80 11.30 0.622 0.331 8.29
+$\text{MAGE}_{\text{DINO}}$ (Hu et al., 2022) 23.72 0.940 0.064 2.11 22.27 0.931 0.075 2.59
+TextOCVP (Ours) 26.99 0.950 0.062 1.36 23.88 0.931 0.078 2.23
 On CATER, we train the models to predict nine future frames given a single reference frame and a text caption.
@@
 Fig.
-4 qualitatively compares TextOCVP and MAGE.
-Our proposed method generates a sequence that is closely aligned to the ground-truth, whereas MAGE predictions feature multiple errors and artifacts, including missing objects, blurry contours, and significant changes on object shapes.
-This highlights the effectiveness of object-centric representations to accurately represent and model object dynamics.
+4(a) qualitatively compares TextOCVP and MAGE on CATER.
+TextOCVP generates a sequence that exhibits coherent object trajectories and accurate motion grounded in the input textual instruction.
+This alignment is made possible by the explicit object-centric design, which disentangles the scene into structured object representations and allows the model to reason about each object’s dynamics independently.
+In contrast, MAGE predictions feature multiple errors and artifacts, including missing objects, blurry contours, and significant changes on object shapes.
+These issues arise from the difficulty of modeling object-specific dynamics in pixel or token space without explicit object-level abstraction.
+These results demonstrate the importance of structured, object-centric representations in capturing the compositional and controllable nature of physical scenes, which is especially critical in tasks involving fine-grained object manipulations.
+Refer to captiont=1t=122101020203030MAGETextOCVP‘the medium green metal sphere is sliding to (2, 1).the small brown metal cube is picked up and placed to (-3, 1)BlurredChanged
+(a) Text-guided video prediction result on CATER. Top row depicts the ground truth frames. TextOCVP predicts sharp and accurate frames, whereas MAGE blurs and misses objects.
 […]
-(a) CATER Evaluation
-CATER1→9 CATER1→19
+(b) Qualitative evaluation on CLIPort for text-driven video prediction.
+Top row depicts the ground truth frames.
+TextOCVP successfully completes the pick-and-place task, whereas in $\text{MAGE}_{\text{DINO}}$ the moved block disappears.
+Figure 4: Qualitative comparison of TextOCVP and baseline methods on text-guided video prediction.
+4.2.2 CLIPort Results
 […]
-OCVP 0.874 0.078 4.16 0.854 0.101 8.08
-Non-OC 0.874 0.092 3.04 0.849 0.112 8.62
-SEER 0.723 0.245 11.23 0.535 0.299 17.29
-MAGE 0.877 0.108 3.46 0.871 0.111 5.88
-TextOCVP 0.922 0.036 2.16 0.902 0.044 5.09
-(b) CLIPort Evaluation
-CLIPort1→9 CLIPort1→19
 […]
-Non-OC 0.901 0.184 8.13 0.872 0.210 13.23
-SEER 0.887 0.141 6.80 0.622 0.331 8.29
-$\text{MAGE}_{\text{DINO}}$ 0.940 0.064 2.11 0.931 0.075 2.59
-TextOCVP 0.950 0.062 1.36 0.931 0.078 2.23
-Table 2: Ablation studies. We ablate different architectural design choices in our proposed TextOCVP model.
-(a) Effect of number of layers in predictor.
+Nevertheless, we observe that TextOCVP often generates visual artifacts and lacks fine-grained textures.
+While these imperfections can affect its quantitative scores, TextOCVP’s object-centric structure enables robust long-term prediction and precise instruction-following behavior.
+4.3 Model Analysis
+4.3.1 Ablation Studies
+We perform several ablation studies to support and validate the architectural choices of our model components and their impact on TextOCVP’s video prediction performance.
+The results are presented in Table 2, and analysed below across four main design axes.
+Further results and analysis are provided in Appendix F.
+Table 2: Ablation studies on key design choices in TextOCVP, including predictor depth (2(a)), text encoder choice (2(b)), residual connections (2(c)), and slot count (2(d)).
+(a) Effect of number of layers ($N_{\text{P}}$) in predictor module.
 CATER1→9
@@
 8 0.922 0.036 2.16
-(b) Evaluation of different text encoders on CATER and CLIPort.
+(b) Impact of different text encoders, including a transformer encoder (TF) and two T5 variants, on CATER and CLIPort.
 CATER1→19 CLIPort1→19
 […]
-Custom TF 0.903 0.045 5.39 - - -
+TF 0.903 0.045 5.39 - - -
 Frozen T5 0.902 0.044 5.09 0.931 0.078 2.23
@@
 […]
-No 0.946 0.066 1.58
-Yes 0.950 0.062 1.36
-(d) Effect of the number of slots $N_{\mathbf{S}}$.
+0.946 0.066 1.58
+0.950 0.062 1.36
+(d) Effect of the number of object slots ($N_{\mathbf{S}}$).
 CLIPort1→9
@@
 10 0.950 0.062 1.36
-CLIPort Results
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-‘put the blue block in the gray bowl
-[…]
-TextOCVPt=1t=12020303040405050Missing
-[…]
-[…]
-[…]
-Nevertheless, we identify that TextOCVP often generates visual artifacts and lacks fine-grained textures, which can affect its quantitative scores despite accurate predictions.
-Further figures and results are provided in Appendix F.
-4.3 Model Analysis
-Ablation Studies
-We perform several ablation studies to support and validate the architectural choices of our model components and their impact on TextOCVP’s video prediction performance.
-The results are presented in Table 2(d).
-Number of Layers (Table 2(a))   TextOCVP performs best with $N_{\text{P}}=8$ predictor layers, outperforming shallower variants.
-We did not explore scaling beyond eight layers due to the significant increase in the number of parameters.
-Residual Connection (Table 2(c))   Incorporating a residual connection in the predictor, $\hat{\textbf{S}}_{t+1}:=\hat{\textbf{S}}_{t+1}+\textbf{S}_{t}$, enhances the model’s performance by improving the temporal consistency of the predicted slots.
-Text Encoder (Table 2(b))   We evaluate the performance of different text encoders, including a vanilla transformer, a frozen T5 encoder, and a fine-tuned (FT) T5.
-A frozen T5 module led to the best TextOCVP performance on the CLIPort dataset.
-Although all encoders led to similar performance on the CATER dataset, we opted for the frozen T5 to avoid unnecessarily increasing the model’s parameter count.
-Number of Slots (Table 2(d))   Using $N_{\mathbf{S}}=10$ slots resulted in better performance for TextOCVP on CLIPort, even though each scene can be described with eight slots (six objects, robot arm and background).
-This observation suggests that additional slots can be used as registers in order to help with internal computations (Darcet et al.
-2024).
-Table 3: Quantitative evaluation on a CLIPort test set with unseen colors. We report the video prediction performance and its drop relative to the known-color dataset.
+Number of Layers (Table 2(a))
+We study the effect of increasing the number of transformer layers ($N_{\text{P}}$) in the predictor module.
+We observe that prediction quality improves with model depth, with $N_{\text{P}}=8$ yielding the best performance across all metrics on the CATER dataset.
+This suggests that deeper predictor models enable more accurate modeling of object dynamics.
+However, we did not explore beyond eight layers due to the substantial increase in computational cost, training time and parameter count.
+Residual Connection (Table 2(c))
+We evaluate the effect of introducing a residual connection in the predictor module, where the updated slot representation is defined as $\hat{\textbf{S}}_{t+1}=\textbf{S}_{t}+f_{\text{pred}}(\textbf{S}_{t})$.
+Adding this residual path improves performance across all metrics on the CLIPort dataset, particularly in improving the temporal consistency of predicted slots.
+This result is consistent with prior findings (Villar-Corrales et al., 2023), which show that residual updates are beneficial for iterative refinement in structured latent spaces.
+Text Encoder (Table 2(b))
+We evaluate the performance of three different text encoders, including a lightweight transformer trained from scratch (TF), a frozen T5 encoder, and a fine-tuned (FT) T5.
+While all three perform similarly on CATER, the frozen T5 text encoder achieves the best results on CLIPort, especially on perceptual metrics such as JEDi and LPIPS.
+Interestingly, fine-tuning the T5 encoder led to slightly degraded performance, likely due to overfitting or interference with the pretrained representations.
+We therefore use the frozen T5 encoder in all main experiments to balance performance and model efficiency.
+Number of Slots (Table 2(d))
+Finally, we explore the impact of varying the number of slots ($N_{\mathbf{S}}$) used to represent the scene.
+Although each CLIPort scene can be described with eight slots—representing six objects, the robot arm, and the background—we find that using ten slots results in notably better performance.
+This observation suggests that the additional slots can function as internal registers, supporting attention routing or acting as a form of cache that aids with internal model computations (Darcet et al., 2024).
+4.3.2 Model Robustness and Generalization
+We evaluate the robustness and generalization of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on two CLIPort evaluation sets: one involving color variations in the text instructions that were not encountered during training (unseen-color), and another featuring scenes with more objects than observed during training, eight instead of six (more-objects).
+These experiments assess model robustness under distribution shifts and test the ability to generalize to unseen visual-linguistic combinations and novel scene configurations.
+Results are summarized in Table 3, reporting both the absolute performance and the relative drop compared to the standard (seen-color or six-object) settings.
+On the unseen-color benchmark (Table 3(a)), TextOCVP consistently outperforms $\text{MAGE}_{\text{DINO}}$ across all metrics and prediction horizons, showing significantly smaller degradation in perceptual similarity (LPIPS).
+This indicates stronger robustness to novel color-object combinations.
+In the more-objects setting (Table 3(b)), TextOCVP again maintains superior performance with a notably smaller decline in LPIPS, demonstrating resilience to higher scene complexity and unseen object counts.
+Qualitative examples in Appendix F.6 further support these findings.
+Overall, these results highlight the advantages of structured object-centric representations, which provide greater generalization and robustness in video prediction tasks, particularly in contrast to holistic scene representations that struggle with novel scene compositions.
+Table 3: Quantitative evaluation on two CLIPort test sets with unseen colors and larger number of objects.
+We report the absolute video prediction performance and the relative drop (in parentheses) compared to the original evaluation setting.
+Best result is marked in bold.
+(a) Evaluation on CLIPort test set with objects of colors unseen during training.
 CLIPort1→9 CLIPort1→19
 […]
-$\text{MAGE}_{\text{DINO}}$ 0.935 -0.5% 0.076 -19% 0.924 -0.7% 0.087 -16%
-TextOCVP 0.946-0.4% 0.066-6.4% 0.927-0.4% 0.083-6.4%
-Model Robustness
-‘the large purple rubber cone is picked up
-and placed to (2, 3). the small gold metal
-snitch is picked up and placed to (-1, 1).’
-Original
-Caption
-a)
-‘the medium cyan rubber sphere is picked up
-and placed to (-2, 2). the medium purple
-metal cone is sliding to (2, 3).’
-Changed
-Objects
-& Actions
-b)
-‘the large purple rubber cone is sliding to
-(-1, 1). the small gold metal snitch is
-rotating. the medium purple metal cone
-is picked up and placed to (-1, -3).’
-More
-Moving
-Objects
-c)
-t=1t=12215153030
-Figure 6: Evaluation of TextOCVP’s controllability on CATER.
-Top row shows the ground truth sequence.
-TextOCVP enables predicting different sequence continuations conditioned on distinct text instructions.
-We evaluate the generalization ability of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set featuring color variations in the text instructions that were not encountered during training.
-This setup assesses the models’ ability to generalize to novel scene configurations and adapt to unseen text-visual combinations.
-Table 3 presents the results on this evaluation set with unseen colors.
-We report both the video prediction performance as well as the performance drop relative to the evaluation set with known colors.
-Our model consistently outperforms $\text{MAGE}_{\text{DINO}}$ for both prediction horizons.
-Most notably, TextOCVP demonstrates significantly higher robustness to objects with previously unseen colors compared to the baseline, showing a much smaller drop in performance for the perceptual LPIPS metric.
-These results highlight the advantages of structured object-centric representations, which provide greater generalization and robustness in video prediction—particularly in contrast to holistic scene representations that struggle with novel scene compositions.
-Controllability
-A key objective of text-driven video prediction is to provide fine-grained control over the prediction process, which is enabled through language instructions describing the objects in the scene and their expected motion.
-In Fig.
-6, we qualitatively showcase the level of control that TextOCVP provides over its generations on the CATER dataset.
-We qualitatively evaluate this ability by generating multiple video continuations from the same reference frame while varying the language instruction.
-These variations include changes to both the target objects and their actions, as well as providing instructions that include a number of actions beyond what the model was trained on.
-As shown in Fig.
-6, TextOCVP successfully identifies the relevant objects and executes the actions described in the text instructions.
-Notably, our model distinguishes between two nearly identical purple cones, despite the identical shapes and color, and correctly generates a sequence with the specified motion.
-These results show that key object attributes, such as size or color, are effectively captured through the text-to-slot attention mechanism, enabling the predictor to accurately forecast the motion of each individual object.
-This highlights the benefits of combining object-centric representations with text guidance for fine-grained, controllable video prediction.
-Interpretability
-Refer to captiont=1t=1Refer to caption55Refer to caption1010Refer to caption2020Refer to caption3030
-‘the cone is sliding to (-3, -3).
-the snitch is picked up and placed to (-1, 2 ).’
-Refer to captionRefer to caption1.00.50.0Refer to caption
+$\text{MAGE}_{\text{DINO}}$ (Hu et al., 2022) 0.935 (-0.5%) 0.076 (-19%) 0.924 (-0.7%) 0.087 (-16%)
+TextOCVP (Ours) 0.946(-0.4%) 0.066(-6.4%) 0.927(-0.4%) 0.083(-6.4%)
+(b) Evaluation on CLIPort test set with a larger object count than training scenes.
+CLIPort1→9 CLIPort1→19
+[…]
+$\text{MAGE}_{\text{DINO}}$ (Hu et al., 2022) 0.929 (-1.2%) 0.088 (-37.5%) 0.920 (-1.2%) 0.094 (-25.3%)
+TextOCVP (Ours) 0.936 (-1.5%) 0.076(-22.6%) 0.921(-1.1%) 0.090(-15.4%)
+4.3.3 Interpretability
+Refer to captiont=1t=1Refer to caption55Refer to caption1010Refer to caption1515Refer to caption2020Refer to caption2525Refer to caption3030‘the cone is sliding to (-3, -3).
+the snitch is picked up and placed to (-1, 2).’ Refer to captionRefer to caption1.00.50.0Refer to caption
 Slot 3
@@
 2
-Figure 7: Text-to-slot attention weights for different objects in a video. Slots representing objects described in the caption focus on relevant text tokens, such as target locations.
-In Fig.
-7, we visualize the text-to-slot attention weights—averaged across attention heads—for different slots in a CATER sequence, revealing how textual information guides model predictions.
-We observe that slots corresponding to objects described in the textual instruction focus on relevant text tokens, such as target coordinate locations.
-These results demonstrate that the text-to-slot attention mechanism effectively aligns the textual information with the object-centric representations, enabling the model to generate accurate video sequences based on the given text instructions, while providing superior interpretability.
-5 Conclusion
-We presented TextOCVP, a novel object-centric model for text-conditioned video prediction.
-Given a single input image and a text description, TextOCVP generates frames by parsing the scene into object representations and modeling their dynamics conditioned on the textual instruction.
-This is accomplished through a text-conditioned object-centric transformer, which predicts future object states by modeling the spatio-temporal relationships between objects while incorporating textual guidance.
-Through extensive evaluations, we demonstrated that TextOCVP outperforms existing models for text-driven video prediction from a single frame, highlighting our model’s ability to predict over long prediction horizons and adapt its predictions based on the provided textual description.
-Moreover, we validated our architectural choices through ablation studies, highlighting the importance of combining textual and object-centric information, as well as demonstrating TextOCVP’s robustness and interpretability.
-With its structured latent space and superior controllability, TextOCVP offers potential for applications in robotics, where it can serve as a controllable world model, enabling efficient planning, reasoning, and decision-making.
-Acknowledgment
-This work was funded by grant BE 2556/16-2 (Research Unit FOR 2535 Anticipating Human Behavior) of the German Research Foundation (DFG)
-Appendix A Limitations and Future Work
-A.1 Limitations
-While TextOCVP demonstrates promising results for text-guided object-centric video prediction, it presents some limitations that we plan to address in future work:
-Prediction Artifacts
-TextOCVP occasionally generates artifacts in the predicted frames, such as blurriness, inconsistent object appearances, lack of textured details, or visual artifacts in the background.
-We believe that these limitations stem from the video rendering module, which might lack the representational power to reconstruct precise image details from the object-centric latent space representation.
-Limited Temporal Consistency
-While TextOCVP produces plausible predictions, we observe that its temporal consistency can degrade when forecasting for long prediction horizons ($T>30$), occasionally resulting in object jittering or instability.
-We attribute this limitation to the fact that TextOCVP is trained to predict only up to nine future frames ($T=9$) and only optimizing reconstruction losses, which do not penalize such temporal inconsistencies.
-A.2 Future Work
-In future work we aim to address these limitations.
-Our TextOCVP model is designed with a modular architecture, enabling for both improvements and flexible swapping of components—such as the parsing, predictor, or video rendering modules—to seamlessly improve the entire pipeline.
-To address the prediction artifacts, we plan to extend our TextOCVP framework with more powerful decoder modules, such as autoregressive transformers (Singh, Wu, and Ahn 2022) or diffusion models (Jiang et al.
-2023), as well as scale our predictor module.
-Furthermore, we plan to incorporate temporal discriminators (Clark, Donahue, and Simonyan 2020) to improve the temporal consistency of the predicted video frames.
-We believe that exploring these architectural modifications can overcome the aforementioned limitations and will enable the scaling of TextOCVP to real-world robotic environments.
-Appendix B Implementation Details
-We employ $\text{TextOCVP}_{\text{SAVi}}$ for the experiments on CATER and $\text{TextOCVP}_{\text{DINO}}$ for experiments on CLIPort.
-Below we discuss the implementation details for each of these variants.
-B.1 $\text{TextOCVP}_{\text{DINO}}$
-The $\text{TextOCVP}_{\text{DINO}}$ variant consists of our proposed text-conditioned predictor module and an object-centric decomposition module that extends the DINOSAUR (Seitzer et al.
-2023) framework for recursive object-centric video decomposition and video rendering.
-Text-Conditioned Predictor
-The predictor is composed of $N_{\text{P}}=8$ identical layers, each containing 8-head attention mechanisms and an MLP with a single hidden layer of dimension 1024 and a ReLU activation function.
-Furthermore, the predictor uses an embedding dimensionality of 512, context window size of ten frames, and applies a residual connection from the predictor input to its output.
-Text Encoder
-$\text{TextOCVP}_{\text{DINO}}$ leverages a pretrained and frozen small version of T5 encoder (Raffel et al.
-2020), which consists of six T5 blocks.
-This text encoder uses a vocabulary with size 32,128.
-Scene Parsing
-The scene parsing module generates $N_{\mathbf{S}}=10$ slots of dimension 128.
-As feature extractor, we use DINOv2 ViT-Base (Oquab et al.
-2024), featuring 12 layers, using a patch size of 14, and producing patch features with dimension $D_{h}=768$.
-The Slot Attention corrector module processes the first video frame with three iterations in order to obtain a good initial object-centric decomposition, and a single iteration for subsequent frames, which suffices to recursively update the slot representation.
-The initial object slots $\textbf{S}_{0}$ are randomly sampled from a Gaussian distribution with learned mean and covariance.
-We use a single Transformer encoder block as the transition function, which consists of four attention heads and an MLP with a hidden dimension of 512.
-Video Rendering
-The video rendering module consists of two distinct decoders.
-First, a four-layer MLP-based Spatial Broadcast Decoder (Watters et al.
-2019) with hidden dimension 1024 reconstructs the patch features from the slots.
-Then, a CNN-based decoder reconstructs full-resolution images from these features.
-It consists of four convolutional layers, each using $3\times 3$ kernels, a ReLU activation function and bilinear upsampling.
-A final convolutional layer and bilinear interpolation are applied to map the outputs to the RGB channels and spatial dimensions of the image.
-Training
-We train our model for object-centric decomposition using video sequences of length five frames for 1000 epochs.
-We use batch size of 16, the Adam optimizer (Kingma and Ba 2015), and a base learning rate of $4\times 10^{-4}$, which is linearly warmed-up for the first 10000 steps, followed by cosine annealing for the remaining of the training process.
-Moreover, we clip the gradients to a maximum norm of 0.05.
-The predictor module is trained given the frozen and pretrained object-centric decomposition model for 700 epochs to predict the subsequent nine frames using a single seed frame.
-The predictor is trained using the same hyper-parameters as for object-centric decomposition.
-In the predictor loss function $\mathcal{L}_{\text{TextOCVP}}$, we set $\lambda_{\text{Img}}=1$ and $\lambda_{\text{Slot}}=1$.
-B.2 $\text{TextOCVP}_{\text{SAVi}}$
-$\text{TextOCVP}_{\text{SAVi}}$ uses the same text-conditioned predictor and text-encoder architectures as $\text{TextOCVP}_{\text{DINO}}$, but employs SAVi (Kipf et al.
-2022) as the object-centric decomposition module.
-Scene Parsing
-The scene parsing module generates $N_{\mathbf{S}}=8$ slots of dimension 128.
-Following Kipf et al.
-(2022), we use as feature extractor a four-layer CNN with ReLU activation function, where each convolutional layer features 32 $5\times 5$ kernels, $\text{stride}=1$, and $\text{padding}=2$.
-The Slot Attention corrector follows the same structure as in $\text{TextOCVP}_{\text{DINO}}$.
-Video Rendering
-Following Kipf et al.
-(2022), we utilize a CNN-based Spatial Broadcast Decoder (Watters et al.
-2019) with four convolutional layers with 32 kernels of size $5\times 5$ , $\text{stride}=1$, and $\text{padding}=2$.
-A final convolutional layer maps from the hidden 32-channel representation to four output channels (RGB + alpha mask).
-Training
-We train our model for object-centric decomposition using video sequences of length ten frames for 1000 epochs, using batch size of 64, and an initial learning rate of $10^{-4}$, which is warmed up for 2500 steps, followed by cosine annealing for the remaining of the training process.
-Moreover, we clip the gradients to a maximum norm of 0.05.
-The predictor module is trained given the frozen and pretrained object-centric decomposition model for 1400 epochs to predict the subsequent nine frames using a single seed frame.
-The predictor is trained using the same hyper-parameters as for object-centric decomposition.
-In the predictor loss function $\mathcal{L}_{\text{TextOCVP}}$, we set $\lambda_{\text{Img}}=1$ and $\lambda_{\text{Slot}}=1$.
-Table 4: Number of learnable parameters in TextOCVP and baselines for experiments on the CLIPort dataset.
-Model # Parameters
-TextOCVP 33.76M
-Non-OC 34.16M
-$\text{MAGE}_{\text{DINO}}$ 32.11M
-SEER 405.89M
-Appendix C Baselines
-[…]
-[…]
-Table 5: Comparison of MAGE variants on CLIPort. $\text{MAGE}_{\text{DINO}}$ clearly outperforms the original MAGE variant.
-Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
-MAGE 7.116 0.453 0.713
-$\text{MAGE}_{\text{DINO}}$ 23.723 0.940 0.064
-[…]
-MAGE is an autoregressive text-guided video prediction framework that utilizes a VQ-VAE (Van Den Oord, Vinyals et al.
-2017) encoder-decoder architecture to learn efficient visual token representations.
-A cross-attention module aligns textual and visual embeddings to produce a spatially-aligned motion representation termed Motion Anchor (MA), which is fused with visual tokens via an axial transformer for video generation.
-For experiments on CATER, we use a codebook size of $512\times 256$ with a downsampling ratio of four, whereas on CLIPort we use a codebook size of $512\times 1024$.
-To ensure a fair comparison with TextOCVP on CLIPort, we replace MAGE’s standard CNN encoder and decoder with the DINOv2 ViT encoder and CNN decoder used in our TextOCVP model.
-We refer to this modified version as $\text{MAGE}_{\text{DINO}}$.
-Table 5 presents a comparison between the original MAGE model and $\text{MAGE}_{\text{DINO}}$ on CLIPort.
-The results demonstrate that $\text{MAGE}_{\text{DINO}}$ significantly outperforms the original variant, enabling a fair comparison with TextOCVP and other baselines on this benchmark.
-C.2 SEER
-SEER is a diffusion-based model for language-guided video prediction.
-It employs an Inflated 3D U-Net derived from a pretrained text-to-image 2D latent diffusion model (Rombach et al.
-2022), extending it along the temporal axis and integrating temporal attention layers to simultaneously model spatial and temporal dynamics.
-For the language conditioning module, SEER introduces a novel Frame Sequential Text (FSText) Decomposer, which decomposes global instructions generated by the CLIP text encoder (Radford et al.
-2021) into frame-specific sub-instructions.
-These are aligned with frames using a transformer-based temporal network and injected into the diffusion process via cross-attention layers.
-We initialize SEER from a checkpoint pretrained on the Something-Something V2 dataset (Goyal et al.
-2017), and further fine-tune it for a few epochs.
-We observed that incorporating a text loss enhanced SEER’s performance, while other hyper-parameters were kept consistent with its original implementation.
-Appendix D Datasets
-D.1 CATER
-CATER (Girdhar and Ramanan 2020) is a dataset that consists of long video sequences, each described by a textual caption.
-The video scenes consist of multiple 3D geometric objects in a 2D table plane, which is split into a $6\times 6$ grid with fixed axis, allowing the exact description of object’s positions using coordinates.
-The text instruction describes the movement of specific objects through four atomic actions: ‘rotate’, ‘pick-place’, ‘slide’, and ‘contain’.
-The caption follows a template consisting of the subject, action, and an optional object or end-point coordinate, depending on the action.
-The movement of the objects starts at the same time step.
-Furthermore, the initial positions are randomly selected from the plane grid, and the camera position is fixed for every sequence.
-In our work, we employ CATER-hard, which is a complete version of the CATER dataset, containing 30000 video-caption pairs, with video frames resized to $64\times 64$.
-It includes 5 possible objects: cone, cube, sphere, cylinder, or snitch, which is a special small object in metallic gold color, shaped like three intertwined toruses.
-Furthermore, every object is described by its size (small, medium, or large), material (metal or rubber), and color (red, blue, green, yellow, gray, brown, purple, cyan, or gold if the object is the snitch), and this description is included in the textual caption.
-Every atomic action is available.
-The ‘rotate’ action is afforded by cubes, cylinders and the snitch, the ‘contain’ action is only afforded by the cones, while the other two actions are afforded by every object.
-Every video has between 3 and 8 objects, and two actions happen to different objects at the same time.
-The vocabulary size is 50.
-D.2 CLIPort
-CLIPort (Shridhar, Manuelli, and Fox 2022) is a robot manipulation dataset, consisting of video-caption pairs, i.e.
-long videos whose motion is described by a textual video caption.
-There are many variants of the CLIPort dataset, but we focus on the Put-Block-In-Bowl variant.
-We generate 21000 video-caption pairs with resolution $336\times 336$.
-Every video contains 6 objects on a 2D table plane, and a robot arm.
-Objects can be either a block or a bowl, and there is at least one of them in every sequence.
-The starting position of each object is random, with the only constraint being that it must be placed on the table.
-Each video describes the action of the robot arm picking a block, and putting it in a specific bowl.
-The video caption follows the template ‘put the [color] block in the [color] bowl’.
-Each individual object in the scene has a different color.
-In the train and validation set, the block and the bowl that are part of the caption can have one of the following colors: blue, green, red, brown, cyan, gray, or yellow, while in the evaluation set with unseen colors they can have blue, green, red, pink, purple, white, or orange color.
-The other 4 objects, called distractors, can have any color.
-During a video sequence, it can be possible that the robot arm goes out of frame, and comes back in later frames, thus requiring the model to leverage long range dependencies.
-The vocabulary size is 15.
-[…]
-CATER1→9 CATER1→19 CLIPort1→9 CLIPort1→19
-[…]
-OCVP 29.08 0.874 0.078 4.16 28.11 0.854 0.101 8.08
-Non-OC 29.68 0.874 0.092 3.04 28.39 0.849 0.112 8.62 23.44 0.901 0.184 8.13 20.14 0.872 0.210 13.23
-SEER 22.05 0.723 0.245 11.23 16.05 0.535 0.299 17.29 21.01 0.887 0.141 6.80 11.30 0.622 0.331 8.29
-MAGE 34.91 0.877 0.108 3.46 34.76 0.871 0.111 5.88 23.72 0.940 0.064 2.11 22.27 0.931 0.075 2.59
-TextOCVP 32.98 0.922 0.036 2.16 31.29 0.902 0.044 5.09 26.99 0.950 0.062 1.36 23.88 0.931 0.078 2.23
-Appendix E Evaluation Metrics
-To measure TextOCVP’s video prediction performance and compare it with existing approaches, we evaluate the visual quality of predicted video frames using popular image- and video-based metrics.
-PSNR and SSIM (Wang et al. 2004) measure pixel-wise and statistical differences between the predicted and ground-truth video frames, respectively.
-LPIPS (Zhang et al.
-2018) is a perceptual metric that measures the visual similarity between two images based on deep features from pretrained neural networks, usually VGG (Simonyan and Zisserman 2015).
-Unlike pixel-wise metrics, LPIPS compares activations from multiple layers and captures differences in texture, structure, and semantics, thus aligning closely with human perception.
-While these metrics focus on frame-level visual quality, video prediction evaluation requires quantifying temporal consistency and motion realism.
-JEDi (Luo et al.
-2025) measures the quality and realism of generated videos by comparing feature distributions of generated and real videos, capturing both visual fidelity and motion dynamics.
-We prefer JEDi over FVD (Unterthiner et al.
-2019), as JEDi is less sensitive to small evaluation datasets, and more stable in low-motion synthetic scenarios.
-In our evaluations and comparisons with baselines, we favor the LPIPS and JEDi scores, which correlate well with human perception, while reporting other metrics for completeness.
-Appendix F Additional Results
-F.1 SAVi vs. DINOSAUR
-Current object-centric approaches for video prediction are limited to relatively simple synthetic datasets, and struggle to scale beyond scenes featuring simple 3D shapes with simple deterministic motion.
-We attribute this limitation primarily to the object-centric modules used for learning object representations.
-Motivated by this observation, we extend the recent DINOSAUR (Seitzer et al.
-2023) framework to recursively process video data and reconstruct video frames from their corresponding object-centric representations.
-To demonstrate the significance of the object-centric module in scaling to more complex datasets, we compare both SAVi and our Extended DINOSAUR trained on CLIPort.
-As illustrated in Fig.
-8, SAVi struggles to accurately represent the objects on the table, missing multiple objects and changing their shape and color.
-In contrast, the Extended DINOSAUR model successfully reconstructs the scene, closely resembling the input, while accurately representing each object.
-The visual features extracted by the DINOv2 (Oquab et al.
-2024) encoder contain high-level semantic information, and during training, the slots are specifically optimized to efficiently encode this information.
-This design enables the model to scale and handle more complex object-centric video data effectively.
-$t=1$
-22
-33
-44
-55
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-GT
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-SAVi
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-Ext. DINO
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-Ext. DINO
-Masks
-Figure 8: Comparison between SAVi and our Extended DINOSAUR (Ext.
-DINO) decomposition modules for reconstructing a CLIPort sequence.
-We visualize the reconstructed frames, as well as the slot masks obtained by Extended DINOSAUR.
-SAVi fails to reconstruct most objects, whereas Extended DINOSAUR accurately reconstructs the scene, while representing each object.
-F.2 Quantitative Comparison
-In the main paper, we present quantitative evaluations on the CATER and CLIPort datasets using three evaluation metrics, shown in two separate tables.
-For completeness, we report in Table 6 a text-guided video prediction evaluation of TextOCVP and multiple baseline models on CATER and CLIPort using four distinct evaluation metrics for different prediction horizons.
-[…]
-F.3 Computational Efficiency
-In Table 7 we report the average inference time for $T=1$ and $T=9$ frame predictions on CATER and CLIPort using a NVIDIA-A6000-48Gb GPU.
-[…]
-Table 7: Average inference (Inf.) time for $T=1$ and $T=9$ frame predictions on CATER and CLIPort.
-CATER Inf. [ms] CLIPort Inf. [ms]
-Model $T=1$ $T=9$ $T=1$ $T=9$
-MAGE 15.4 $\pm$ 1 111.1 $\pm$ 11 34.4 $\pm$ 3 198.8 $\pm$ 2
-[…]
-Table 8: Quantitative evaluation on a CLIPort test set with more objects. We report the video generation performance and its drop relative to the original dataset.
-CLIPort1→9 CLIPort1→19
-[…]
-$\text{MAGE}_{\text{DINO}}$ 0.929 -1.2% 0.088 -37% 0.920 -1.2% 0.094 -25%
-TextOCVP 0.936 -1.5% 0.076 -23% 0.921 -1.1% 0.090 -15%
-F.4 Robustness to Number of Objects
-We evaluate the performance of TextOCVP and $\text{MAGE}_{\text{DINO}}$ on a CLIPort evaluation set consisting of scenes with a larger number of objects than encountered during training, i.e.
-the 2D table plane contains 8 objects instead of the default 6 objects.
-This evaluation benchmarks the robustness of models when generalizing to scenes with a previously unseen number of objects.
-The quantitative results are presented in Table 8.
-We report both the video prediction performance, as well as the performance drop relative to the original evaluation set.
-We observe that our model outperforms $\text{MAGE}_{\text{DINO}}$ on every metric for both prediction horizons.
-Notably, TextOCVP demonstrates significantly higher robustness to scenes with more objects, as reflected by the smaller decline in performance for the perceptual LPIPS metric.
-This result is further illustrated in Fig.
-9, where we show qualitative comparisons of video generations for scenes with 8 objects.
-As observed, our model correctly predicts sequences following the motion described in the text instructions, whereas $\text{MAGE}_{\text{DINO}}$ fails to generate accurate sequences according to the descriptions.
-These results further demonstrate the effectiveness of object-centric representations for video prediction, as TextOCVP is able to generalize to scenes with more objects by simply increasing the number of slots.
-This flexibility is enabled by initializing object slots via sampling from a learned Gaussian, allowing the use of a variable number of slots at test time while breaking symmetry and preserving permutation invariance.
-Although the object slots do not specialize during training, they reliably bind to meaningful entities through iterative attention during inference.
-This results in robust and scalable scene decomposition, enabling accurate modeling of complex scenes with varying object counts.
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the gray block in the brown bowl Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-$\text{MAGE}_{\text{DINO}}$
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOCVP
-t=1t=1551010202030304040a)
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the yellow block in the gray bowl Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-$\text{MAGE}_{\text{DINO}}$
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-TextOCVP
-t=1t=110102020303040405050b)
-Figure 9: Qualitative evaluation of $\text{MAGE}_{\text{DINO}}$ and TextOCVP on CLIPort sequences with more objects than those seen during training.
-TextOCVP correctly generates sequences according to the text instructions, whereas $\text{MAGE}_{\text{DINO}}$ misses the target bowl.
-Table 9: Impact of visual artifacts on TextOCVP’s performance on CLIPort. Cropping the artifact-prone bottom part of the predicted frames leads to a notable improvement in TextOCVP’s performance.
-CLIPort1→9 CLIPort1→19
-[…]
-Full Image 0.950 0.062 0.931 0.078
-Excluded Bottom 0.953 0.050 0.932 0.069
-F.5 Impact of Visual Artifacts
-As already discussed in the main paper, TextOCVP occasionally generates visual artifacts on the CLIPort dataset, most noticeably in the bottom-center region of the frame, where blurry patches often appear.
-To assess the impact of these visual artifacts on TextOCVP’s quantitative performance, we evaluate our model after removing the last bottom rows from the predicted frames—an area that contains only background pixels and the artifacts.
-As shown in Table 9, removing these rows leads to improved results, particularly for the perceptual LPIPS metric.
-These findings verify that TextOCVP generates future frame predictions that closely follow the text description, and that its overall performance is underestimated due to localized visual artifacts.
-Mitigating such artifacts through improved architectures or training strategies presents an interesting direction for future work.
-F.6 Text-To-Slot Attention
-An additional advantage of using object-centric representations is improved interpretability.
-This is illustrated by visualizing the text-to-slot attention weights, which reveal how textual information guides the model’s predictions.
-In Fig.
-11, we visualize the text-to-slot attention weights from different cross-attention heads for a single slot that represents a rotating red cube.
-We observe that the slot attends to distinct relevant text tokens from the language description on each attention head, including the object shape (head 2), size (head 1) and the action taking place (head 4).
-Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
-‘put the gray block in the brown bowl
-[…]
-TextOCVPt=1t=12020303040405050
-Figure 10: Qualitative result on CLIPort. Top row shows ground truth frames. TextOCVP completes the pick-and-place task, whereas MAGE fails to predict the robot motion.
-F.7 Additional Qualitative Evaluations
-Fig.
-10 shows an example where we evaluate TextOCVP and $\text{MAGE}_{\text{DINO}}$ for text-guided video prediction over a long prediction horizon of 50 frames.
-$\text{MAGE}_{\text{DINO}}$ fails to complete the task outlined in the textual description, as it stops generating consistent robot motion after 30 frames.
-In contrast, TextOCVP successfully predicts future frames where the robot completes the pick-and-place task.
-Figs. 12 and 13  show qualitative evaluations on CATER in which both MAGE and TextOCVP successfully predict sequence continuations following the instructions from the textual description.
-Fig. 14 illustrates an example where MAGE fails to generate a correct sequence, while TextOCVP successfully completes the task described by the text.
-Figs.
-15 and 16  show examples of TextOCVP’s control over the predictions.
-In both sequences, TextOCVP generates a correct sequence given the text instructions, and seamlessly adapts its generations to a modified version of the textual instructions.
+(a) Text-to-slot attention weights, averaged across heads, for different objects in a CATER sequence.
 ‘the small red rubber cube is rotating.
@@
 rotating
-Figure 11: Visualization of the text-to-slot attention weights between an object slot and the textual tokens.
-We visualize the text-to-slot attention weights for four different heads.
-Our predictor module focuses on distinct and relevant text tokens on each head, including the object size, shape or the action.
+(b) Text-to-slot attention weights between an object slot and the textual tokens for four different heads.
+Figure 5: Visualization of text-to-slot attention in TextOCVP.
+(a) Slots attend to relevant text tokens, grounding objects to their described motions.
+(b) Distinct attention heads focus on complementary textual cues, such as object attributes and actions.
+Fig.
+5(a) visualizes the text-to-slot attention weights—averaged across heads—for different slots in a CATER sequence, highlighting how textual instructions guide the model’s predictions.
+Slots corresponding to described objects attend strongly to the relevant tokens in the language, such as actions or target coordinates.
+To further dissect this behavior, Fig.
+5(b) shows the attention distribution of individual text-to-slot attention heads for a slot representing a rotating red cube.
+Different heads specialize in distinct attributes of the text, including the object’s shape (Head 2), size (Head 1), and the described motion (Head 4).
+Together, these results show that the text-to-slot attention mechanism effectively aligns textual information with object-centric representations.
+This enables accurate video prediction conditioned on natural language, and provides a degree of interpretability by revealing which parts of text influence each object representation.
+4.3.4 Controllability
+Refer to captiont=1t=12215153030Refer to captionOriginal Caption i) ‘the large purple rubber cone is picked up and placed to (2, 3).
+the small gold metal snitch is picked up and placed to (-1, 1).’ Refer to captionChanged Actions ii) ‘the large purple rubber cone is sliding to ( -1 , -3 ).
+the small gold metal snitch is rotating.’ Refer to captionChanged Objects & Actions iii) ‘the medium cyan rubber sphere is picked up and placed to (-2, 2).
+the medium purple metal cone is sliding to (2, 3).’ Refer to captionMore Moving Objects iv) ‘the large purple rubber cone is sliding to (-1, 1).
+the small gold metal snitch is rotating.
+the medium purple metal cone is picked up and placed to (-1, -3).’
+(a) Qualitative evaluation of TextOCVP’s controllability on CATER.
+Top row shows the ground truth sequence.
+We underline the changed actions with respect to the original caption.
+TextOCVP demonstrates fine-grained control by predicting different sequence continuations from the same reference frame, each conditioned on a different instruction.
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the cyan block in the brown bowl Original Caption Refer to captionRefer to captionRefer to captionRefer to caption‘put the blue block in the brown bowl Changed Block t=1t=11010202030304040
+(b) Qualitative evaluation of TextOCVP’s controllability on CLIPort.
+Top row shows ground truth frames.
+TextOCVP correctly generates sequences where the robot picks up the correct block and places it into the specified bowl.
+Figure 6: Qualitative evaluation of TextOCVP’s controllability on CATER and CLIPort datasets.
+A key objective of text-guided video prediction is to provide fine-grained control over the prediction process via language instructions that specify the relevant objects and their actions.
+Fig.
+6(a) illustrates TextOCVP’s controllability on CATER.
+Starting from the same reference frame, we generate multiple sequence continuations by varying the natural language instruction.
+These variations include changing the target objects and their actions, as well as instructions that specify a greater number of actions than seen during training.
+As shown in Fig.
+6(a), TextOCVP successfully identifies the relevant objects and executes the described actions.
+Notably, it distinguishes between two nearly identical purple cones, despite their identical shapes and color, and generates sequences consistent with the specified motions.
+A similar experiment on CLIPort is shown in Fig.
+6(b).
+Given a single frame with multiple colored blocks and bowls, modifying the instruction determines which block the robot arm picks and the destination bowl.
+In both cases, TextOCVP correctly selects the specified block, places it in the instructed bowl, and adapts the arm trajectory to the described action.
+These results show that key object attributes, such as size or color, are effectively captured through the text-to-slot attention mechanism.
+This enables accurate, per-object motion forecasting and highlights the benefit of combining object-centric representations with language guidance for controllable video prediction.
+Further quantitative and qualitative evaluations of TextOCVP’s controllability are provided in Appendices F.5 and F.10, respectively.
+5 Conclusion
+In this work, we presented TextOCVP, an object-centric model for text-conditioned video prediction.
+Given a single input image and a natural language description, TextOCVP generates future frames by parsing the scene into slot-based object representations and modeling their dynamics conditioned on the text instruction.
+This is accomplished through a text-conditioned object-centric transformer that predicts future object states by modeling spatio-temporal relationships between objects while incorporating textual guidance.
+Through extensive evaluations, we demonstrated that TextOCVP outperforms other existing approaches for text-driven video prediction from a single frame, highlighting our model’s ability to predict over long prediction horizons and adapt its predictions based on the provided description.
+Moreover, we validated our architectural choices through ablation studies, highlighting the importance of combining textual and object-centric information, and demonstrating strong robustness and interpretability.
+With its structured latent space and superior controllability, TextOCVP offers a promising step toward controllable object-centric manipulation in simulated robotic environments, supporting more efficient planning, reasoning and decision-making.
+Acknowledgment
+This work was funded by grant BE 2556/16-2 (Research Unit FOR 2535 Anticipating Human Behavior) of the German Research Foundation (DFG).
+Computational resources were provided by the German AI Service Center WestAI.
+Appendix A Limitations and Future Work
+A.1 Limitations
+While TextOCVP demonstrates promising results for text-guided object-centric video prediction, it presents some limitations, which we plan to address in future work:
+Prediction Artifacts
+TextOCVP occasionally generates artifacts in the predicted frames, such as blurriness, inconsistent object appearances, lack of textured details, or visual artifacts in the background.
+We believe that these limitations stem from the video rendering module, which might lack the representational power to reconstruct precise image details from the object-centric latent space representation.
+Limited Temporal Consistency
+While TextOCVP produces plausible predictions, we observe that its temporal consistency can degrade when forecasting for long prediction horizons ($T>30$), occasionally resulting in object jittering or instability.
+We attribute this limitation to the fact that TextOCVP is trained to predict only up to nine future frames ($T=9$) and only optimizing reconstruction losses, which do not penalize such temporal inconsistencies.
+Inherent Limitations of Slot-based Models
+Slot-based object-centric models, while effective for disentangling scene structure and producing interpretable object-level representations, come with several inherent limitations.
+They often struggle to represent fine textures, objects with intricate geometry, small visual details, or tightly interlocking and highly deformable objects, as these phenomena exceed the granularity and capacity of a single slot embedding.
+In such cases, relevant information is often spread across multiple slots, leading to oversegmentation issues, or compressed into an overly coarse latent, leading to imperfect and blurred reconstructions, lack of representation detail, or unstable object binding.
+More fundamentally, current slot-based architectures typically operate at a single spatial scale, which prevents them from naturally encoding part-whole hierarchies, capturing object structure at different levels of detail, or representing complex compositional structures.
+Addressing these challenges through hierarchical object-centric parsing, adaptive slot resolution, or richer generative decoders remains an important future direction for object-centric learning and video prediction.
+A.2 Future Work
+In future work we aim to address these limitations.
+Our TextOCVP model is designed with a modular architecture, enabling for both improvements and flexible swapping of components—such as the parsing, predictor, or video rendering modules—to seamlessly improve the entire pipeline.
+To address the prediction artifacts, we plan to extend our TextOCVP framework with more powerful decoder modules, such as autoregressive transformers (Singh et al., 2022) or diffusion models (Jiang et al., 2023), as well as scale our predictor module.
+Furthermore, we plan to incorporate temporal discriminators (Clark et al., 2020) to improve the temporal consistency of the predicted video frames.
+We believe that exploring these architectural modifications can overcome the aforementioned limitations and will enable the scaling of TextOCVP to real-world robotic environments.
+Appendix B Implementation Details
+We employ $\text{TextOCVP}_{\text{SAVi}}$ for the experiments on CATER and $\text{TextOCVP}_{\text{DINO}}$ for experiments on CLIPort.
+Below we discuss the implementation details for each of these variants.
+B.1 $\text{TextOCVP}_{\text{DINO}}$
+The $\text{TextOCVP}_{\text{DINO}}$ variant consists of our proposed text-conditioned predictor module and an object-centric decomposition module that extends the DINOSAUR (Seitzer et al., 2023) framework (Extended DINOSAUR) for recursive object-centric video decomposition and video rendering.
+Text-Conditioned Predictor
+The predictor is composed of $N_{\text{P}}=8$ identical layers, each containing 8-head attention mechanisms and an MLP with a single hidden layer of dimension 1024 and a ReLU activation function.
+Furthermore, the predictor uses an embedding dimensionality of 512, context window size of ten frames, and applies a residual connection from the predictor input to its output.
+Text Encoder
+$\text{TextOCVP}_{\text{DINO}}$ leverages a pretrained and frozen small version of T5 encoder (Raffel et al., 2020), which consists of six T5 blocks.
+This text encoder uses a vocabulary with size 32,128.
+Scene Parsing
+The scene parsing module generates $N_{\mathbf{S}}=10$ slots of dimension 128.
+As feature extractor, we use DINOv2 ViT-Base (Oquab et al., 2024), featuring 12 layers, using a patch size of 14, and producing patch features with dimension $D_{h}=768$.
+The Slot Attention corrector module processes the first video frame with three iterations in order to obtain a good initial object-centric decomposition, and a single iteration for subsequent frames, which suffices to recursively update the slot representation.
+The initial object slots $\textbf{S}_{0}$ are randomly sampled from a Gaussian distribution with learned mean and covariance.
+We use a single Transformer encoder block as the transition function, which consists of four attention heads and an MLP with a hidden dimension of 512.
+Video Rendering
+The video rendering module consists of two distinct decoders.
+First, a four-layer MLP-based Spatial Broadcast Decoder (Watters et al., 2019) with hidden dimension 1024 reconstructs the patch features from the slots.
+Then, a CNN-based decoder reconstructs full-resolution images from these features.
+It consists of four convolutional layers, each using $3\times 3$ kernels, a ReLU activation function and bilinear upsampling.
+A final convolutional layer and bilinear interpolation are applied to map the outputs to the RGB channels and spatial dimensions of the image.
+Training
+We train our model for object-centric decomposition using video sequences of length five frames for 1000 epochs.
+We use batch size of 16, the Adam optimizer (Kingma and Ba, 2015), and a base learning rate of $4\times 10^{-4}$, which is linearly warmed-up for the first 10000 steps, followed by cosine annealing for the remaining of the training process.
+Moreover, we clip the gradients to a maximum norm of 0.05.
+The predictor module is trained given the frozen and pretrained object-centric decomposition model for 700 epochs to predict the subsequent nine frames using a single seed frame.
+The predictor is trained using the same hyper-parameters as for object-centric decomposition.
+In the predictor loss function $\mathcal{L}_{\text{TextOCVP}}$, we set $\lambda_{\text{Img}}=1$ and $\lambda_{\text{Slot}}=1$.
+B.2 $\text{TextOCVP}_{\text{SAVi}}$
+$\text{TextOCVP}_{\text{SAVi}}$ uses the same text-conditioned predictor and text-encoder architectures as $\text{TextOCVP}_{\text{DINO}}$, but employs SAVi (Kipf et al., 2022) as the object-centric decomposition module.
+Scene Parsing
+The scene parsing module generates $N_{\mathbf{S}}=8$ slots of dimension 128.
+Following Kipf et al.
+(2022), we use as feature extractor a four-layer CNN with ReLU activation function, where each convolutional layer features 32 $5\times 5$ kernels, $\text{stride}=1$, and $\text{padding}=2$.
+The Slot Attention corrector follows the same structure as in $\text{TextOCVP}_{\text{DINO}}$.
+Video Rendering
+Following Kipf et al.
+(2022), we utilize a CNN-based Spatial Broadcast Decoder (Watters et al., 2019) with four convolutional layers with 32 kernels of size $5\times 5$ , $\text{stride}=1$, and $\text{padding}=2$.
+A final convolutional layer maps from the hidden 32-channel representation to four output channels (RGB + alpha mask).
+Training
+We train our model for object-centric decomposition using video sequences of length ten frames for 1000 epochs, using batch size of 64, and an initial learning rate of $10^{-4}$, which is warmed up for 2500 steps, followed by cosine annealing for the remaining of the training process.
+Moreover, we clip the gradients to a maximum norm of 0.05.
+The predictor module is trained given the frozen and pretrained object-centric decomposition model for 1400 epochs to predict the subsequent nine frames using a single seed frame.
+The predictor is trained using the same hyper-parameters as for object-centric decomposition.
+In the predictor loss function $\mathcal{L}_{\text{TextOCVP}}$, we set $\lambda_{\text{Img}}=1$ and $\lambda_{\text{Slot}}=1$.
+Table 4: Number of learnable parameters in TextOCVP and baselines for experiments on CLIPort.
+Model # Parameters
+TextOCVP 33.76M
+Non-OC 34.16M
+$\text{MAGE}_{\text{DINO}}$ 32.11M
+SEER 405.89M
+Appendix C Baselines
+[…]
+[…]
+[…]
+MAGE is an autoregressive text-guided video prediction framework that utilizes a VQ-VAE (Van Den Oord et al., 2017) encoder-decoder architecture to learn efficient visual token representations.
+A cross-attention module aligns textual and visual embeddings to produce a spatially-aligned motion representation termed Motion Anchor (MA), which is fused with visual tokens via an axial transformer for video generation.
+For experiments on CATER, we use a codebook size of $512\times 256$ with a downsampling ratio of four, whereas on CLIPort we use a codebook size of $512\times 1024$.
+To ensure a fair comparison with TextOCVP on CLIPort, we replace MAGE’s standard CNN encoder and decoder with the DINOv2 ViT encoder and CNN decoder used in our TextOCVP model.
+We refer to this modified version as $\text{MAGE}_{\text{DINO}}$.
+Table 5 presents a comparison between the original MAGE model and $\text{MAGE}_{\text{DINO}}$ on CLIPort.
+The results demonstrate that $\text{MAGE}_{\text{DINO}}$ significantly outperforms the original variant, enabling a fair comparison with TextOCVP and other baselines on this benchmark.
+MAGE and $\text{MAGE}_{\text{DINO}}$ share several architectural similarities with our proposed approach—using similar encoder and decoder modules, text-conditioning, and an autoregressive transformer for prediction.
+However, these models differ from TextOCVP in two fundamental ways:
+Scene representation:  TextOCVP operates on object-centric slot representations, whereas MAGE and $\text{MAGE}_{\text{DINO}}$ rely on holistic, dense scene tokens (vector-quantized image latents).
+Whereas slots provide a factorized latent structure with one representation per object, the holistic VQ-token grid entangles object information across many spatial tokens without explicit boundaries.
+This distinction enables a clean comparison between object-centric and holistic autoregressive models.
+Mechanism for text conditioning:  Both methods use cross-attention to incorporate language guidance, but differ in how textual information interacts with the predictor.
+MAGE and $\text{MAGE}_{\text{DINO}}$ compute a single global Motion Anchor via one cross-attention step between text and image latents, and inject this global signal uniformly into all decoding steps.
+In contrast, TextOCVP applies text-to-slot cross-attention within every transformer block, allowing the model to repeatedly integrate and select the textual information most relevant at each processing stage.
+Table 5: Comparison of MAGE variants on CLIPort. $\text{MAGE}_{\text{DINO}}$ clearly outperforms the original MAGE variant across all metrics.
+Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
+MAGE 7.116 0.453 0.713
+$\text{MAGE}_{\text{DINO}}$ 23.723 0.940 0.064
+C.2 SEER
+SEER is a diffusion-based model for language-guided video prediction.
+It employs an Inflated 3D U-Net derived from a pretrained text-to-image 2D latent diffusion model (Rombach et al., 2022), extending it along the temporal axis and integrating temporal attention layers to simultaneously model spatial and temporal dynamics.
+For the language conditioning module, SEER introduces a novel Frame Sequential Text (FSText) Decomposer, which decomposes global instructions generated by the CLIP text encoder (Radford et al., 2021) into frame-specific sub-instructions.
+These are aligned with frames using a transformer-based temporal network and injected into the diffusion process via cross-attention layers.
+We initialize SEER from a checkpoint pretrained on the Something-Something V2 dataset (Goyal et al., 2017), and further fine-tune it for a few epochs.
+We observed that incorporating a text loss enhanced SEER’s performance, while other hyper-parameters were kept consistent with its original implementation.
+C.3 Non-OC
+Non-OC is a variant of our proposed TextOCVP model in which the slot-based object-centric latent representations are replaced with a single, high-dimensional slot embedding.
+This design allows us to isolate the contribution of object-centric structure in the latent space.
+Non-OC mirrors the TextOCVP architecture, using the same visual backbone, text-guided autoregressive predictor and decoder, and it is trained with identical hyper-parameters and training schedule.
+The only difference lies in the scene parsing module: instead of slot attention, Non-OC applies an additional convolutional block followed by average pooling to produce a single latent vector per frame.
+This results in one 512-dimensional embedding, in contrast to the set of 128-dimensional object slots used in the object-centric model.
+Appendix D Datasets
+D.1 CATER
+CATER (Girdhar and Ramanan, 2020) is a dataset that consists of long video sequences, each described by a textual caption.
+The video scenes consist of multiple 3D geometric objects in a 2D table plane, which is split into a $6\times 6$ grid with fixed axis, allowing the exact description of object’s positions using coordinates.
+The text instruction describes the movement of specific objects through four atomic actions: ‘rotate’, ‘pick-place’, ‘slide’, and ‘contain’.
+The caption follows a template consisting of the subject, action, and an optional object or end-point coordinate, depending on the action.
+The movement of the objects starts at the same time step.
+Furthermore, the initial positions are randomly selected from the plane grid, and the camera position is fixed for every sequence.
+In our work, we employ CATER-hard, which is a complete version of the CATER dataset, containing 30000 video-caption pairs, with video frames resized to $64\times 64$.
+It includes 5 possible objects: cone, cube, sphere, cylinder, or snitch, which is a special small object in metallic gold color, shaped like three intertwined tori.
+Furthermore, every object is described by its size (small, medium, or large), material (metal or rubber), and color (red, blue, green, yellow, gray, brown, purple, cyan, or gold if the object is the snitch), and this description is included in the textual caption.
+Every atomic action is available.
+The ‘rotate’ action is afforded by cubes, cylinders and the snitch, the ‘contain’ action is only afforded by the cones, while the other two actions are afforded by every object.
+Every video has between 3 and 8 objects, and two actions happen to different objects at the same time.
+The vocabulary size is 50.
+D.2 CLIPort
+CLIPort (Shridhar et al., 2022) is a robot manipulation dataset, consisting of video-caption pairs, i.e.
+long videos whose motion is described by a textual video caption.
+There are many variants of the CLIPort dataset, but we focus on the Put-Block-In-Bowl variant.
+We generate 21000 video-caption pairs with resolution $336\times 336$.
+Every video contains 6 objects on a 2D table plane, and a robot arm.
+Objects can be either a block or a bowl, and there is at least one of them in every sequence.
+The starting position of each object is random, with the only constraint being that it must be placed on the table.
+Each video describes the action of the robot arm picking a block, and putting it in a specific bowl.
+The video caption follows the template ‘put the [color] block in the [color] bowl’.
+Each individual object in the scene has a different color.
+In the train and validation set, the block and the bowl that are part of the caption can have one of the following colors: blue, green, red, brown, cyan, gray, or yellow, while in the evaluation set with unseen colors they can have blue, green, red, pink, purple, white, or orange color.
+The other 4 objects, called distractors, can have any color.
+During a video sequence, it can be possible that the robot arm goes out of frame, and comes back in later frames, thus requiring the model to leverage long range dependencies.
+The vocabulary size is 15.
+Appendix E Evaluation Metrics
+To measure TextOCVP’s video prediction performance and compare it with existing approaches, we evaluate the visual quality of predicted video frames using popular image- and video-based metrics.
+PSNR and SSIM (Wang et al., 2004) measure pixel-wise and statistical differences between the predicted and ground-truth video frames, respectively.
+LPIPS (Zhang et al., 2018) is a perceptual metric that measures the visual similarity between two images based on deep features from pretrained neural networks, usually VGG (Simonyan and Zisserman, 2015).
+Unlike pixel-wise metrics, LPIPS compares activations from multiple layers and captures differences in texture, structure, and semantics, thus aligning closely with human perception.
+While these metrics focus on frame-level visual quality, video prediction evaluation requires quantifying temporal consistency and motion realism.
+JEDi (Luo et al., 2025) measures the quality and realism of generated videos by comparing feature distributions of generated and real videos, capturing both visual fidelity and motion dynamics.
+We prefer JEDi over FVD (Unterthiner et al., 2019), as JEDi is less sensitive to small evaluation datasets, and more stable in low-motion synthetic scenarios.
+In our evaluations and comparisons with baselines, we favor the LPIPS and JEDi scores, which correlate well with human perception, while reporting other metrics for completeness.
+$t=1$
+22
+33
+44
+55
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+GT
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+SAVi
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Ext. DINO
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Ext. DINO
+Masks
+Figure 7: Comparison between SAVi and our Extended DINOSAUR (Ext.
+DINO) decomposition modules for reconstructing a CLIPort sequence.
+We visualize the reconstructed frames, as well as the slot masks obtained by Extended DINOSAUR.
+SAVi fails to reconstruct most objects, whereas Extended DINOSAUR accurately reconstructs the scene, while representing each object.
+Appendix F Additional Results
+F.1 SAVi vs. DINOSAUR
+Current object-centric approaches for video prediction are limited to relatively simple synthetic datasets, and struggle to scale beyond scenes featuring simple 3D shapes with simple deterministic motion.
+We attribute this limitation primarily to the object-centric modules used for learning object representations.
+Motivated by this observation, we extend the recent DINOSAUR (Seitzer et al., 2023) framework to recursively process video data and reconstruct video frames from their corresponding object-centric representations.
+To demonstrate the significance of the object-centric module in scaling to more complex datasets, we compare both SAVi and our Extended DINOSAUR trained on CLIPort.
+As illustrated in Fig.
+7, SAVi struggles to accurately represent the objects on the table, missing multiple objects and changing their shape and color.
+In contrast, the Extended DINOSAUR model successfully reconstructs the scene, closely resembling the input, while accurately representing each object.
+The visual features extracted by the DINOv2 (Oquab et al., 2024) encoder contain high-level semantic information, and during training, the slots are specifically optimized to efficiently encode this information.
+This design enables the model to scale and handle more complex object-centric video data effectively.
+F.2 Computational Efficiency
+In Table 6 we report the average inference time for $T=1$ and $T=9$ frame predictions on CATER and CLIPort using a NVIDIA-A6000-48Gb GPU.
+[…]
+Table 6: Average inference (Inf.) time for $T=1$ and $T=9$ frame predictions on CATER and CLIPort.
+CATER Inf. [ms] CLIPort Inf. [ms]
+Model $T=1$ $T=9$ $T=1$ $T=9$
+MAGE 15.4 $\pm$ 1 111.1 $\pm$ 11 34.4 $\pm$ 3 198.8 $\pm$ 2
+[…]
+F.3 Quantitative Comparison
+In the main paper, we present quantitative evaluations on the CATER and CLIPort datasets using three evaluation metrics, shown in two separate tables.
+For completeness, we report in Table 7 a text-guided video prediction evaluation of TextOCVP and multiple baseline models on CATER and CLIPort using four distinct evaluation metrics for different prediction horizons.
+[…]
+[…]
+[…]
+F.4 Object-Centric Evaluation
+A key advantage of object-centric approaches in video prediction is their ability to generate segmentation masks alongside frame predictions.
+For TextOCVP, we derive predicted segmentation masks by applying an argmax operation over all object slot masks produced by the decoder during prediction and subsequently filtering the resulting masks by assigning values with a small magnitude to the background.
+These predicted masks are evaluated against ground-truth segmentation masks using the Intersection over Union (IoU) metric.
+We compare our model against two baselines on CLIPort.
+The Copy-Seed baseline simply replicates the predicted segmentation mask from the seed frame across all frames, while the OC-module-only variant generates segmentation masks by feeding ground-truth frames directly into the object-centric module, i.e., Extended DINOSAUR.
+As shown in Table 8, TextOCVP outperforms both baselines across prediction horizons of $T=9$ and $T=19$ frames, demonstrating the model’s ability to model object dynamics, maintaining spatial consistency and object coherence over time.
+The predicted masks of TextOCVP align more closely with the ground-truth segmentations, benefiting from better temporal stability throughout the prediction horizon.
+In Figs.
+11 and 12 , we illustrate TextOCVP’s object-centric behavior on CATER and CLIPort by showing the predicted segmentation masks, as well as different per-object predictions.
+Table 8: Object-centric evaluation on CLIPort.
+Comparison of Intersection over Union (IoU) scores between predicted and ground-truth segmentation masks for different prediction horizons ($T=9$ and $T=19$).
+CLIPort1→9 CLIPort1→19
+Method $\uparrow IoU\uparrow$
+Copy-Seed 0.540 0.525
+OC-module-only 0.554 0.553
+TextOCVP 0.573 0.569
+F.5 Quantitative Evaluation of Controllability
+In Sec.
+4.3.4, we qualitatively show how TextOCVP is able to adapt its predictions based on the language instruction it receives as input.
+We further make an initial attempt to quantitatively evaluate the controllability of TextOCVP.
+Starting from a CLIPort evaluation set of 100 sequences, we create three language instruction variants for each sequence, differing in the specified target bowl.
+Given the same initial frame, TextOCVP then generates future predictions conditioned on the adapted instructions.
+To assess performance, we ground the slot masks produced by the video rendering decoder to their corresponding objects in the scene and compute two key distances: (1) between the masks of the robot arm and the specified block, and (2) between the picked block and the target bowl.
+These measures are used to estimate the pick-and-place success rate.
+A generated sequence is considered successful only if the robot arm remains sufficiently close to the correct block over multiple frames and the block is significantly close to the target bowl toward the end of the sequence.
+As shown in Table 9, TextOCVP demonstrates consistent success rates across different instruction variations, indicating strong robustness and fine-grained controllability.
+Given identical starting scene, the model effectively adapts its predictions to the changing text instructions, successfully placing the block into the specified bowls in most of the cases.
+We note that an equivalent experiment could not be performed with $\text{MAGE}_{\text{DINO}}$, as it does not generate object masks during prediction.
+Table 9: Quantitative evaluation of TextOCVP’s controllability on CLIPort.
+We report the mean pick-and-place success rates (averaged over 5 runs) for the original evaluation set and three instruction variants differing in the target bowl.
+TextOCVP maintains consistent performance across variants, reflecting strong robustness to instruction changes.
+Pick-and-Place success rate
+Method Original set Variant 1 Variant 2 Variant 3 Mean variants
+TextOCVP 0.83 0.86 0.80 0.81 0.82
+F.6 Robustness to Number of Objects
+In Sec.
+4.3.2, we quantitatively assess the generalization and robustness of TextOCVP in video prediction tasks involving novel scene compositions.
+Our results highlight the benefits of object-centric representations over holistic scene-based approaches.
+This finding is further illustrated in Fig.
+9, which presents qualitative comparisons of video generations for scenes containing eight objects, in contrast to the six-object configurations seen during training.
+As observed, our model correctly predicts sequences following the motion described in the text instructions, whereas $\text{MAGE}_{\text{DINO}}$ fails to generate accurate sequences according to the descriptions.
+These results further demonstrate the effectiveness of object-centric representations for video prediction, as TextOCVP is able to generalize to scenes with more objects by simply increasing the number of slots.
+This flexibility is enabled by initializing object slots via sampling from a learned Gaussian distribution, allowing the use of a variable number of slots at test time while breaking symmetry and preserving permutation invariance.
+Although the object slots do not specialize during training, they reliably bind to meaningful entities through iterative attention during inference.
+This results in robust and scalable scene decomposition, enabling accurate modeling of complex scenes with varying object counts.
+F.7 Impact of Visual Artifacts
+As already discussed in the main paper, TextOCVP occasionally generates visual artifacts on the CLIPort dataset, most noticeably in the bottom-center region of the frame, where blurry patches often appear.
+To assess the impact of these visual artifacts on TextOCVP’s quantitative performance, we evaluate our model after removing the last bottom rows from the predicted frames—an area that contains only background pixels and the artifacts.
+As shown in Table 10, removing these rows leads to improved results, particularly for the perceptual LPIPS metric.
+These findings verify that TextOCVP generates future frame predictions that closely follow the text description, and that its overall performance is underestimated due to localized visual artifacts.
+Table 10: Impact of visual artifacts on TextOCVP’s performance on CLIPort. Cropping the artifact-prone bottom part of the predicted frames leads to a notable improvement in TextOCVP’s performance.
+CLIPort1→9 CLIPort1→19
+[…]
+Full Image 0.950 0.062 0.931 0.078
+Excluded Bottom 0.953 0.050 0.932 0.069
+F.8 Impact of Video Rendering Module
+In the previous section (Appendix F.7), we described how visual artifacts affect TextOCVP’s performance on the CLIPort dataset.
+We argue that these artifacts mostly originate from limitations in the video rendering module, and that they can be mitigated by adopting a more expressive decoder architecture.
+To support this claim, we conducted an additional experiment on CLIPort in which we replaced the simple CNN decoder—described in Appendix B.1—with a more expressive alternative, while keeping the rest of TextOCVP’s architecture unchanged.
+The new decoder, inspired by VQGAN (Esser et al., 2021), integrates convolutional, residual, non-local attention, and upsampling blocks, and contains nearly twice as many learnable parameters as the original module ($\sim 25\text{M}$ vs.
+$\sim 13\text{M}$).
+TextOCVP models using both decoder variants are qualitative compared in Fig.
+10, which shows that the more expressive decoder substantially reduces the visual artifacts observed in the original model’s predictions.
+In Table 11 we quantitatively compare both decoder variants.
+The more expressive decoder model achieves noticeably better LPIPS and JEDi scores, demonstrating superior perceptual frame and video quality.
+Overall, these findings indicate that employing a more capable decoder can mitigate some of TextOCVP’s visual limitations.
+They also highlight the benefits of our modular architecture, which enables such improvements to be incorporated seamlessly without modifying the underlying predictor or representation.
+Exploring more powerful rendering modules—such as diffusion-based decoders (Jiang et al., 2023) or refining the VQGAN-style architecture (Esser et al., 2021)—is a promising direction for future work.
+Table 11: Comparison of TextOCVP with a variant using a more expressive video rendering module on CLIPort.
+The stronger, VQGAN-inspired (Esser et al., 2021), decoder leads to improved perceptual quality.
+CLIPort1→9 CLIPort1→19
+TextOCVP Decoder $downarrow LPIPS\downarrow JEDi\downarrow$
+Simple CNN 0.062 1.36 0.078 2.23
+VQGAN-based 0.043 0.86 0.065 2.22
+F.9 Evaluation of Predictions over Time
+In Fig.
+8, we present plots of LPIPS and slot-mask IoU scores over prediction horizons of up to $40$ future frames.
+These curves show that our model’s predictions steadily deviate from the ground truth as the prediction horizon increases, reflecting the natural accumulation of errors inherent to autoregressive prediction.
+We also note that quantitative metrics such as LPIPS or IoU become less informative when predicting beyond 2020-$30$ future frames.
+At these prediction horizons, even slight differences in object velocity or trajectory can compound and lead to large discrepancies in frame-wise metrics, despite still producing qualitatively consistent futures.
+This behavior is typical of autoregressive video models and reflects the sensitivity of pixel-level scores to small temporal deviations, rather than a qualitative failure of the predicted dynamics.
+Refer to caption0.060.080.100.120.140.161510152025303540Predicted time-step
+LPIPS $\downarrow$
+(a) LPIPS ($\downarrow$) metric across prediction horizons
+Refer to caption0.480.500.520.540.560.581510152025303540Predicted time-step
+IoU $\uparrow$
+(b) IoU ($\uparrow$) metric across prediction horizons
+Figure 8: LPIPS and slot-mask IoU metrics across prediction horizon of 40 future frames. The plots show both the average values (bold line) as well as the standard deviation (shaded areas).
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the gray block in the brown bowl Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+$\text{MAGE}_{\text{DINO}}$
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+TextOCVP
+t=1t=1551010202030304040a)
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the yellow block in the gray bowl Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+$\text{MAGE}_{\text{DINO}}$
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+TextOCVP
+t=1t=110102020303040405050b)
+Figure 9: Qualitative evaluation of $\text{MAGE}_{\text{DINO}}$ and TextOCVP on CLIPort sequences with more objects than those seen during training (eight instead of six).
+TextOCVP correctly generates sequences according to the text instructions, whereas $\text{MAGE}_{\text{DINO}}$ misses the target bowl.
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the red block in the brown bowl Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+TextOCVP w/
+VQGAN dec.
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+TextOCVP w/
+original dec.
+t=1t=1551010202030304040a)
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘put the yellow block in the gray bowl Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+TextOCVP w/
+VQGAN dec.
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+TextOCVP w/
+original dec.
+t=1t=110102020303040405050b)
+Figure 10: Qualitative comparison between original TextOCVP implementation and a TextOCVP variant with a more expressive, VQGAN-inspired, video rendering module.
+While both variants correctly predict the described motion, the new decoder improves the predicted frames’ quality by mitigating most of visual artifacts, especially on the bottom-center region and in the robot arm.
+F.10 Additional Qualitative Evaluations
+Fig.
+13 shows an example where we evaluate TextOCVP and $\text{MAGE}_{\text{DINO}}$ for text-guided video prediction over a long prediction horizon of 50 frames.
+$\text{MAGE}_{\text{DINO}}$ fails to complete the task outlined in the textual description, as it stops generating consistent robot motion after 30 frames.
+In contrast, TextOCVP successfully predicts future frames where the robot completes the pick-and-place task.
+Figs. 14 and 15  show qualitative evaluations on CATER in which both MAGE and TextOCVP successfully predict sequence continuations following the instructions from the textual description.
+Fig. 16 illustrates an example where MAGE fails to generate a correct sequence, while TextOCVP successfully completes the task described by the text.
+Figs.
+17 and 18  show examples of TextOCVP’s control over the predictions.
+In both sequences, TextOCVP generates a correct sequence given the text instructions, and seamlessly adapts its generations to a modified version of the textual instructions.
+GT
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption‘the large purple rubber cone is picked upand placed to (2, 3).
+the small gold metal snitch is picked up and placed to (-1, 1).’Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionPredictionsRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Slot
+Masks
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Object 1
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Object 2
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Object 3
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+Object 4
+t=1t=1551010151520203030
+Figure 11: TextOCVP’s object-centric behavior on a CATER sequence.
+The first row shows the ground truth sequence, followed by TextOCVP’s predicted frames and segmentation masks.
+The subsequent rows display the reconstructed objects from four of the predicted slots across various time steps, highlighting the ability of TextOCVP to model the dynamics of individual objects in the scene through slot representations.
+GT
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption ‘put the cyan block in the red bowl Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption PredictionsRefer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Slot
+Masks
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Object 1
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Object 2
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Object 3
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Object 4
+t=1t=1551010151520203535
+Figure 12: TextOCVP’s object-centric behavior on a CLIPort sequence.
+The first row shows the ground truth sequence, followed by TextOCVP’s predicted frames and segmentation masks.
+The subsequent rows illustrate the represented objects from four of the predicted slots across various time steps.
+Although only eight slots are required for this dataset (six objects, one robot arm, and one background), we use ten slots in CLIPort experiments which proved to beneficial; the two extra slots represent background.
+We emphasize that the object segmentations are computed at the patch level, thus the pixelated appearance in the visualizations.
+Refer to captionRefer to captionRefer to captionRefer to captionRefer to caption
+‘put the gray block in the brown bowl
+[…]
+TextOCVP t=1t=12020303040405050
+Figure 13: Qualitative result on CLIPort. Top row shows ground truth frames. TextOCVP completes the pick-and-place task, whereas $\text{MAGE}_{\text{DINO}}$ fails to predict the robot motion.
 Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption